A Selected Bibliography of Publications by, and about, Lise Meitner

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 June 2024 Version 1.17

Title word cross-reference

$21.95 [Bal12]. $29.99 [Sap14]. $30.00 [Kra02]. $34 [Hof97]. 4n + 1 [HM35a]. $7.50 [Lin60]. $8.50 [Nig03]. = mc^{2} [For80]. ^{261,262} [NFG^{+}08]. 265−−279 [SN17]. ^{re} [Rad04]. _{1} [HM23, Mei23b, Mei23a]. _{2} [HM13a, HM23]. α [HM20b, Mei06a, Mei07, Mei15a, Mei22c, MF24, Mei26d, MF26b, Mei28a, Mei28c, MP31, MP33a, MD35a, MSH38, SN17]. β [vBHM11a, vBHM11b, vBHM11c, vBHM13, vBHM14, vBHM15, FR13a, HM08c, HM08b, HM09d, HM09e, HM10, HM12b, HM20b, HM24, HM25b, Mah22, Mei06a, vBHM12, Mei15b, Mei22d, Mei22b, Mei23d, Mei23b, Mei23a, Mei24a, Mei26c, Mei28c, Mei28b, Mei29e, Mei29d, MO30, Mei32a, MD35b, Mei37, Mei43a, Mei43b, Mei43d, Mei43c]. · [BFG^{+}33]. γ [HM23, MH13, Mei22c, Mei22e, Mei22b, Mei24d, Mei24a, Mei25, Mei28c, Mei28b, Mei28d, Mei29e, Mei29b, Mei29a, MH30b, MH30a, MH31b, MH31a, MP31, Mei32a, MH32, MP33b, MW33, MHK33, MK33, MP33a, Mei34c, MD35d, Mei37, Mei43a, Mei43b, Mei43c]. K [JKS15].

-decay [Mah22, SN17]. -Euro [Her06]. -Feinstruktur [MP31]. -particles [Mei28a]. -Radiation [MF24, MH30b, Mei43a, Mei29d, MP33b]. -Ray [Mei23d, Mei43b, Mei43d]. -Rays [Mei43b, HM12b, Mei06a, MH31a, MW33, Mei34c, MSH38]. -Raysand [Mei43c]. -RaySpectra [Mei43c]. -Spektrum [MP31]. -Strahl-Spektra [Mei22c]. -Strahl-Spektren [Mei22d]. -Strahlen [vBHM11a, vBHM11b, vBHM11c, vBHM13, vBHM14, vBHM15, HM08c, HM08b, HM12b, HM20b, HM23, vBHM12, MH13, Mei15b, MF24, Mei24d, Mei25, Mei26d, MF26b, Mei28c, Mei29e, Mei29d, MO30, Mei32a, MP33b, MP33a, Mei37, Mei43a, HM09d, Mei06a, Mei07, Mei15b, Mei22e, Mei22b, Mei24a, Mei26c, Mei28b, Mei29b, Mei29a, MH31a, MW33, MHK33, MK33, MP33a, Mei34c, MD35d, MSH38]. -Strahlenspektrum [HM24, Mei23b, Mei23a, Mei28c, Mei28b, Mei28d]. -Strahlenspektrums [Mei23d]. -Strahlenumwandlungen [Mei28c]. -Strahlspektren [HM25b, Mei43a]. -Strahlung [HM09e, HM10, Mei15a, Mei25, MH32, Mei22c, Mei29e, MH30b, MH30a, MH31b]. -Zerfall [MD35a, MD35b].

/Brochure [Ano79a]. /Butterflies [Ano46b]. /Celebration [Ano79a]. /Das [DRF90a, DRF90b]. /Death [Ano46b]. /Deutsche [Ano79a]. /Deutsches [DRF90a, DRF90b]. /Dr. [Ano46b]. /Einstein [Ano79a]. /Ernst [HFK58]. /Experiment [Ano46a]. /Feier [Ano79a]. /German [Ano79a]. /Goethe [Ano46b]. /Hahn [Bas62]. /Hasse [KZN^{+}88]. /Hubert [Ano79a]. /Lise [HFK58]. /Max [Bas62]. /Nobel [Ano46b]. /Pasteur [Ano46b]. /Phys. [Bas62]. /Physical [Bas62]. /President [Ano46b]. /The [Ano46b]. /Zentrum [DRF90a, DRF90b].

0 [All96, Gav04, Gav05, Gru04, Her06, Hes96, KK97a, Kra02, Rou96, Her06]. 0-19-513748-5 [Gav04]. 0-520-0890-5 [Rou96]. 0-520-08906-5 [Hes96, KK97a]. 0-521-01706-8 [Gav05]. 0-674-07404-1 [All96]. 0-7167-3103-7 [Kra02].

1 [All96, B^{+}79, Swe06]. 1-84046-574-3 [Swe06]. 1/B [Ano46a]. 1/Zerstörung [Ano46a]. 100th [Ano79b, Ano79a, B^{+}79, EHM^{+}79, Lam79]. 125th [Lem03]. 12th [B^{+}79]. 13 [Rad99]. 1924 [Ern93]. 1925 [Mei26e]. 1928 [Mei28a]. 1930 [Ano31]. 1932 [For80]. 1933 [For80]. 1939 [MF69]. 1943 [Har43]. 1945 [DF14]. 1946 [CSW97b]. 1947 [Lem10]. 1948 [Lem98, WSB^{+}99, Hie00]. 1948/Messing [WSB^{+}99]. 1949 [Mei49]. 1959 [Mei59]. 1960 [Sim14a]. 1960s [Mla98, Mon98]. 1968 [Fri70]. 1969 [Ano69]. 1972 [Whi00]. 1979 [B^{+}79, Lam79]. 1988 [MTVB88]. 1988/Green [MTVB88]. 1988/Grünes [MTVB88]. 19th [DS93]. 1st [Bri61].

2 [Sap14]. 2002 [Bar05, vO02]. 2003 [Lem03]. 2004 [Ano04]. 2006 [Fre06]. 2008 [Lei09b]. 2010 [KLR13, Lei09a]. 2012 [Ano13]. 2014 [Zam14]. 20th [Bri61, DS93, FPLR59a, Sim02, Rei07]. 21 [Ano69]. 27 [Har43]. 28 [Har43].

28.9.43 [Har43]. 29.50 [Bri61].

3 [Bal12, Gru04, Her06, Swe06, Her06]. 3-7776-1380-0 [Her06, Her06]. 3-932513-46-0 [Gru04]. 34.95$ [KK97a]. 37 [Kra81c]. 38 [Hah40].

4 [Bre36]. 4.50 [Dan36]. 46 [Kra81c, Mei42b, Mei45b].

5 [Gav04, Hes96, KK97a, Rou96]. 50 [Dan36]. 50th [Mei52b]. 526pp [Wal96].

60th [Mei39b, MWF55, ML56]. 63 [Mei43b].

7 [Kra02]. 70th [HFK58, Mei49]. 75 [Bas62]. 75th [Bas62, Str54].

8 [Gav05, Mei49]. 8/Richardson [KZN^{+}88]. 80th [Bri61, FPLR59b, HFK58, Lin60, Mei59, Mei79]. 85th [Mei64d].

91 [vG32, HM32]. 93 [Nod34]. 95 [Hof97]. 978 [Bal12, Sap14]. 978-0-19-976641-3 [Bal12]. 978-1-4516-6043-2 [Sap14].

=beam [Mei28c].

A. [HM32]. Abb [Ras81, Sch60]. Abbildungen [Mei29f, Mei32b]. Abfolge [Kei14]. able [Ker91]. above [HM32, Mei22a]. Abscheidung [HM19d]. Abschließende [Kra81c]. Absolute [For80, MO30]. Absorbance [MH31b]. Absorption [HM08c, Mei06a, Mei28b, HM09d, Mei15b]. Absorptions [Mei28b]. Absorptions-spektrum [Mei28b]. Absorptionsgesetz [HM09d, Mei15b, MH31b]. Abspaltung [MSH38]. abstract [Cun13]. Abstracts [Ano12]. Abzweigungsverhältnis [HM22]. Academic [Mei29f, Ano69]. Academy [Bar05, LMOS15]. Accident [Pop16]. Accidental [PC02]. according [Mei28e]. Acoustics [Ano79a]. act [Stu13]. Actinium [HM19c, Mei18b, HM08a, HM08b, HM13b, HM18, HM19a, HM22, Mei25, HM08a]. Actiniumreihe [Mei25, HM22]. Actiniums [HM13b, HM18, Mei18b, HM08a, HM08b, HM19c, HM19a]. Action [Whi00]. Active [HS89, Mei12, vBHM11c, HM12b]. activity [Ren04]. Adolf [Kac96, Lip59]. Advancement [Neu21]. advances [MTVB88]. Affirmative [Whi00]. After [Den99, Kae48, Mei25]. Against [Kri99, Ano46b]. Age [Rif06, Sap14, Acz09, Hah79, Her79, Her86, Nel14, Rif99, Ban04, Abb99, Gol01, McL01, Ste00, Wes01, Rad04]. ago [Bar99, Hah67, Sim14b]. Ajzenberg [Whi00]. Ajzenberg-Selove [Whi00]. Akademische [Mei29f, Ano69]. aktiven [vBHM11c, HM12b, Mei12]. Aktuelles [DRF90a, DRF90b]. Akustik [Ano79a]. Albert [Ano79b, Ano79a, B^{+}79, For80, Lam79, See56, Ano79a, B^{+}79, EHM^{+}79, For80, Her79, Lam79, See56]. Alexander [WHFB53]. alkaline [HS39]. allemand [Rad04]. allgemeine [MD35e]. Allocation [HM23]. almost [Her06]. alt [Gre84]. Aluminium [Mei34b]. aluminum [Mei34b]. always [Hen16, Seg93]. am [Ano69, Bre36, Har43, Lam79, Mei26e, Mei36a, Mei59]. Ambrosius [Mei32b]. America [Whi00, Stu85]. American [Mei28a]. Amerika [Ano46a]. Amerika/Nobel [Ano46a]. Amerika/Nobel-Preis [Ano46a]. Amerikanischer [Ano46a]. Amherst [Bar05]. among [Rom88]. Analogien [HM10]. Analogies [HM10]. Analyses [Rad99, Rad04]. Anecdotes [PC02]. Anne [Nig03]. Annette [Wol10]. anniversary [Ano46b, Bas62, Mei52b, Neu21]. Annotated [Lem10]. Announced [Ano39]. Anonymous [Far05, Swe06]. Anregung [MP33b]. Anteil [Mei56, Mei68]. Anton [Fis10, Fis12, Fis10, Fis12]. Anwendung [HM20b, Mei52a]. Apfel [Büh98b, SRRM^{+}97]. appearance [MSH38]. apple [Büh98b]. application [HM20b, Mei52a]. April [B^{+}79]. Arbeit [HM25a, Mei22a]. Arbeiten [HM21a, Hah40, Mei26d]. Arbeitet [Ano46a, Ano46b]. Arbeitet/Präsident [Ano46a, Ano46b]. Arbeitsgemeinschaft [Ano79a]. Archibald [Rad04]. Archives [Bar05, Cra02, For80]. area [KZN^{+}88]. arrangements [HJ43]. Arten [Mei21c]. Arthur [For80, Cap57a, Cap57b]. Artificial [GLR12, HM35a, Sim07b, Dan36, HM25a, HMS35, HM35b, HM35c, HM37, HM38, Mei34g, Mei34e, MD35c, HM36, Hah43, MD35i, MD35j, Mei36b, MSH38]. Association [Ano79a]. Astronautenzentrum [DRF90a, DRF90b]. Astronomer [PC02]. Astronomisches [Ano46a]. Atoll [Ano46a]. Atom [Ano39, Kae45, Lau51, Mei46a, Mei57, Sch57, Gol87, Hah43, Mei17, Mei57]. Atomarten [HM36]. Atombombe [Büh98a, Pop16, Mei56, Mei68]. Atombomben [Ano46a, Ano46b]. Atombomben/Nobel [Ano46a, Ano46b]. Atombomben/Nobel-Preis [Ano46a, Ano46b]. Atome [BFG^{+}33, Der07, DRF90a, DRF90b, Gol87, Mei24b]. Atome/Europäisches [DRF90a, DRF90b]. Atomenergie [HM54]. Atomforscher [Hof23]. Atomforschers [Hof78]. Atomforschung [Ano46a, Mei32a]. Atomforschung/Schiffssteuerung [Ano46a]. Atomforschung/Wieder [Ano46a]. Atomgewichtes [Mei31a]. Atomic [Ano16, FR13c, Gol90, Lau51, Lin60, Mei34d, MD35k, Pop21, Rei84, Sap14, SM08, WM15, WHF^{+}96, Büh98a, Cra69, Dan36, FPLR59b, Hah79, Her79, Her86, Joh17, Kei14, Ker91, Mei21a, Mei21b, Mei24d, Mei25, Mei26d, Mei27b, Mei31a, MP32, MD35c, Mei50b, Mei56, Mei68, Nel14, Pop16, Ree15a, Ban04, HM54, Mei26e, Mei34a, PMFB33]. Atominnern [Mei27b]. Atomkern [Mei34a]. Atomkerne [MD35c, Dan36, Mei26d, MD35k, PMFB33]. Atomkernen [MP32]. Atomkernprozesse [Mei34d]. Atomkernprozessen [Mei52a]. Atomkerns [Mei50b]. Atomkernspaltung [Kei14]. Atomkonstitution [Mei21a, Mei21b]. Atomkontrolle [Ano46a]. Atomkontrolle/Ehrung [Ano46a]. Atomphysikerin [Ker91, Ker98, KZN^{+}88]. Atoms [BFG^{+}33, Der07, Mei24b]. Atomumwandlungen [Hah43]. Atomvorgänge [Mei26e]. Atomzeitalter [Her86]. Atomzeitalters [Hah79]. Atomzerfall [Mei25, Mei24d]. Attempt [Mei45a, Fri02]. attempts [Nav06]. Attention [Ken54]. attribution [Alp89]. Audio [Gru04]. Audio-CD [Gru04]. Aufbau [Dan36, Mei26d, Mei27b, MD35c]. Aufbruch [Lah91]. Aufl [Bri61]. Auflage [Mei27a]. Aufnahme [Rec91a, Rec91b]. Aufnahmen [vLM27]. Auftreten [MSH38]. auftretenden [MP33c]. Auger [Bre36]. Ausbeute [HJ43]. Ausblick [Zim79]. Ausgrenzung [uRLS15]. aussendende [HM20b]. Ausstellungskatalog [Lem03]. Ausstoßung [HM09a, RM09]. Austria [DRRMH^{+}09]. Austrian [McL99]. Auswanderung [Sim14a]. author [MWZG61]. Autobiographical [For80]. autobiography [Seg93]. Autors [MWZG61]. Autors/Weitere [MWZG61]. Award [Ano67, Ano46b, Ano66]. awarded [Ano79a, Lei09b].

B [Ano46a, Mei27a, Mei29f, Mei32b]. Back [Mei64c, Mei64a, Mei64b, Vog06, Wol10]. background [Ree15b]. Bad [Ano46a, Ano46b, Ano46b]. Bande [Lem10]. Banesh [For80]. Barium [HS89]. Barth [Mei32b]. Baruch [Ano46a]. Basel [Ras81]. Battle [Kae45]. Bau [MTVB88]. Bayrischer [Ano46a]. Bd [Mei29f]. Be [Lev97, Ker91]. Beach [Ras81]. Beam [Mei22c, Mei43a, Mei28c]. beams [Mei28c]. became [Ree15a]. become [Ker91]. becomes [Her86]. bedeutenden [Kra78b]. Bedeutung [Mei23c, Mei31a, Mei32a]. Before [Whi00]. Began [FW67, Kae48, Hah67]. begann [Hah67]. Begründer [Hah79]. Behavior [MF26b, HMS37, Mei41c]. behaviors [HMS34]. Behaviour [Mei41a]. Behind [Bar05, Kra02, Fri01]. Beim [Ano46a, HM10, HMS35, Mei24d, MF26b, Mei36b]. being [Mei58]. Beispiel [SK80a, SK80b]. Beiträge [Bri61, FPLR59a, Lip59, Sch60]. Beiträgen [Mei05]. belated [Pea15]. Bemerkungen [HMS35, Kra81c, Mei26b, Mei48, HM25a, HM32]. ber [HMS35, Mei26d]. Berechnung [vLM27, HJ43]. Bergakademie [DRF90a, DRF90b]. Bericht [Ano46a, DRF90a, DRF90b, Hah40, Har43, Mei10]. Berichtigung [MF26a]. Berkeley [Hof97, KK97a]. Berlin [Ano79a, B^{+}79, Dan36, DRF90a, DRF90b, For80, Hah67, Har43, KLR13, Lem03, Wol10, Ano69, Ano79a, Ano97, Ano10, Bas62, B^{+}79, HSMH36, Hah67, Har43, Kra83, Lam79, LMOS15, Lem03, Mei54, Sim03, Vog06]. Berlin-Dahlem [HSMH36, Mei54]. Berlin/CERN [DRF90a, DRF90b]. Berlin/Hubert [Ano79a]. Berlin/Hubert-Schardin-Medaille [Ano79a]. Berliner [Ano46a, Bey09, DRF90a, DRF90b, Vog06, Wol10]. Berlins [LMOS15]. Bernardini [Bre36]. berücksichtigt [CSW97a]. Besprechungen [Har43, BBT^{+}22, BEE^{+}28, BvLM23, GOB^{+}32, HMB20, MG27, MnBB28, MJG^{+}31, MhKG37, RMJ27, SM26, ZSM32, vLvRM^{+}23]. Bestätigung [Mei24e]. bestimmter [Mei29d]. Bestimmung [Mei26c, MO30]. Bestrahlung [HMS36, HS39, MH36, MSH38]. Beta [Jen00, Mei10, Wat83]. beta-ray [Wat83]. Beta-Strahlen [Mei10]. between [Lam79, Mei22b, Mei24a, Mei31c, MP32]. Bevollmächtigte [Mai07]. bewegten [Ano12]. Beyond [Cal17a, HMS36]. Beziehung [Mei29c]. Bh [NFG^{+}08]. Bias [Hoo03, Ore03]. Big [Ano39, JMH^{+}05]. Bikini [Ano46a]. Bikini-Atoll [Ano46a]. Bilder [Hof97]. Bildern [Hah79]. Bildung [HM35a].

Bildungsfragen [MWZG61]. Biografie [Hof78]. biografischer [Der07]. Biographical [Bar05, Der07, Kur02]. Biographie [Hah79, Ras81, Kra88]. Biographies [McL99]. Biography [Hof78, Hah79]. Biologists [Dei96, All96]. Birth [Kra78a]. Birthday [Lem03, Str54, B^{+}79, FPLR59b, HM31a, HFK58, Mei39b, MWF55, Mei59, Mei64d, Mei79, Mei49, Bri61, Lin60]. Birthdays [EHM^{+}79, Ano79b, Ano79a, Lam79, ML56, FKEE48, ST68, WHFB53]. bis [Büh98b, Ern92, Ern93, Fis10, Fis12, For80, Mor98, Sim94a, SRRM^{+}97]. bisher [HM35a]. Bissigkeiten [MWZG61]. Bl [Kra81c]. Bl. [Kra81c]. Blättern [Lah91]. Bleiarten [HM33]. Board [Lei09b]. bodies [Mei06d]. Bohr [Ano71, KZN^{+}88, Ben00, SM08]. Bomb [Kae45, Ken54, Rei84, WM15, WHF^{+}96, Büh98a, Joh17, Mei56, Mei68, Pop16, Ree15a]. Bombardment [Fri39, MF39b, Mei41a]. Bombs [Pop21, Ano46b]. Bond [Lem10]. Bonnet [MWZG61]. Bonzenfreies [MWZG61]. Book [Abb99, AS96, All96, Bal12, Bey96, Bey09, For80, Gav04, Gav05, Gis97, Gol96, Gol01, Hes96, Hie00, Joh96, KK97b, KK97a, Kel96, Kra02, Kri99, Lei68, Lin60, Lip59, McL01, Mon98, Nig03, Nye97, Osg66, Rec02, Rou96, Sap14, Sim94a, Ste00, Swe06, Wal96, Whi00, Gri91, Bri61, Dan36, Mei28a, SRRM^{+}97, Wol10, WSB^{+}99]. Books [Bar05, For80, Swe06]. Bothe [Bre36]. boundaries [Rei07]. branching [HM22]. Braunschweig [Bri61, Sch60]. breakthroughs [Lig05]. Bretscher [Bre36]. bridging [Rei07]. Brief [Ano66, KZN^{+}88]. Briefe [Ern92, Ern93, Hen16, Hof23, Sim94a]. Briefwechsel [Hie00, Lem98, Lem10, WSB^{+}99]. Bright [See56]. Bringing [Stu85]. Britain [Ber16d]. brochure [Mei32b, Ano79a]. Broglie [For80]. brosch [Her06, Mei32b]. Broschüre [Ano79a]. Brown [KZN^{+}88]. Bruxelles [Far01]. Buch [Gri91]. Bührke [SRRM^{+}97]. Bürger [WSB^{+}99]. Büsten [Rec91a, Rec91b]. busts [Rec91a, Rec91b]. Butterflies [Ano46b].

C [Bal12, For80, HM08a, MF26b, Mei28a, MF26b, HM08a, HM09c]. Calculation [HJ43, vLM27]. California [Hof97, KK97b, KK97a, Rou96, Wal96, Rad99]. Call [Fre07, Lei09a]. Cambridge [Bar05, Gav05]. Cancer [Kae39]. Candidates [JKS15]. Cannonball [Ano39]. Capture [Mei40, Mei41b, Mei48]. career [Sch91]. Caricature [NCSW98]. Carl [KF01]. Cartan [For80]. Case [Ken54]. Catalog [Lem03, Mei28a, For80]. Cavendish [Osg66, Nav06]. CD [Gru04]. Celebrating [Lam79]. Celebration [Ano79a, EHM^{+}79]. Census [Bar05, CHU87]. Centenary [For80, Kra78a, Kra78b]. Center [For80, SKDA15]. centrifuge [HJ43]. Centuries [DS93]. Century [Bar05, Bri61, Den99, Dar04, FPLR59a, Lig05, Rei07, Sim02]. CERN [DRF90a, DRF90b]. certain [Mei29d]. Chadwick [Osg66]. chain [And73]. chains [SN17]. change [Ano79a]. changed [Ber11, Con16]. changer [Ree15a]. Characteristics [And98]. cheerfulness [Her06]. Chemical [Mei26e, Rei07, HM19d, HM19b, HMS34, HMS37, Mei28a, Mei36c]. Chemie [Bri61, Föl91a, FPLR59a, HSMH36, Hah40, Lip59, Mei54, Ras81, Sch60, Sim07a]. Chemie-Nobelpreis [Föl91a]. Chemiker [Ano46a, Ano46b].

Chemiker/Das [Ano46a, Ano46b]. chemische [Mei36c]. Chemischen [Mei26e, HM19d, HMS34]. chemisches [HM19b, HMS37]. chemist [Ano46b]. Chemistry [AS11, Bri61, HSMH36, Lei68, Mei54, CHU87, FPLR59a, Bar05, Hah40, SKDA15, Sim05, Sim07a, Sto97]. Choices [Whi00]. city [Ren04]. Clara [Hen16]. Clay [Bre36]. Cleavage [Mei50b]. Cloud [MF24, Mei28e]. Cockcroft [Bre36]. codiscoverer [CSW97b]. Collaboration [SKDA15, Hah67]. colleagues [Vog06, Vog14, Wol10]. Collected [KZN^{+}88, Osg66]. Collections [For80]. Colloquium [MWZG61]. Colloquium/an [MWZG61]. Cologne [Gru04]. Columbia [Ano39, Lau51]. commemoration [Hei69]. Commemorative [B^{+}79]. commentary [Ern92]. comments [Mei26b, HM25a]. communication [HM19d, HM21b, HM32, HM35c]. Community [Mai07]. Comp [Mei28a]. companions [Mei05]. Company [Mei29f]. Comparison [NFG^{+}08, NGD^{+}09]. compiled [For80]. complementary [NFG^{+}08, NGD^{+}09]. complete [Hof23]. Complex [HM09c, HM13b]. Compton [Mei24e]. Comptoneffekt [Mei24e]. Computational [SKDA15]. Computer [DRF90a, DRF90b]. Concerning [HS39]. Concluding [Kra81c]. conclusion [Mei24e]. Condensed [KZN^{+}88]. conditions [Kra83]. conduction [Mei06d]. Conference [Ano03, KLR13, Far01]. confirmation [Mei24e]. Conflicts [Hof05]. Consensus [Jen00]. consideration [CSW97a]. Constants [Ano31]. constitution [Mei21b, Mei21a]. Construction [MTVB88]. contemporary [Hof23]. content [HM21b, HM22, HM20a]. context [Ren04]. Continuous [Mei23d]. contributions [Jon90, Mei05, FPLR59a, Bri61]. Controversy [Jen00, Wat83]. conversation [Lah91]. conversion [HM25a, HMS35, HMS36, Mei34f]. conversions [HM09a, Mei28c, RM09]. Copenhagen [Ano71, Seg07]. copper [Mei46b]. corpuscular [Mei31b]. Correction [MF26a]. Correlation [Mei22c]. Correspondence [Ern92, Lem10, Lem98]. cosmic [Mei23c, Mei29c, Mei29c]. Cosmology [Rei72]. Countered [NCSW98]. Craig [Sap14]. Crawford [Bar05]. created [Acz09]. creating [Ren04]. creation [Stu13]. Credit [Joh96]. Cropper [Gav04, Rec02]. Cross [Mei40, Mei48]. Cross-Sections [Mei40]. Cultural [B^{+}79, Lem03]. Curie [Mei34g, Cal17a, Con16, LR10, Mei34g]. Curie/Ferdinand [ML56]. Curies [Kae48].

D [Bre36, HM09b, Mei15b, vBHM11b, HM09b, Mei15b]. DAGA [Ano79a]. dagegen [CSW97a]. Dahlem [Hen16, HSMH36, Hen16, Mei54]. Dampf [Mei28e]. dans [Dem15]. daraus [Mei29e]. Dardo [Bar05]. Dargebracht [Neu21]. Dark [See56]. Darstellung [For80, Mor11]. daß [Mei25]. Date [Ano79a]. Davies [KZN^{+}88]. Dawn [Abb99, Gol01, McL01, Rif06, Ste00, Wes01, Rad04, Rif99]. Day [Lau51]. Days [WM15, MLT01]. DDR [Bas62, Bas62]. DDR/Physical [Bas62]. Dead [Fri50]. Death [Ano46b]. Debever [For80]. debut [Dur03]. Decay [HM25b, Jen00, Mei12, Mei45b, vBHM13, HM09b, Mah22, Mei09a, Mei09b, Mei11b, Mei21c, Mei24c, Mei24d, MD35a, MD35b, SN17]. December [Lem03]. Declaration [KF01]. decomposition [Mei25]. dedicated [FPLR59b, Lin60]. Deerfield [Ras81]. defeat [Ols63]. Defeats [Kri99]. Defining [Mla98, Mon98]. Deichmann [All96]. Delbrück [Dan36]. Deposition [HM19d]. design [Büh98a]. detect [Nav06]. determination [Mei26c, MO30]. deuterons [Mei46b]. Deutsche [Ano46a, Ano79a]. Deutschen [Ano46a, Rec91a, Rec91b, Hof23, Wei90]. Deutsches [DRF90a, DRF90b, Sim10]. Deutschland [Ano46a, Sch91]. Deutschland/Amerikanischer [Ano46a]. Deutschland/Max [Ano46a]. Deutung [Mei21c, Mei23d, HM24, Mei23b, Mei23a]. Develop [WHF^{+}96]. development [Mei24c, Mei52b]. devotion [RCRC97]. Dezember [Lem03]. DFG [DRF90a, DRF90b]. Did [Pop21, Büh98a, CSW97b, Ols63]. Dies [Fri68]. dieser [Mei43a]. Dietrich [Her06, Kra81c, Kra81c]. Differences [Mei45a]. Different [MF26b, HJ43, Mei21c, Mei34e]. direction [Osg66]. director [Lah91, Ano56]. discovered [Mei34g]. discoverer [Bar00]. Discoveries [Laz82, Cal17a, Lig05, McG93b, McG98b]. discovers [Ell08, Trö09]. Discovery [FW67, Gar63a, Gra64, GLR12, HM32, Hoo02, Mei62, Mei63a, Mei63b, Mor98, Sim91a, Sim91b, Sim98a, Sim98b, Sim00, Sim07b, SK80b, Ano46b, FW85, GA71, vG32, Kae48, Kei14, Kra83, Pea15, Per97, Sim89, Sim14b, SP15, SK80a]. discussions [RS80]. Disintegration [MF39a, Mei45c, MF68, MF69]. dispersion [Mei07, Mei26d]. Display [Sim10]. Disputes [Ken54]. Distinguished [Fri73]. distributing [MH13]. distribution [HM12b, Mei29e]. Division [Fri39]. DM [Bri61, Ras81, Sch60]. Do [FH04]. doctoral [Mei52b]. Documentary [For80, GA71]. documents [CSW97b, Hah79, Hof23]. Doktor [Mei52b]. Doktor-Jubiläum [Mei52b]. Doktorarbeit [WSB^{+}99]. Dokumente [For80, Hof23]. Dokumenten [Hah79]. Doubtful [Kae39]. d’ouvrages [Rad99, Rad04]. Dr [Ano46a, Ano46b, Hen16]. Dr. [Ano46b, Hen16, Ken54]. Drama [Ken54, Fri02]. Dramatic [Sap14, Nel14]. Drawing [NCSW98]. drei [Ell08]. Drop [Ano94, Stu94]. Drucktechnische [MWZG61]. Dukas [For80]. dunkle [See56]. Dunlap [All96]. Dunning [Bre36]. Durable [HMS38]. Durch [Ano46a, Ano49a, Ano49b, HM35a, HM35b, HM35c, MF26b, Mei31b, MP33b, Mei34f, MHS37, MSH38, Mei43a, SK80a, SK80b]. Durchgang [MF26b]. during [MP33c]. DVT [DRF90b]. DVT-Preis [DRF90a].

Each [Ano39, MH13]. Early [Nav06, SK80b, LR10, SK80a]. earth [HS39]. earths [Mei41a, Mei41c]. Eckener [Ano46a, Ano46b, Ano46b]. ed [Bri61, Rad04]. Edited [Bre36, Her06, Lin60, Mei29f, For80]. Edith [Hen16, Hen16]. editing [Ern92]. Edition [Ern92, Mei28a, Sim94a, Mei27a]. Editor [Bar05, KK97b]. Editorial [Rom88]. education [MWZG61]. Effect [BM84, Fri17, HJ43, Mei24e, MW33]. effects [Mei28a]. efficiency [HJ43]. Egon [Mei27a]. ehnjährigen [Neu21]. Ehren [Ano69]. Ehrensaal [Rec91a, Rec91b]. ehrt [DRF90a, DRF90b]. Ehrung [Ano46a, Rec91a, Rec91b]. Eidgenössischen [Bre36, Mei36a]. Eigenschaften [HM19d, HM20c, HM21b, Mei11a]. eigentlichen [HM09e]. ein/Gesamtdeutsche [DRF90a, DRF90b]. Einfache [Mei43a, Mei11b]. Einfangquerschnitten [Mei48]. Einfluß [Mei43a]. Einführung [Mor98]. eingeweiht [DRF90a, DRF90b]. eingeweiht/Studienförderpreis [DRF90a, DRF90b]. einige [Mei06b, Mei06c, Mei11b, HMS35, Mei26b, Mei48, Mei54]. einiger [HM08c, Mei41c]. Einleitung [MD35g, MD35h]. Einstein [Ano79b, Ano79a, B^{+}79, Chi08, For80, Lam79, Ano79a, B^{+}79, EHM^{+}79, For80, Her79, JMH^{+}05, Lam79, RS80, See56, For80]. Einstein-Marić [Chi08]. Einstein-Symposion [Ano79a]. Einstein. [For80]. einzelnen [HM12b, HJ43, MH13]. EIR [MTVB88]. electromagnetic [Mah22]. Electron [Ano39, MP33c]. Electrons [Mei43a, Mei43b, Mei43c, MP33b, Mei34b, Mei43d, BFG^{+}33, Mei28a]. Elektronen [MP33b, Mei34b, Mei43a, BFG^{+}33]. Elektronenbahnen [MP33c]. Elektronenmikroskopie [DRF90a, DRF90b]. Element [HM18, HM32, Mei18b, Nod34, Sto97, vG32, HM19b, HM19b]. Elemente [HMS36, Mei34f, Mei26b, Mei34a]. Elementen [MH32, Mei34e, Mei48]. Elements [Ano92, vG32, HM32, HMS36, LT13, Lev97, Sim00, Sim07b, Kan12, Mei26b, MH32, Mei34a, Mei34e, Mei34f, Mei36c, Mei48, Sce13]. Elie [For80]. Elisabeth [Bar05, Lem10, Sch97, Bar05]. Emigration [Sim14a]. Emilio [Seg93]. Emission [Mei28d, Mei28d]. emitted [Mei25, Mei42b, Mei42a]. emittiert [Mei25]. emittierten [Mei42b]. emitting [HM20b]. employees [Mei05]. Encyclopedia [DS93, Kur02, Bar05]. Energie [Mei28c, MO30, Mei31c]. Energieentwicklung [Mei24c]. Energien [LMOS15]. Energies [LMOS15]. Energiespektrum [Mei34b]. Energieverteilung [Mei29e]. Energy [Ano39, Mei41b, Mei45a, Mei62, Mei63a, Mei63b, Seg85, HM54, Mei24c, Mei28c, MO30, Mei31c, Mei34b, Wat83, Mei29e]. England [Hof23, Hof23]. enlarged [Mei28a]. Enrico [Ano03, Hah62]. Entdecker [Ell08, Trö09]. entdeckte [Mei34g]. Entdeckung [Ano46a, vG32, HM31b, HM32, Kei14, Mor98, Per97, SK80a, SK80b, Sim91b]. Entering [Den99]. entrance [Vog06, Wol10]. entstehenden [HS39]. Entstehung [Mei22d]. Entstehungsgeschichte [HM33]. Entwicklung [Mei52b]. entworfen. [Büh98a]. Epic [Sap14, Nel14]. episode [Gri91]. Epoch [Ano39]. eponymous [KZN^{+}88]. EPS [Ano13, HJP17]. EPS-NPD [Ano13]. Epsilon [Hof23, Hof23]. era [Nel14, Sap14]. Erdalkalimetalle [HS39]. Erden [Mei41c]. Erfahrung [Mei22a]. erfolgtem [Mei25]. Erforschung [Fis10, Fis12, Str54]. ergänzt [Hof23]. ergeben [Mei06b, Mei06c]. Ergebnisse [HM12a, Ano24a, Ano24b, KZN^{+}88]. erhielt [CSW97a]. Erho [HJ43]. erhöht [MTVB88]. erhöht/Océ [MTVB88]. erhöht/Océ-van [MTVB88]. erinnert [For80]. Erinnerungen [HFK58, Her06, Mei54, Mei05, HR01]. Erkenntnis [Lam79]. Erkenntnisse [Kei14, Hah75]. Erklärung [KF01]. Erlebnisse [Hah75]. Ernest [Mac11].

Ernst [HFK58, ST68, FS06, HFK58, Sim94a]. Eröffnet [Ano46a]. Eröffnet/88 [Ano46a]. Errata [Ano94]. Erratum [Mei43b]. erregten [Mei43a]. Erregung [Mei34e]. erreicht [DRF90a, DRF90b]. erst [Mei25]. Erste [Ano46a]. ersten [MWF55]. erstmals [Hof23]. erweitert [MTVB88]. erweitert/Geheimer [MTVB88]. Erwin [McL99]. erzählt [FS06]. erzeugt [MHS37]. escape [Sim90]. Essay [Bar05]. essays [FPLR59b, Lin60]. est [Dem15]. establishment [Ols63]. Euro [Gru04, Her06]. Europa [DRF90a, DRF90b]. Europa/Risikostudie [DRF90a, DRF90b]. Europäisches [DRF90a, DRF90b]. European [AS11, Fre07, Lei09b]. Everything [KZN^{+}88]. Evidence [Fri39, MF24, HS39, Mei25, vBHM11b, HM09c]. Exact [Ano24a, Ano24b]. Exakten [Ano24a, Ano24b]. Examining [MH30a]. Example [SK80b, SK80a]. Excellence [Bar05, Kra02, Fri01]. Exception [Sim05]. Exceptional [Sim05]. Excerpt [SH02a]. excitation [MP33b, MP33c]. Excited [Mei43b, Mei43c]. excitement [Mei34e]. exclusion [uRLS15]. Exhibit [For80]. Exhibition [Lem03, B^{+}79, For80]. exhibitions [For80]. Exil [Sim14a, Vog14]. Exile [Sim94b, Sim14a, Vog14]. Existence [HS89, HS39]. expand [Mei27a]. expanded [MTVB88]. expanded/Secret [MTVB88]. Expelled [MP01]. experience [Mei22a]. Experiences [Hah75]. Experiment [Ano46a]. Experimental [Mei29f, KZN^{+}88, Mei26c]. Experimentalists [Gea14]. Experimentalphysik [Mei29f]. Experimentelle [Mei26c]. Experiments [Gea14, Mei28e, MP34]. Exposure [HM09a, RM09]. external [Kra83].

F [Bri61, Lin60, Mei29f, Mei32b, Mei34g, MH67, Ras81, Sch60]. F. [Mei34g, MH67]. Fachausstellung [DRF90a, DRF90b]. Fachliteratur [Mor98]. facility [Wei90]. facsimiles [Bey49]. Fact [Pei86]. Failed [WHF^{+}96, Ber16a]. Faksimiles [Ras81]. Fall [Sap14, Nel14]. Fallout [Sim91a, Sim91b, Sim91b]. False [Alp89]. Family [Kel96]. Fara [Swe06]. Farben-Fernsehen [Ano46a]. Farm [Hof23, Hof23]. Farm-Hall-Protokolle [Hof23]. fast [Her06, Mei31b]. faster [Mei48]. Faszination [WSB^{+}99]. Faust [Seg07]. Fay [Whi00]. Fear [WM15]. Februar [Ano69]. February [Ano69, MF69]. Federal [Bre36, Mei36a]. fehlenden [HM35a]. Feier [Ano79a, Lam79, EHM^{+}79]. Feinstruktur [MP33a, MP31]. Female [Whi00]. Feministinnen [Sch91]. feminists [Sch91]. Ferdinand [ML56]. Fermi [Ano03, Ano66, Hah62, Sim03]. Fernsehen [Ano46a]. Fernsehserie [Gri91]. Festschrift [Neu21]. field [Mei11c]. Fifty [Kae48]. fin [Ren04]. fin-de-siècle [Ren04]. Final [Sto97]. Finally [Sto97]. findings [Kei14]. fine [MP31, MP33a]. Finnland [DRF90a, DRF90b]. Fired [Mei43a]. firm [Ano46b]. First [Ken54, Sim96b, WM15, Hof23, MWF55, HM19d]. Fischer [FS06]. Fission [FW67, Gar63a, Gra64, Hah62, HS89, Mei39a, MF39c, Mei45a, Mei50a, Mor98, Ree05, Sim91a, Sim91b, Sim96b, Sim98a, Sim98b, Sim00, Sim03, Sim07b, Sim10, Sim12, SK80b, Stu85, Stu94, Alp89, Bar00, Ber16c, CSW97b, Ell08, Fri74b, FW85, GA71, Ham02, Kei14, Kra83, KF01, Laz82, LPS^{+}15, MF39b, MLT00, MLT01, Pea15, Pei86, Per97, Sim89, Sim07a, Sim14b, SP15, SK80a, Stu13, Trö09, Whe09, Zim79, Ano94, CSW97a]. flash [Gar63b]. Flaute [WSB^{+}99]. Flaute/Lemmerich [WSB^{+}99]. Fleischmann [Bre36]. Florida [Ras81]. Florida-Basel [Ras81]. Flügge [Sch98]. Focuses [Ken54]. foiled [Wat97]. Folge [Gri91]. folgende [Mei29e]. Folgerung [Mei24e]. Folgerungen [Mei06b, Mei06c]. Force [Ano39]. Förderung [Neu21]. forge [Fri02]. Forgetting [Sim12]. forgotten [Brü00]. form [MTVB88]. Formation [HM35a, HS89]. Formula [MH30b]. formulas [Mei06b, Mei06c, MH30a]. Forschung [Ano46a, Ano46b, Mai07, HM12a, Hof05]. Forschung/Goethe [Ano46a, Ano46b]. Forschung/Goethe-Preis [Ano46a, Ano46b]. Forschungsreaktor [DRF90a, DRF90b]. Forschungsunternehmen [Ano46a, Ano46b]. Forschungszentrum [DRF90a, DRF90b]. Fortschritt [MTVB88]. Found [Sim96b]. Foundations [Bey49, Mor98]. founder [Hah79]. four [Cal17a, Cra92, Trö09]. Frage [HM13b, HM19b]. Fragen [MWZG61]. fragment [Brü00, LPS^{+}15]. fragmentation [MD35j]. Fragments [HS89]. Franchetti [Bre36]. Franck [Gea14]. Frank [For80]. Frau [Der07, Gru04, Mei03]. Frauen [DS93, Chi94, Föl91b]. Frauenstudium [Sch91]. Frédéric [HFK58, HFK58, LR10]. Freed [Ano39]. freedom [EMvL79]. Freeman [Bar05, Kra02]. Freiberg [DRF90a, DRF90b]. Freiheit [EMvL79]. Freitod [Hen16]. French [For80, Dem15, Gol87]. Fresnel} [Mei06b, Mei06c]. Fresnelschen [Mei06b, Mei06c]. Freundschaft [Lem10, uRLS15]. Frieden [HM54]. Friedman [Bar05, Kra02]. Friedr [Sch60]. Friedrich [Bri61, For80, Lip59]. friends [Gol96]. Friendship [Lem10, uRLS15]. Friendships [Cro01b]. Frisch [Bri61, CSW97a, Lin60, Lip59, Sch60, CSW97a, Gar63a, Sim96b, Stu13]. Fritz [CSW97a, FH81, Ras81, CSW97a, Kra81b, Kra81a, Kra81c]. fröken [FH04]. frühes [SK80a, SK80b]. fundamental [Bey49]. Fundamentals [HM12a]. Further [MWZG61, MP34]. fusion [MTVB88, Rei72]. fusion/Nobel [MTVB88]. Future [Fri12, Zim79].

G [Bre36, Mei27a, Mei32b]. Galilei [Büh98b, SRRM^{+}97]. Galileo [Gav04, Gav05, Rec02, Büh98b, Cro01a, Jam04b]. Gambini [SRRM^{+}97]. game [Ree15a]. game-changer [Ree15a]. Gamow [MP31, Rei72]. gamowsche [MP31]. ganze [Sim07a]. gas [Mei28e]. Gase [MF26b]. Gases [MF26b]. Gasgemische [Mei28e]. Gauge [SRRM^{+}97]. Geb [Mei27a, Mei29f, Bri61, Hen16, KK97a, Mei32b, Sch60]. Gebiet [Mei11c]. Geburtstag [Ano79b, B^{+}79, Bri61, For80, Kra78b, Lem03, Mei39b, Mei59, HM31a, Mei49, Mei64d, Mei79, Str54, Sch60]. Geburtstage [Ano79a, EHM^{+}79, Lam79]. Gedächtnisausstellung [B^{+}79, For80]. Gedenkfeier [Ano69]. Gedenkworte [Hei69]. geeignete [Mei28e]. Gegen [Ano46a, Ano46b]. Gegendarstellung [Kra81c]. geh [Dan36]. Gehalt [HM21b, HM20a]. gehalten [Bre36, Mei26e]. Geheimer [MTVB88]. geht [Der07]. Geiger [Bre36, DRF90a, DRF90b]. Gelehrten [Ano46a].

Gelehrten/Neuer [Ano46a]. Gelehrter [Ano46a]. Gelehrter/Farben [Ano46a]. Gelehrter/Farben-Fernsehen [Ano46a]. gelingen [Ker91]. Gemeinschaftsforschung [Mai07]. Gender [Hoo03, Ore03]. General [For80, MD35e, Lee00]. genesis [HM33]. genius [Gar63b, Mac11]. Geometrical [KZN^{+}88]. geopolitical [Ree15a]. George [Bar05, Rei72]. Gerald [For80]. gerettet [Sim07a]. Gerhard [FKEE48]. German [Ano24a, Ano24b, Ano31, Ano46b, Ano49a, Ano49b, Ano56, Ano59b, Ano69, Ano79b, Ano79a, Ano10, Ano12, ABB^{+}36, vBHM11a, vBHM11b, vBHM11c, vBHM13, vBHM14, vBHM15, Bas62, BBT^{+}22, B^{+}79, BEE^{+}28, BvLM23, BFG^{+}33, Bre36, Bri61, Büh98a, Büh98b, Cap57b, Chi94, CSW97a, CSW98, DF14, Dan36, Der07, DS93, DRF90b, EHM^{+}79, EMvL79, Ell08, Ern92, Ern93, FKEE48, FS06, Fis10, Fis12, Föl91b, FBLB^{+}35, FPLR59a, Gre84, Gri91, vG32, Gru04, GOB^{+}32, HM08a, HM08c, HM08b, HM09d, HM09a, HM09c, HM09b, HM09e, HM10, HM12a, HM12b, HM13b, HM13a, HM14, HM18, HM19d, HM19b, HM19c, HM19a, HM20b, HM20c, HM21a, HM21b, HM22, HM23, HM24, HM25a, HM25b, HM31a, HM32, HM33, HMS34, HMS35, HM35a, HM35b, HM35c, HM36, HSMH36, HMS36, HMS37, HM37, HMS38, HM38, HS39, Hah40, Hah43, HM54]. German [HFK58, Hah67, Hah75, Hah79, Har43, HJ43, Hei69, Hen16, Her86, HR01, Her80, Her10, Her06, Hof78, Hof05, Hof23, HMB20, Kei14, Ker91, Ker98, Kra78b, Kra81b, Kra81a, Kra81c, KZN^{+}88, KF01, Lah91, Lam79, LMOS15, Lem98, Lem03, Lem10, Mai07, Mei06a, Mei06b, Mei06c, Mei06d, Mei07, Mei09a, Mei09b, Mei10, Mei11b, Mei11a, FM11, Mei11c, Mei12, MH13, Mei15b, Mei18a, Mei18b, Mei18c, HM20a, Mei21a, Mei21b, Mei21c, Mei22d, Mei22a, Mei22c, Mei22e, Mei22b, Mei23c, Mei23d, Mei23b, Mei23a, Mei24c, Mei24e, MF24, Mei24d, Mei24b, Mei24a, Mei25, Mei26e, MF26a, Mei26b, Mei26c, Mei26a, Mei26d, MF26b, Mei27b, MG27, Mei27a, MnBB28, Mei28e, Mei28c, Mei28b, Mei28d, Mei29e, Mei29c, Mei29f, Mei29d, Mei29b, Mei29a, MO30, MH30b, MH30a, MH31b]. German [Mei31a, MJG^{+}31, Mei31b, MH31a, MP31, Mei31c, Mei32a, Mei32b, MH32, MP32, MP33b, MW33, MP33c, MHK33, MK33, MP33a, Mei34a, Mei34g, Mei34e, Mei34b, Mei34d, Mei34c, Mei34f, MP34, MD35c, MD35e, MD35g, MD35h, MD35i, MD35j, MD35f, MD35k, MD35d, MD35a, MD35b, Mei36b, Mei36a, MH36, Mei36c, MhKG37, Mei37, MHS37, MSH38, Mei39b, Mei41c, Mei42b, Mei43a, Mei45b, Mei48, Mei49, Mei50b, Mei52a, Mei52b, Mei54, MWF55, ML56, Mei56, Mei57, Mei58, Mei59, MWZG61, Mei63b, Mei64d, Mei68, Mei79, Mei01, Mei03, Mei05, MH67, MTVB88, MLT00, MLT01, Mor98, Mor11, Neu21, Nod34, Ols63, Per97, PMFB33, Pop16, PdK23, Rec91a, Rec91b, RMJ27, Rif90, Rif92, RM09, Sch91, Sch57, See56, SH02a, SKDA15, Sim91b, Sim98b, Sim01]. German [Sim07a, Sim12, Sim14a, uRLS15, SRRM^{+}97, SM26, Str54, ST68, SK80a, SK80b, TF99, Trö09, Vog06, Vog14, Vor10, Wal03, WHFB53, Wei90, Wol10, WSB^{+}99, ZSM32, Zim79, vLM27, vLvRM^{+}23, vM55]. Germany [Lin60, Ber16a, Sch91, Sim90, WHF^{+}96]. Ges [Bas62]. Ges. [Bas62]. Gesamtdeutsche [DRF90a, DRF90b]. Geschäftsführer [Lah91].

Geschichte [KZN^{+}88, LMOS15]. Geschlechterverhältnisse [SRRM^{+}97]. Geschwindigkeit [Mei29d]. Gesellschaft [Bas62, DRF90a, DRF90b, Mai07, Mei26e, Neu21, Vog06, Wol10, Bey09]. Gesellschaft/Bergakademie [DRF90a, DRF90b]. Gesellschaft/Hahn [Bas62]. Gesetz [EMvL79]. Gespräch [Lah91]. Gift [MP01]. gleichnamigen [KZN^{+}88]. gleichzeitig [HM20b]. Glimpses [For80]. global [Ree15a]. goes [Der07]. Goethe [Ano46b]. Goethe-Preis [Ano46a, Ano46b]. Got [Sim96b, Ols63]. Göttingen [KF01]. Göttinger [KF01]. Gravity [SRRM^{+}97]. Gravity/MuPAD [SRRM^{+}97]. Great [Cro01a, Far05, Kra78a, Swe06, Büh98b, Kae48, Kra78b, Lig05, Ols63, Ber16d, Gav04, Rec02]. Green [MTVB88]. Greiner [Lei09b]. Grinten [MTVB88]. Grinten-Preis [MTVB88]. Grosse [HM32, HM32]. Großforschungseinrichtung [Wei90]. Grösste [Ano46a, Ano46b]. ground [JKS15, Ken54]. Group [Far01]. Grundlagen [Mor98, HM12a]. Grünes [MTVB88]. Guide [Ber16b]. Guido [FKEE48]. Gustav [FKEE48, WHFB53].

H [Bar05, Bre36, Gav04, Kra02, Mei29f, Rec02, Ken54]. H-Bomb [Ken54]. habe [Büh98a]. Haber [Hen16, Hen16]. Hahn [Ano46a, Ano49a, Ano49b, Ano59b, Ano69, Ano79b, Ano79a, B^{+}79, Bri61, CSW97a, Ern92, Ern93, Föl91a, For80, Gre84, Hei69, Hen16, Her80, Her10, Her06, Hof78, Hof05, Kra81c, KF01, Lah91, Lam79, Lin60, Mei39b, Mei49, Mei52b, MWF55, Mei59, Mei64d, Mei79, Mei05, MH67, MLT00, MLT01, Sch60, Sim94a, Sim07a, Sto89, SK80a, SK80b, Tru96, Wal03, Zim79, Ano49a, Ano49b, Ano59a, Ano69, Ano79a, Ano97, Ano10, Bas62, B^{+}79, CSW97a, EHM^{+}79, Ern92, Ern93, FR13a, FKEE48, FPLR59b, Gre84, Hah67, Hah79, Hei69, Hen16, Her79, Her80, Her10, Hof78, Hof05, Kra81c, KF01, Lah91, Lam79, Mei39b, Mei49, Mei52b, Mei59, Mei64d, Mei79, MH67, MLT00, Mor11, Ree05, She83, Sim03, Sim06, Sim07a, Sim12, uRLS15, SK80a]. Hahn [SK80b, Vor10, Wal03, Wal06, Wei90, Zim79, Her06]. Hahn-Meitner-Institut [Bas62, Ano97, Gre84]. Hahn-Meitner-Instituts [Lah91]. Hahn-Meitner-Straßmann [Kra81c]. Hahn/Lise [Sto89]. Hailed [Ano39]. Halban [Bre36]. Hall [Hof23, Rec91a, Rec91b]. Handbook [Mei29f]. Handbuch [Mei29f]. Hannelore [Nig03]. Hans [MWF55, MWF55]. hard [Mei34c]. Hardy [Nig03]. Hargittai [Bar05]. Harms [Mei29f]. harter [Mei34c]. Harvard [All96]. Hasse [KZN^{+}88]. Haug [SRRM^{+}97]. Hauptportal [Bey09, Vog06, Wol10]. Hawking [KZN^{+}88, Cro01a, Gav04, Rec02]. Heads [Gri91]. Heat [Mei06d]. heavier [Mei48]. Heavy [Fri39, MH32]. Heilbron [KK97b, KK97a]. Heinrich [MWF55]. Heisenberg [DRF90a, DRF90b, KF01, Rec91a, Rec91b, BM84, Cap57a, Cap57b, KF01, Rec91a, Rec91b]. Heisenberg/DVT [DRF90a, DRF90b]. Heisenberg/DVT-Preis [DRF90a]. held [Bre36, Mei26e]. Helen [For80]. Helle [See56]. Her [Kra78a, CSW97b, Der07, Kra78b, Vog06, Vog14, Wol10]. Heraeus [DRF90a, DRF90b]. Herausgeg [Sch60, Bri61]. Herausgegeben [Mei29f].

Herewith [Ols63]. Heritage [B^{+}79, Lem03]. Herneck [For80]. heroine [Kra00]. Herstellung [HM09b, HM19b]. Herstellungsmethoden [Mei11b]. Herta [For80]. Hertz [Gea14]. Hevesy [Mei32b]. High [SRRM^{+}97, JKS15]. high- [JKS15]. Hintereingang [Bey09, Vog06, Wol10]. Hintertreppe [Fis10, Fis12]. Hiroshima [Lau51]. Hirzel [Her06]. Historic [HJP17]. Historical [Cra02, NCSW98, Seg85, Ben00, Wat83, Bar05]. History [Bal12, Fri12, KK97b, KK97a, KZN^{+}88, LMOS15, Sim10, WP85, Ano46b, Fri02, Gou98, GA71, Her86, Her80, Her10, KLR13, KHFA67, L’A07, LR10, Pop16, Rad99]. Hitler [Tru96, Dei96, Kac96, MP01, Pop16, Pop21, All96]. Hitlers [Pop16]. hochauflösende [DRF90a, DRF90b]. Hochschule [Bre36, Mei36a]. Hoffman [For80]. Hoffmann [FKEE48]. Hoffmann/Guido [FKEE48]. Höhenstrahlung [Mei29c]. homogeneous [Mei26c]. homogener [Mei26c]. honor [Ano69, Rec91a, Rec91b]. House [HJP17]. HQ [KLR13]. HQ-3 [KLR13]. Hrsg [Bre36, Her06]. Huber [WSB^{+}99]. Hubert [Ano79a]. Hubert-Schardin-Medaille [Ano79a]. Human [For80, Gav05, Mei58, Mei01]. Humanities [For80]. hundertsten [Kra78b]. Hundred [Bar05]. hung [HJ43]. Hupfeld [BM84]. Hydrogen [Ken54].

I. [Mei34g]. Ich [Büh98a]. Ida [Hoo03, Jon90, Ore03]. idea [JMH^{+}05]. Ifetime [HM19c]. ihr [CSW97a, HMS37, MF26b]. ihre [Bey09, HM23, HMS34, Kra78b, Mei24e, Mei26e, Mei29c, Vog06, Vog14, Wol10]. ihrem [Mei31b, Neu21]. ihren [Der07, Mei22c, Neu21]. ihrer [Mei21c]. II [For80, HM21b, HM35c, SK80b, Ber16d, Mei06d, Sim12, SK80b]. illus [Nig03]. illustrations [Mei29f]. im [B^{+}79, Bre36, Chi94, DS93, DRF90a, DRF90b, Föl91b, HM13b, Kei14, Lah91, Mai07, Mei32b, Sim07a, Vog14, Zim79, Der07, Kra81b, Kra81a, Ras81]. Images [For80]. Immerwahr [Hen16]. implications [Mei06b, Mei06c]. importance [Mei23c, Mei31a, Mei32a]. Inauguration [Rec91a, Rec91b]. Inconvenient [Sim10]. increase [HJ43]. increased [MTVB88]. increased/Océ [MTVB88]. index [Bar05]. individual [HM12b]. Industrielandschaft [DRF90a, DRF90b]. Influence [Mei43a]. Infrared [MTVB88]. Infrarot [MTVB88]. Infrarot-Observatoriums [MTVB88]. Infrarot-Observatoriums/Neu [MTVB88]. Ing [Hen16]. inhomogenen [Mei06d]. inhomogeneous [Mei06d]. Injustice [CSW97b, CSW96]. inner [MW33]. innere [MW33]. innerer [Mei31b]. Innovation [WSB^{+}99]. insights [Hah75]. instead [Ano46b]. Institut [Ano10, Bas62, Bre36, HSMH36, Lah91, Mei36a, Mei54, MTVB88, Sim07a, Ano59a, Ano97, Gre84]. Institute [Ano56, Ano71, Ano10, Bas62, Bre36, Gre84, HSMH36, Hah40, Mei36a, Mei54, MTVB88, Sim05, Sim07a, Lah91, MTVB88]. Institute [?]StuttgartMeyer-Ter-Vehn:1988:ESP. Instituten [Neu21]. institutes [Neu21]. Instituts [Hah40, Lah91]. Intellectual [Jon90]. Intellectuals [DF14]. Intellektuelle [DF14]. interaction [MP32]. interactions [Mei31c]. interdisciplinarity [uRLS15]. Interdisciplinary [SK80b, SK80a]. interdisziplinärer [SK80a, SK80b]. Interdisziplinarität [uRLS15]. Interference [Mei43b, Mei43c, Mei43d]. interior [Mei27b]. Internal [Kra83, Mei31b]. International [Ano03, KLR13]. Internationale [DRF90a, DRF90b]. internationalism [Cra92]. interned [Hof23]. internierten [Hof23]. Interpretation [Ano94, HM24, Mei23d, Mei23b, Mei23a, Stu94, CSW97b, Mei21c, Stu13]. Interscience [Lin60]. Introduction [Cal17b, MD35g, MD35h, Mor98, L’A07]. inventory [KHFA67]. Inverse [SRRM^{+}97]. Investigation [Mei43b, Mei43c, Mei43d]. investigations [Mei36c]. Ioan [Gav05]. Ionen [BFG^{+}33, FM11]. Ionisierungswahrscheinlichkeit [Mei31b]. ionization [Mei31b]. Ionizing [Den99]. Ions [BFG^{+}33, FM11]. Irène [Con16, LR10, ML56]. Irradiated [HS89]. Irradiation [MH36, HMS36, HS39, MHS37, MSH38, Mei41c, Mei46b]. Irrwege [Mei63b]. ISBN [All96, Bal12, Gav04, Gav05, Gru04, Her06, Hes96, KK97a, Kra02, Rou96, Sap14, Swe06]. Isolation [HM32, vG32]. Isolierung [vG32, HM32]. isomerer [MSH38]. isomeric [MSH38]. Isotopes [HS89, HM33, Mei26b, SN17]. Isotopie [Mei26b]. Israeli [SKDA15]. issues [MWZG61]. István [Bar05]. IV [DRF90a, DRF90b, Far01].

J [Bre36, KK97b, KK97a]. Jaenicke [KZN^{+}88]. jahr [Ano46b]. Jahre [Ano46a, Bas62, FKEE48, For80, Gre84, HFK58, Sch91, WHFB53, HFK58, MWF55, ST68, WHFB53, ML56]. jahre/Alexander [WHFB53]. Jahre/Ernst [HFK58]. Jahre/Gustav [FKEE48]. Jahre/Lise [FKEE48, WHFB53]. Jahre/Max [FKEE48]. Jahren [Ern92, Ern93, Hah67, Sim94a, For80]. Jahrhundert [DS93]. Jahrhunderts [Sch60, Bri61, FPLR59a, Lip59]. jährigen [Mei52b]. James [Gav05, Osg66]. Japanischen [Ano46a]. Jauho [SRRM^{+}97]. Jeanne [Chi08]. jenseits [HMS36]. Jesenská [Chi08]. Jewish [DF14, DS93]. Joan [For80]. Johann [Mei32b]. John [Rad04]. Joliot [HFK58, Mei34g, ML56, HFK58, LR10, Mei34g]. Joliot-Curie [ML56]. Joliot-Curie/Ferdinand [ML56]. Joliot/Lise [HFK58]. Jost [Hie00]. journals [Bey49]. Journey [FR13c]. Jr [Bre36]. Jubiläum [Neu21, Mei52b]. jüdische [DF14, DS93]. Juli [B^{+}79, Bre36, Mei26e]. Jülich [DRF90a, DRF90b]. Julius [Dan36]. July [B^{+}79, Bre36, KLR13, Mei26e]. jun [Bre36]. June [B^{+}79, KLR13, Bre36]. Juni [B^{+}79, Bre36]. Juni-4 [Bre36]. Jürgen [DRF90b]. Jürgen-Geiger-Preis [DRF90a]. Just [Bar99].

Kaiser [Hah40, Mai07, Mei54, Neu21, Sim07a, Wol10, HSMH36, Hah40, Mai07, Mei54, Neu21, Sim05, Sim07a, Vog06, Wol10, Bey09]. Kaiser-Wilhelm-Gesellschaft [Mai07, Vog06, Wol10, Bey09]. Kaiser-Wilhelm-Institut [Mei54, Sim07a]. Kaiserslautern [B^{+}79]. Karl [WHFB53]. Karrierefrauen [Sch91]. Keeping [Bal14]. Kein [Mei56, Mei68]. Kernenergie [Mei63b]. Kerner [KZN^{+}88]. Kernforschung [Ano10, Bas62, Gre84]. Kernforschungsinstitutes [Ano56]. Kernfusion [MTVB88]. Kernfusion/Nobelpreis [MTVB88]. Kernkraftwerke [DRF90a, DRF90b]. Kernkraftwerke/Computer [DRF90a, DRF90b]. Kernmodell [Mei22a]. Kernphysik [ABB^{+}36]. Kernphysik. [Bre36]. Kernphysikalische [Mei36a]. Kernspaltung [CSW97a, MLT00, MLT01, Mor98, Sim91b, SK80a, SK80b, Zim79, Ell08, Per97, Sim98b, Trö09]. Kernstruktur [Mei21c, Mei26a]. Kernumwandlungen [Dan36, MD35c]. Key [Ken54]. Kharkiv [PC02]. Kienle [MWF55, MWF55]. Kindness [Her06]. Kinetics [SRRM^{+}97]. Klein [MH30b, MH30a, MH30b, MH30a]. kleinsten [Fis10, Fis12]. Knots [SRRM^{+}97]. knowledge [Lam79, Mai07]. Kohlrausch [Mei29f]. Kolleginen [Vog06]. Kolleginnen [Bey09, Vog14, Wol10]. Kolloquium [MWZG61]. Kolloquium/Drucktechnische [MWZG61]. Köln [DRF90a, DRF90b, Gru04]. Köln/20 [DRF90a, DRF90b]. Kommentierter [Lem10]. Kommentierung [Ern92, Sim94a]. komplexen [HM09c, HM13b]. Konflikte [Hof05]. Konstanten [Ano31]. kontinuierlichen [Mei23d]. Köpfe [Gri91]. Korpern [Mei06d]. Korpuskularstrahlen [Mei31b]. kosmische [Mei23c]. kosmischen [Mei29c]. Krafft [Ras81]. kriegsrelevanter [Mai07]. Kulturbesitz [B^{+}79, Lem03]. Kunde [Ano46a]. Kungälv [HJP17]. Künstliche [Hah43, Dan36, HM25a, HM35a, HM35b, HM35c, HM37, HM38, Mei34g, MD35c, HM36, MD35i, MD35j, Mei36b, MSH38]. künstlichen [HMS35]. künstlicher [Mei34e]. Kurian [Bar05]. kurze [KZN^{+}88]. Kurzfassungen [Ano12]. kurzlebiges [HM08a, HM09b]. Kurzporträt [Ker91]. kurzwellige [MH31b]. kurzwelliger [MH30b, MH30a, MH31a, MH32, MHK33, MK33].

L [Ano59b, Bre36, KK97b, KK97a, Mei28a, MH67, Rou96]. L. [MH67]. Laboratory [Nav06, Osg66]. Lady [BM84]. Lamla [HFK58, ST68, HFK58]. Langer [Mei18b, HM18, MF24]. langlebiges [HMS38]. langsamer [Mei48]. largest [Ano46b]. Laue [Ano79b, Ano79a, B^{+}79, Bri61, FKEE48, For80, Hie00, Lam79, Lem98, Lin60, Sch60, WSB^{+}99, Ano49a, Ano49b, Ano79a, B^{+}79, EHM^{+}79, FPLR59b, Her79, Lam79, Lem98, RS80]. Laue/Einstein [Ano79a]. Laue/Einstein-Symposion [Ano79a]. Laue/Otto [FKEE48]. Laureates [Bar05, Dar04]. Laves [Bri61, Lin60, Lip59, Sch60]. Law [MH31b, EMvL79, HM09d, Mei15b, MH31a]. Lead [Mei40, HM33]. Leaders [Lau51, Ols63]. Leading [Gav04, Rec02, Cro01a]. leap [Fis10, Fis12]. Leben [DS93, For80, Hof78, Kra81b, Kra81a, Ras81, Rif90, Rif92, Sim01]. Lebensdauer [HM19c, HM21b, Mei18a, HM20a, HM18, Mei18b]. Lebensgeschichte [Ker98, KZN^{+}88]. Lecture [Mei26e, Mar54]. Lectures [HR01, Mei36a, Bre36, Mei11c]. legend [MLT00, MLT01]. Legende [MLT00, MLT01]. Lehrbuch [Mei32b]. Leibniz [LMOS15]. Leibnizschen [LMOS15]. Leipzig [Mei27a, Mei29f, Mei32b]. Leiterin [Ano56]. Leitfaden [WSB^{+}99]. Lemmerich [WSB^{+}99, Hie00]. Leseprobe [SH02a]. Leserbriefe [TF99]. letter [Dem15]. Letters [For80, Hen16, Hof23, TF99, Ern93, Lah91]. Leuna [Har43, Har43]. levels [Mei31b]. Lewin [AS96, Bey96, Gis97, Gol96, Hof97, Joh96, Kel96, Kri99, Mor11, Nye97, Rad99, Wal96, Whi00]. Lewis [For80]. Lexikon [DS93]. Liberate [Ano39]. Library [B^{+}79]. Licht [MTVB88]. Liebenswürdigkeit [Her06]. Life [AS96, Bey96, DS93, Gav04, Gis97, Gol96, Hes96, Hof97, Joh96, KK97b, KK97a, Kel96, Kra78a, Kra81a, KZN^{+}88, Kri99, Mar54, Mei18b, Nye97, Rec02, Whi00, Cro01a, HM18, Hof78, Ker98, Kra81b, HM20a, Rad99, Rif90, Rif92, Rou96, Sim96a, Sim01, Sim02, Wal96]. Lifetime [Mei18a, HM21b]. Light [MTVB88]. Lihotzky [Chi08]. like [Gol96]. Lind [Mei28a]. Liquid [Ano94, Stu94]. Liquid-Drop [Ano94, Stu94]. Lise [Ano38, Ano46a, Ano49a, Ano49b, Ano69, Ano79b, Ano79a, Ano04, Ano13, B^{+}79, Bey09, Bri61, Büh98a, Büh98b, Cap57a, Cap57b, Chi08, CSW97a, Dan36, Der07, FKEE48, For80, HFK58, Hah67, Hei69, Hen16, Her80, Her10, Her06, Ker91, Ker98, KZN^{+}88, Lam79, Lei09a, Lem10, Lin60, MLT01, Mor11, Rec91a, Rec91b, Sch60, SH02a, Sim14a, uRLS15, SRRM^{+}97, Sto89, Str54, TF99, Vog06, Vor10, WHFB53, Wol10, Zam14, Zim79, AS96, Ano49a, Ano49b, Ano56, Ano67, Ano69, Ano79a, Ban04, Bar00, B^{+}79, Ben00, Büh98a, Büh98b, Cal17c, Cap57b, Con16, Cra69, CSW97a, CSW97b, CSW98, Cro01b, Der07, EHM^{+}79, Ern92, Ern93, FR13b, Föl91a, Fre06, Fre07, FPLR59b, Fri68, Fri70, Fri73, Fri74a, Gar63a, Gol01, Gru04, HFK58, Hah67, Ham02, Hei69]. Lise [Hen16, Her79, Her80, Her10, Jam04a, Jon90, Kac96, Kan12, Kar79, Ker91, Ker98, Kra78a, Kra78b, KZN^{+}88, Kra88, Kra00, Kri99, Lam79, Lei09b, Lem98, Lem03, Lem10, Mah22, McG93a, McG98a, McL99, MWF55, Mei64a, Mei64b, Mei05, MLT00, Mor11, Off11, Org46, PC02, Per97, Rec91a, Rec91b, Ree05, Rif90, Rif92, Rif99, Rif06, Rom88, Sch97, SH02b, SH02a, SKDA15, Sim89, Sim90, Sim91a, Sim91b, Sim94b, Sim96a, Sim98a, Sim98b, Sim01, Sim02, Sim05, Sim14a, uRLS15, SP15, Str54, ST68, Vog06, Vog14, Vor10, Wat83, Wat97, Whi00, WSB^{+}99, Zim79, vO02, Abb99, Bey96, Gis97, Gol96, Gru04, Her06, Hes96, Hie00, Hof97, Joh96, KK97b, KK97a, Kel96, Kri99, McL01, MLT01, Nig03, Nye97, Rad99, Rad04, Rou96, Sim94a, Ste00, TF99, Wal96, Wes01]. Listen [Gru04]. Literatur [FBLB^{+}35]. Literature [Mor98, FBLB^{+}35]. lived [HM08a, HM09b, JKS15]. lives [McG93b, McG98b]. London [Bas62, Bas62]. London/Max [Bas62]. Long [Mei18b, MF24, HM18, JKS15]. long-lived [JKS15]. Long-Range [MF24]. Looking [Mei64c]. looks [Mei64a, Mei64b]. Loops [SRRM^{+}97]. Lord [Osg66, HM31a]. Louis [For80]. Low [KZN^{+}88]. Ltd [Mei29f]. Lübbig [SRRM^{+}97].

M [Bre36, Mon98, Mei29f, Sch60]. M. [Mei27a]. Macroscopic [KZN^{+}88]. Magneten [DRF90a, DRF90b]. magnetic [vBHM13, vBHM14, vBHM15, vBHM11a]. magnetische [vBHM13, vBHM14, vBHM15, vBHM12, vBHM11a]. main [Vog06, Wol10]. makers [Ano99]. Making [Ano39, Gou98, LT13]. Malley [Bal12]. managing [Lah91]. Manhattan [Ree14, Ree15a]. Manne [Kac96]. manufacturing [Mei11b]. Manuscripts [For80]. Many [Tel55, Rom88].

Map [PC02]. Marc [Bar05, Kra02]. March [Ano16, B^{+}79, Cap57a, Cap57b, Mei39b, Mei49, Cap57b, Lam79, Mei59]. March/Der [Cap57a, Cap57b]. March/The [Cap57b]. Margaret [Whi00]. Margarete [Chi08]. Marić [Chi08]. Marjorie [Bal12]. März [B^{+}79, Lam79, Mei39b, Mei49, Mei59]. mass [Mei31c]. Masse [Mei31c]. Materie [HM09a, RM09]. math [You08]. Matilda [Fri17]. Matter [KZN^{+}88, Whi00, HM09a, RM09]. matters [Rei72]. Matthäus [Wei90, Wei90]. Mauro [Bar05]. Max [Ano46a, Ano46b, Ano49a, Ano49b, Ano79b, Ano79a, Bas62, B^{+}79, Bri61, Dan36, FKEE48, Fis10, Fis12, For80, Hie00, Lam79, Lem98, Lin60, MTVB88, Sch60, WSB^{+}99, Ano46b, Ano49a, Ano49b, Ano79a, Bas62, B^{+}79, EHM^{+}79, Fis10, Fis12, FPLR59b, Her79, HR01, Lam79, Mei58, Mei01]. Max-Planck-Institute [MTVB88]. Measurements [Mei43b, Mei43c, Mei43d]. Measuring [Mei43a]. Mechanism [FW67, Whe09]. Medaille [Ano49a, Ano49b, Ano79a]. medal [Ano79a, Ano49a, Ano49b]. Medicine [Ano46b, Cra02, Bar05]. Medizin [Ano46a, Ano46b, SRRM^{+}97]. Medizin/Pasteur [Ano46a, Ano46b]. Mediziner [Ren04]. Mediziner-Viertel [Ren04]. meetings [Har43]. Meets [Rei84]. Megavoltage [Rob95]. meinen [For80]. Meissner [WHFB53]. Meitner [Abb99, AS96, Ano46a, Ano49a, Ano49b, Ano56, Ano69, Ano79a, Ano97, Ano04, Ano10, Ano13, Bas62, B^{+}79, Bey96, Bey09, Bre36, Bri61, Büh98a, Cap57a, Cap57b, Chi08, CSW97a, CSW98, Dan36, Der07, DRF90a, DRF90b, Ern92, Ern93, FKEE48, Föl91a, Fre06, Fre07, Fri70, Gis97, Gol96, Gol01, Gre84, Gru04, HFK58, Hen16, Her80, Her10, Her06, Hes96, Hie00, Hof97, Joh96, Kac96, KK97b, KK97a, Kel96, Ker91, Kra78b, Kra81c, Kra88, Kri99, Lah91, Lam79, Lei09a, Lei09b, Lem98, Lem03, Lem10, Lin60, McL01, MH67, MLT00, MLT01, Mor11, Nye97, Per97, Rad99, Rad04, Rec91a, Rec91b, Rif90, Rif92, Rou96, Sch60, SKDA15, Sim91b, Sim94a, Sim98b, Sim01, Sim14a, uRLS15, Ste00, Str54, ST68, SK80a, SK80b, TF99]. Meitner [Vog06, Vog14, Vor10, Wal96, WHFB53, Wei90, Wes01, Whi00, Wol10, WSB^{+}99, Zam14, Zim79, vO02, Ano38, Ano49a, Ano49b, Ano56, Ano59a, Ano59b, Ano67, Ano69, Ano79b, Ano79a, Ban04, Bar00, B^{+}79, Ben00, BM84, Brü00, Büh98a, Büh98b, Cal17c, Cap57b, Con16, Cra69, CSW97a, CSW97b, CSW98, Cro01b, Der07, Dur03, EHM^{+}79, Ern92, Ern93, FR13a, FR13b, For80, Fri02, FH04, Fri12, Fri17, FPLR59b, Fri68, Fri73, Fri74a, Gar63a, HFK58, Hah67, Ham02, Hei69, Hen16, Her79, Her80, Her10, Jam04a, Jon90, Kan12, Kar79, Ker91, Ker98, Kra78a, Kra78b, KZN^{+}88, Kra00, Kri99, Lam79, Lem03, Mah22, McG93a, McG98a, McL99, MWF55, Mei64a, Mei64b, MH67, MLT00, NCSW98, Off11, Org46, PC02, Per97, Rec91a, Rec91b, Ree05, Rif90, Rif92, Rif99]. Meitner [Rif06, Sch97, SH02b, SH02a, Sim89, Sim90, Sim91a, Sim91b, Sim94b, Sim96a, Sim96b, Sim98a, Sim98b, Sim01, Sim02, Sim03, Sim05, Sim14a, uRLS15, SP15, SRRM^{+}97, Sto89, Str54, SK80a, SK80b, Stu13, Tru96, Vog06, Vog14, Vor10, Wat83, Wat97, Wei90, Zim79, Nig03]. Meitner-Institut [Ano10, Lah91, Ano59a]. Meitner-Max [Hie00].

Meitner-Minerva [SKDA15]. Meitner/Ernst [ST68]. Meitner/Hawking [KZN^{+}88]. Meitner/Memel [SRRM^{+}97]. Meitners [Rom88]. Memel [SRRM^{+}97]. Memoirs [Whi00]. memorandum [Ber11]. Memorial [Mar54, Ano69, Rei72]. memoriam [Kar79, See56]. Memories [Her06, Mei05, Mei54]. Memory [Fri12, Her79, See56, Sim06]. men [Rom88]. Mensch [Mei58, Mei01]. Menschen [Ano12]. merged [MTVB88]. Merit [Cap57b]. Merite [Cap57a, Cap57b]. Merton [Wei90, Wei90]. Mesothor [Mei18a]. Mesothorium [Mei18a]. Messing [WSB^{+}99]. Messung [Mei43a]. metals [HS39]. Method [MF24, Mei43a, Mei43b, Mei43c, HM09b, Mei28e, Mei31b, Mei43d]. Methode [HM09b, Mei31b, Mei43a]. methods [Mei11b]. Meyer [Mei27a]. MGM [Rei84]. Mie [FKEE48, WHFB53]. Mikroskop [DRF90a, DRF90b]. Milena [Chi08]. Mileva [Chi08]. Millennium [Den99, Ano99]. Min [Gru04]. mind [Seg93]. Mineralien [HM20a]. minerals [HM21b, HM20a]. Minerva [SKDA15]. Minns [FH04]. mir [Ker91]. Miss [Fri02, Ano38, FH04, Fri12]. Missing [HM35a]. Mitarbeitern [Mei05]. Mitglied [DRF90a, DRF90b]. Mitglied/Neue [DRF90a, DRF90b]. Mitteil [HM35c]. Mitteilung [HM19d, HM21b, HM32]. mittels [HS39]. mixtures [Mei28e]. Mladjenovic [Mon98]. Model [Ano94, Mei50a, Stu94, LPS^{+}15, Mei22a, Mei50b]. modern [Mac11, Mei31a]. modernen [Mei31a]. mögliche [Mei23d]. Möglichkeit [HM19b, Mei21c]. momentous [McG93b, McG98b]. moments [Büh98b]. money [MTVB88]. Monograph [Mei28a, SH02a]. Monographie [SH02a]. Morality [Sch97]. Moskauer [DRF90a, DRF90b]. motion [Seg93]. move [Lah91]. moved [Ano12]. Moving [Wei72, Wei85]. Mt [NGD^{+}09, SN17]. München [Kra81c]. Münchner [Mei26e]. Munich [Kra81c, Mei26e]. MuPAD [SRRM^{+}97]. Museum [B^{+}79, DRF90a, DRF90b, Rec91a, Rec91b, Sim10]. Museums [Rec91a, Rec91b]. Muttersubstanz [HM18, Mei18b]. Myers [KZN^{+}88]. Mysterious [Bal12]. Mystery [FR13c].

N [For80]. N.Y [Bar05]. Nach [Ano46a, Ano31, HM13b, HM19b, KZN^{+}88, MF24, Mei25, Mei28e]. Nachbetrachtung [MLT01]. Nachweis [HS39, MF24, Mei25, vBHM11b, HM09c, Mei31b]. Named [Ano92, Lev97]. Namengebung [Wei90]. Names [Sto97]. naming [Wei90]. National [PC02, Bal14]. Nationalism [Cra92]. Nationalsozialismus [Mai07]. Natur [HM09c, HM13b]. Natural [Dan36, Föl91b, MD35c]. Nature [HM09c, Mei46a, HM13b, Mei17, Bal14, MF69, Sim96b]. Natürliche [Dan36, MD35c]. Naturwissenschaft [SRRM^{+}97]. Naturwissenschaften [Ano24a, Ano24b]. Naturwissenschaftlerin [Kra78b]. naturwissenschaftlerinnen [Föl91b]. Nazi [Ber16a, Mai07, MP01, WHF^{+}96]. Nazi-related [Mai07]. Nazism [Ols63]. near [Ano46b]. Nebelmethode [MF24, Mei28e]. necessary [Mei24e]. Neglect [Hoo02]. Nelson [Osg66, Sap14]. neptunium [Nod34]. Neu [MTVB88]. Neuburger [Mei22a]. Neuburgersche [Mei22a]. Neue [Ano46a, DRF90a, DRF90b, HM09b, HM10, Lah91, Sch57, HMS36, MH36]. neuen [Lah91]. Neuer [Ano46a]. Neuere [Mei26d, Mei34d]. Neues [Mei18b, HM08a, HM18, HMS38]. neutrino [Nav06]. Neutron [BM84, Fri39, Kae39, HMS36, HS39, MP33c, MHS37, MF39b, Mei41a, Mei41c, Mei48]. Neutronen [HM35a, HM35c, HMS36, HS39, MP32, MSH38, Mei48, HM35b, Mei34f, MP34, MH36]. Neutronen-Bestrahlung [HMS36]. Neutronenanregung [MP33c]. Neutronenbestrahlung [MHS37, Mei41c]. Neutrons [HM35a, HS89, MH36, MF39a, Mei40, MF68, MF69, HM35b, HM35c, MP32, Mei34f, MP34, MSH38]. Newer [Mei34d]. News [Ano66, MTVB88, Stu85, DRF90b]. Newton [Büh98b]. Newtons [Büh98b, SRRM^{+}97]. ni [FH04]. nicht [Büh98a, CSW97a]. Niederschlags [vBHM11c, HM12b, Mei12]. Niels [Ano71, Ben00, KZN^{+}88, SM08]. Nishina [MH30b, MH30a, MH30b, MH30a]. Niveaus [Mei31b]. No [Kra81c, Swe06, Mei56, Mei68]. Nobel [Bar05, Kra02, TF99, Ano46b, Bar05, CHU87, Cra92, CSW96, CSW97a, CSW97b, CSW98, Cra02, Dar04, Föl91b, Fri01, Har02, Kur02, McG93b, MB94, McG98b, McL99, MTVB88]. Nobel-Frauen [Föl91b]. Nobel-Preis [Ano46a, Ano46b]. Nobelist [Wat97]. Nobelpreis [CSW97a, Föl91a, MTVB88, CSW98]. Nobelpreis” [TF99]. Nochmals [Kra81c]. Noddack [Hoo03, Jon90, Ore03]. Nominations [Fre07, Lei09a]. Nominators [Bar05, CHU87]. Nominees [Bar05, CHU87]. non [Mah22]. non-radiative [Mah22]. note [Ben00, HM22, Mei22a]. notes [Bar05, For80]. Notiz [HM22, HM31b, Mei22a]. notwendige [Mei24e]. novel [Der07]. November [Lem03]. NPD [Ano13]. Nr [Kra81c]. Nuclear [Abb99, Ano56, Ano94, Ano10, Bas62, Ber16b, Fre06, Fre07, Fri68, Fri73, Gol01, Gra64, Gre84, Jen00, Kel96, KZN^{+}88, Lei09a, Lei09b, McL01, Mei34d, MF39a, Mei50a, Mei62, Mei63a, Mei63b, MF68, MF69, Mla98, Mon98, Mor98, Ree05, Rif06, Ros72, Seg85, She83, Sim91b, Sim98a, Sim98b, Sim00, Sim07b, Sim10, Sim12, Ste00, Stu94, Wes01, Zam14, Acz09, And73, Ano46b, Bar00, Bey49, CSW97b, Dan36, Ell08, Fri74b, GA71, HM36, Ham02, Hof78, Hof23, Kae48, Kei14, Ker98, Kra00, Mei21c, Mei22a, Mei32a, MD35c, Mei52a, MTVB88, MLT00, MLT01, Per97, Rad04, RCRC90, Ree15a, Rif99, Sim14b, Stu96, Stu13, Trö09, Zim79, ABB^{+}36, Bre36, Lei09b, Mei26a, Mei36a]. Nuclear-Fission [Sim10, Sim12]. Nuclear-Physics [Fri68]. Nuclei [Fri39, MD35k, JKS15, Mei26d, MP32, MD35c, Mei46b, Dan36]. nucleii [PMFB33]. Nucleus [FR13c, Mei39a, MF39c, Mei34a, Mei50b]. nur [CSW97a].

O [Ano59b, Bri61, Lin60, MH67, Sch60]. O. [Gar63a, MH67]. obiger [HM32]. Obituary [FH81]. observables [LPS^{+}15]. Observatoriums/Neu [MTVB88]. Observatory [MTVB88]. Observatory/News [MTVB88]. occasion [FPLR59b, Lin60, Mei52b, Neu21]. occurred [Mei25]. occurring [MP33c]. Océ [MTVB88]. Océ-van [MTVB88]. October [Ano03, Far01]. Odds [Kri99]. off [MSH38]. Öffentlichkeit [DRF90b, DRF90a]. old [Gre84].

Oliphant [Bre36]. once [Ols63]. once-great [Ols63]. Oncology [Den99]. One [Bar05, Ber16b]. only [Ano39, CSW97a, Cun13, Mei25]. Operation [Hof23]. Oppenheimer [Ken54]. Optics [SRRM^{+}97]. Optik [DRF90a, DRF90b]. orbits [MP33c]. Order [Cap57b]. ORF [Gri91, Gri91]. Origin [Ano94, Mei22d, Stu94, Fri74b, HM19a]. originally [Bey49]. other [Hof23, Rei72]. Otto [Ano49a, Ano49b, Ano69, Ano79b, Ano79a, B^{+}79, Bri61, CSW97a, Ern92, Ern93, FKEE48, Föl91a, For80, Hei69, Her80, Her10, Her06, Hof05, Kra81c, KF01, Lam79, Lin60, Lip59, Mei49, Mei52b, MWF55, Mei05, MLT00, MLT01, Mor11, Sch60, Sim94a, Sim07a, uRLS15, Vor10, Ano49a, Ano49b, Ano69, Ano79a, B^{+}79, CSW97a, EHM^{+}79, Ern92, Ern93, FPLR59b, Hah67, Hah79, Hei69, Her79, Her80, Her10, Hof78, Hof05, Kra81c, KF01, Lam79, Mei39b, Mei49, Mei52b, MWF55, Mei59, Mei64d, Mei79, MLT00, Ree05, She83, Sim06, Sim07a, Sim12, uRLS15, Sto89, Vor10, Wal03, Wal06, Zim79]. Otto-Hahn-Preise [MWF55]. our [Mac11, MLT00, MLT01]. Oxford [Bal12, Bar05, Gav04]. Oxford/New [Bar05].

P [Bre36, Bri61, For80, Lin60, Sch60]. P. [Bre36]. Paarläufe [FS06]. Pacific [Ken54]. pages [Gav04, Gav05, Her06, Kra81c, Swe06]. pair [FS06]. Paneth [Bri61, Lin60, Lip59, Mei32b, Sch60]. pantheon [Gav04]. paperback [Gav05, Her06, Nig03, Swe06]. Papers [Osg66]. Parallelism [For80]. parent [HM18, Mei18b]. part [Mei56, Mei68]. particles [Fis10, Fis12, Mei28a]. Passing [MF26b]. Pasteur [Ano46a, Ano46b, Ano46b]. Path [Ree05, Her86]. Patricia [Abb99, McL01, Rad04, Ste00, Swe06, Wes01, Gol01]. Paul [FKEE48, Lip59, MTVB88, MTVB88]. Paul-Scherrer-Institut [MTVB88]. peace [HM54]. Pechblende [HM19d, Mei15a]. Pechblenderückständen [HM22]. Peculiarities [PdK23]. People [Tel55, Ano12]. periodic [HM13b, Mei34a]. Periodischen [HM13b]. periodisches [Mei34a]. Personal [HFK58]. Personality [PC02]. Persönliche [HFK58]. Perspective [Seg85, Wat83, Zim79]. Philipp [For80]. Photoeffekt [MW33]. photoelectric [MW33]. Photographic [MF24]. Photographischer [MF24]. Phys [Bas62, Kra81c, Mei43b, Kra81c]. Phys. [Bas62]. Physical [Bas62, Fre07, Fri39, Lei09b, Mei29c, Bas62, Lah91, Mei36a]. Physicist [Ber16b, Gav04, Ker91, Ker98, Kra78b, KZN^{+}88, Whi00]. Physicists [DRRMH^{+}09, Gav04, Gav05, McL99, Rec02, Cro01a, Jam04b]. Physics [Ano03, Bar05, Bri61, CSW97b, Cro01b, Cun13, Fri12, Fri68, KZN^{+}88, Lei09b, Lin60, Mei36a, Mla98, Mon98, She83, ABB^{+}36, Bal14, Ber16c, Bey49, Bre36, Büh98b, Cal17a, CHU87, Dar04, FPLR59a, FPLR59b, Gou98, Her79, Her86, Jon90, Kae48, KLR13, KHFA67, Mei31a, Mic48, Rad99, Ree15a, Sch57, Seg07, Sim94b, Sim96a, Sim01, Sim02, WP85, Wei72, Wei85, Bar05, Bre36, KZN^{+}88, Mei29f, Whi00, AS96, Bey96, Gis97, Gol96, Hes96, Hof97, Joh96, KK97b, KK97a, Kel96, Kri99, Nye97, Wal96, Rou96]. Physics/Thorsen [KZN^{+}88]. Physik [Bri61, Büh98b, DRF90a, DRF90b, FPLR59a, Her86, Lip59, Sch60, Mei31a, Sch57, Sim01, SRRM^{+}97]. Physikalische [Bas62].

Physikalischen [Bre36, DRF90a, DRF90b, Mei36a, Mei29c, Lah91]. pictures [Hah79, Mei32b]. Pinzette [DRF90a, DRF90b]. Pioneer [Fri68, Fri73, RCRC90, Cra69, Ham02, RCRC97]. Pioneers [Hab79, Gol87, Sch91]. Pionierinnen [Sch91]. Pionniers [Gol87]. Pitchblende [HM19d]. pitchblue [HM22]. Planck [Ano46a, Ano46b, Ano49a, Ano49b, Bas62, Fis10, Fis12, HR01, Mei58, Mei01, MTVB88, Ano46b, Ano49a, Ano49b, Fis10, Fis12, HR01, Mei58, Mei01, RS80]. Planck-Gesellschaft [Bas62]. Planck-Gesellschaft/Hahn-Meitner-Institut [Bas62]. Planck-Medaille [Ano49a, Ano49b]. Planck/Todesurteil [Ano46a, Ano46b]. plenipotentiaries [Mai07]. point [LPS^{+}15]. Politics [Bar05, Kra02, Sch97, Sim12, Fri01, Sim06, Sim14b]. Population [Bar05, CHU87, Cra92]. portal [Vog06, Wol10]. portrait [Föl91b, Ker91, Ano59b]. Portraits [For80, Gri91, LMOS15, Far01]. Porträt [Ano59b, Föl91b]. Porträts [LMOS15, Sch60]. position [Ano79a, HM13b, Her80, Her10]. positive [MP33b, Mei34b]. positiven [Mei34b]. positiver [MP33b]. positron [LR10]. Possibilities [Kae39]. possibility [HM19b, Mei21c]. Possible [Mei23d]. Posthumous [Joh96]. Postwar [CSW97b]. Pour [Cap57a, Cap57b]. Power [Ano16, Kae39, KZN^{+}88]. Pp [All96, Rou96, Bal12, Bar05, Bri61, Kra02, Lin60, Nig03, Sap14]. Präparat [Hah40]. Präsident [Ano46a, Ano46b]. Preceded [Ken54]. Precipitation [Mei12, vBHM11c, HM12b]. predict [LPS^{+}15]. Preis [Ano46a, Ano46b, CSW97a, Dan36, DRF90a, DRF90b, Mei32b, MTVB88, Sch60]. Preis/Verwundbarkeit [DRF90a, DRF90b]. Preise [MWF55]. Preisgeld [MTVB88]. Preiswerk [Bre36]. Prematurity [Hoo02]. Preparation [PC02, Hah40]. Presentation [Ano49a, Ano49b]. Presented [Neu21]. President [Ano46b]. Press [All96, Bal12, Bar05, Gav04, Gav05, Hof97, KK97a, Rou96, Wal96]. Preussischer [B^{+}79, Lem03]. Price [Dan36, Mei32b, CSW97a]. Primal [KZN^{+}88]. primären [Mei29e, MO30]. primärer [Mei43a]. Primary [Mei43a, Mei43b, Mei43c, Mei29e, MO30, Mei43d]. privat [For80]. Prize [Ano46b, Ano13, Bar05, Fre06, Fre07, Kra02, Lei09a, Lei09b, Zam14, vO02, MTVB88, Ano04, CSW97a, CSW97b, CSW98, Fri01, Lei09b, McG93b, MB94, McG98b, MWF55, MTVB88, TF99]. Prizes [Bar05, Cra02, CHU87, Bar05, Har02]. probability [Mei31b]. problem [MD35e, SRRM^{+}97]. Problem/Bührke [SRRM^{+}97]. Problemlage [MD35e]. Proceedings [Ano03]. Process [Ano39, Gar63a, Mei41b, CSW97b, Kei14, SRRM^{+}97]. Process/Haug [SRRM^{+}97]. Processes [Mei34d, MH36, Mei45a, HMS36, Mei23c, Mei24c, Mei26e, Mei29c, Mei36b, MSH38, Mei52a]. Produced [HS89, MHS37, Mei46b]. Produces [Ano39]. producing [HM09b]. Product [HMS38, HM08a, HM09b, MH13]. production [HM19b, NFG^{+}08, NGD^{+}09]. Products [HM25b, Mei39a, MF39c, vBHM13, HM09b, HM12b, HMS35, HM37, HM38, Mei09a, Mei09b, Mei11b, MF39b]. Produkt [HM08a, HM09b]. Produkte [HM12b, MH13]. Prof [Ano46a]. professions [Mei60]. Professor [Ano59b, Lei09b]. program [Ber16a]. Project [PdK23, Ree14, Ree15a, Sim12]. Prometheus [Bar05]. Prominence [Sim05]. Prominent [Sim05]. Promotionswillige [WSB^{+}99]. Promotionswillige/Quadbeck [WSB^{+}99]. Promotionswillige/Quadbeck-Seeger [WSB^{+}99]. Proof [HS89, Mei31b]. Properties [HM20c, HM19d, HM21b, Mei11a]. Protactinium [HM19b, HM19c, HM20c, HM19d, HM21b, HM22, Mei18c, Mei28c, Mei18c, HM20a]. Protactiniumgehalt [HM22]. Protactiniums [Mei28c, HM20c, HM31b]. Protaktiniums [HM19d, HM21b]. Protocols [Hof23]. Protokolle [Hof23]. Proving [Ken54]. Prozess [Kei14]. Prozesse [HMS36, Mei23c]. Prüfung [MH30b, MH30a]. Prussian [B^{+}79, Lem03]. publication [Bal14]. publications [For80]. Publisher [Dan36]. Publishers [Lin60]. Publishing [Mei29f, Bal14]. Pullin [SRRM^{+}97].

Quadbeck-Seeger [WSB^{+}99]. Quantensprung [Fis10, Fis12]. Quantum [SKDA15, SRRM^{+}97, Fis10, Fis12, KLR13, KHFA67, SRRM^{+}97]. Quest [KZN^{+}88]. question [HM13b, HM19b]. Questioning [SRRM^{+}97]. questions [And73, MWZG61]. questions/Further [MWZG61]. Qui [Dem15]. Quick [Fri50].

R [Bre36, Bri61, Gar63a, KK97b, KK97a, Lin60, Rou96, Sch60]. Ra [Mei15b, Mei15b]. Race [Lau51, Rob95, WM15, Bal14]. Radar [Ano46a]. Radar/Zwei [Ano46a]. Radiance [Sap14, Nel14]. Radiation [vBHM11b, Den99, HM09e, HM24, HM25b, Mei10, Mei22d, Mei22c, Mei23b, Mei23a, MF24, Mei28b, Mei28d, MH30b, MH31b, MD35d, Mei43a, HM10, Mei25, Mei29e, Mei29c, Mei29d, Mei29a, MH30a, MH32, MP33b]. Radiations [Mei42a]. radiative [Mah22]. Radioactinium [HM25b, vBHM13, HM13b]. Radioactiniums [vBHM13, HM13b]. Radioactive [Con16, HM12a, HM35a, Mei10, Mei18b, Mei22d, Mei46b, HM09a, HM09b, HM18, HM21a, HM36, HM37, HM38, Mei11b, Mei11a, FM11, Mei21c, Mei24c, RM09, Ano31, Mei36c]. Radioactivity [FR13a, GLR12, L’A07, Mor98, Str54, Mei11c, Mei23c, Mei34g, Mei34e, MD35i, RCRC97, Mei21a, Mei21b, MD35f, Mei32b, Bal12]. Radioactivity. [Mei29f, Mei27a]. radioaktive [HM36, HM37, HM38, FM11, Mei36c]. radioaktiven [Ano31, HM12a, HM21a, HM35a, Mei10, Mei11a, Mei21c, Mei24c]. radioaktiver [HM09a, HM09b, Mei11b, Mei22d, RM09]. Radioaktives [Mei18b, HM18]. Radioaktivität [Mei23c, Mei34e, Str54, Mei11c, Mei21a, Mei21b, Mei27a, Mei32b, Mei34g, MD35i, MD35f]. Radioaktivität. [Mei29f]. Radioaktivitätsforschung [Mor98]. Radiochemie [Mei52b]. radiochemistry [Mei52b]. Radioelectronics [PC02]. Radioelemente [HM08c]. radioelements [HM08c]. Radioisotope [vM55]. Radioisotopes [vM55]. Radiothor [vBHM15, Mei18a]. Radiothorium [Mei18a, Mei28b, Mei28d, vBHM15]. Radiothors [Mei28b, Mei28d]. Radiozünder [Ano46a]. Radium [vBHM11b, HM09c, HM24, Kae48, MO30, vBHM11a, HM09a, Mei09a, Mei09b, RM09, Ano46b, vBHM11b, HM09c, MO30]. Radiumfund [Ano46a, Ano46b]. Radiums [vBHM11a, HM09a, HM09e, Mei09a, Mei09b, RM09]. Range [MF24, Mei26c, Mei29d, vLM27]. rare [Mei41a, Mei41c]. Raster [DRF90a, DRF90b]. Raster-Tunnel-Mikroskop [DRF90a]. ratio [HM22]. Ray [Mei23d, Mei43b, Mei43d, Wat83]. Rays [HM23, Kae39, MF26b, Mei43b, Rob95, vBHM11a, vBHM11c, vBHM13, vBHM14, vBHM15, HM08c, HM08b, HM09d, HM12b, HM20b, Mei06a, Mei07, MH13, Mei15b, Mei22e, Mei22b, Mei24d, Mei24a, Mei25, Mei26c, Mei26d, Mei29e, Mei29b, MO30, Mei31b, MH31a, Mei32a, MW33, MHK33, MK33, MP33a, Mei34c, Mei37, MSH38, Mei42b, Mei09a, Mei09b]. Raysand [Mei43c]. RaySpectra [Mei43c]. Reaction [MF39a, MF68, MF69, And73]. reactions [NFG^{+}08, NGD^{+}09]. Readers [TF99]. Reality [RS80, MH67]. realm [Kae48]. Recalls [Lau51]. receive [CSW97b]. received [CSW97a]. Recently [CSW97b]. Recheninstitut [Ano46a]. Recheninstitut/Zwei [Ano46a]. recoil [Mei52a]. Recollections [HR01, HFK58]. Record [KSFN90, NCSW98]. recording [Fri74b]. recordings [vLM27]. Reden [HR01]. reflection [Mei06b, Mei06c]. Reflexionsformeln [Mei06b, Mei06c]. Refuge [DF14]. Refugee [Seg85]. Regime [MP01]. Reich [Sim06]. Reichweite [MF24, Mei26c, Mei29d]. Reichweitestreuung [vLM27]. Reihe [HM35a]. Reihen [MSH38]. Reinhard [Lei09b]. related [Mai07]. relation [Mei29c]. Relationship [Mor11, Sch97, Mei22b, Mei24a]. Relationships [KZN^{+}88]. Relativity [For80]. Release [Mei45a]. released [CSW97b]. relevant [Mai07]. Remarkable [Jam04b, Cal17a, Gav05]. remarks [HMS35, Kra81c, Mei48, HM32]. remember [FH04]. Remembering [Fri12, Fri02]. Reminiscences [Ros72]. Renneberg [SRRM^{+}97]. reply [Kra81c]. Report [Lee00, Mei10, KHFA67, Hah40, Har43, Kae45]. reported [Bey49]. Representation [Mor11]. Repression [Wal06, Wal03]. Reprinted [MF69]. Research [Ano56, Ano10, Bas62, For80, HM12a, Mor98, Ano46b, KZN^{+}88, Mai07, Mei32a, Wei90, Gre84, Hof05]. researchers [Hof23]. Researching [Str54]. residues [HM22]. Resistance [Hoo02]. Resonance [Mei41b]. Response [Ore03]. Responsibility [Wal06, Hof05, KF01, Lam79, Wal03]. Rest [Sim96b]. resulting [HS39, Mei29e]. Results [HM12a, Ano24a, Ano24b, KZN^{+}88]. Rev [Mei43b]. reveal [CSW97b]. Revealed [Kae45]. Review [Abb99, AS96, All96, Bal12, Bar05, Bey96, Bey09, Bri61, Dan36, Gav04, Gav05, Gis97, Gol96, Gol01, Hes96, Hie00, Joh96, KK97b, KK97a, Kel96, Kra02, Kri99, Lin60, Lip59, McL01, Mei28a, Mon98, Nig03, Nye97, Osg66, Rec02, Rou96, Sap14, Sim94a, Ste00, Swe06, Wal96, Wol10, MLT01, Ols63, Wes01, BBT^{+}22, BEE^{+}28, BvLM23, GOB^{+}32, HMB20, MG27, MnBB28, Mei29f, MJG^{+}31, Mei32b, MhKG37, Mor98, RMJ27, SM26, ZSM32, vLvRM^{+}23]. Reviews [For80, Whi00, WSB^{+}99, SRRM^{+}97]. revised [Mei28a]. Revolution [Kae48]. revolutionized [Con16]. revolutions [And98]. Rezension [Bri61, Dan36, Mei27a, Mei29f, Mei32b, Wol10]. Richardson [KZN^{+}88]. richtet [DRF90a, DRF90b]. Rife [Abb99, Gol01, Rad04, McL01, Ste00, Wes01]. Right [Mei62, Mei63a, Mei63b]. Righting [KSFN90]. Rise [Sap14, She83, Nel14]. Risikostudie [DRF90a, DRF90b]. rivalry [Acz09]. Rivals [Gol90]. Road [Bar05, Har02]. Roads [Mei62, Mei63a, Mei63b]. Robert [Bar05, For80, Kra02, Lip59]. Role [Ken54, Mai07, Mei24d]. Rolle [Mai07, Mei24d]. Roman [Der07]. Rome [Ano03]. Roosevelt [Dem15]. Rosbaud [Bri61, Lin60, Sch60, Lip59]. Rossiter [Whi00]. route [Her79]. Rowohlt [Nig03]. Rückstoßes [Mei52a]. rule [HM20b]. runs [FS06]. Russische [Ano46a]. Russischer [Ano46a]. Ruth [Bey96, Gis97, Gol96, Hof97, Joh96, Kel96, Kri99, Nye97, Rad99, Wal96, Whi00, AS96, Mor11]. Rutherford [HM31a, Osg66, HM31a, Mac11, Mar54].

S [Bre36, Bri61, KK97a, Ras81, Sch60]. Sabine [Sim94a]. Sanitt [SRRM^{+}97]. saved [Sim07a]. Scandinavia [DF14]. Scandium [Mei42a, Mei45b, Mei45c]. scattered [Ols63]. Scattering [MH30b, MH30a, MH31a, MH32, MK33, Mei34c, MHK33]. Schalenmodell [Mei50b]. Schardin [Ano79a, Ano79a]. Schatten [Kra81b, Kra81a, Ras81, Chi94]. Scheme [Mei12, Mei45b, Mei45c]. Scherrer [MTVB88]. Schiemann [Lem10, Lem10, Sch97]. Schiffssteuerung [Ano46a]. Schilpp [For80]. Schmetterlinge [Ano46a, Ano46b]. schnelle [Mei31b]. schneller [Mei48]. school [MWZG61]. Schrödinger [McL99]. Schul [MWZG61]. Schul- [MWZG61]. Schütte [Chi08]. Schütte-Lihotzky [Chi08]. Schweden [Sim14a]. Schweidler [Mei27a]. schweren [MH32]. schwereren [Mei48]. Science [Bar05, Cra02, Den99, Far05, Fre06, Hab79, Kae39, KSFN90, KK97b, KK97a, LMOS15, Lei09b, MB94, Sch97, Sim14b, Con16, Cra92, FS06, Fri01, Gri91, Har02, Her80, Her10, Lig05, McG93b, McG98b, Mei03, Pop16, Rad99, RCRC90, RCRC97, Ree15b, Rif90, Rif92, Stu96, Wal96, You08, vM55, Bar05, EMvL79, Fre07, Lei09a, MLT01, Neu21, SRRM^{+}97, Zam14, Bal12, Kra02, Swe06, Gru04]. Sciences [Ano24a, Ano24b, Rei07]. Scientific [Hoo02, Acz09, And98, Bey49, Ols63, Osg66, Ren04, RS80, Vor10]. Scientist [Kae45, Kra78a, Ols63, Hof78, Hof05, Ker91, KF01, Lam79]. Scientists [Bar05, Far05, MP01, Seg85, Whi00, Föl91b, Har02, Kur02, Bar05, Swe06]. scission [LPS^{+}15]. scission-point [LPS^{+}15]. Scribner [Sap14]. Search [Gea14, Sim00, Sim07b]. sechzigsten [HM31a, Mei39b]. Second [Mei27a, Mei28a, HM35c, HM21b, Sim07a]. Secondary [Mei43a, Mei43b, Mei43c, Mei43d]. Secret [Kae45, MTVB88]. Secretary [Lee00]. Sections [Mei40, Mei48]. Security [Ken54]. Seeger [WSB^{+}99]. Seglers [WSB^{+}99]. Segrè [Seg93]. Sein [For80, HM21b, HM20a]. Seine [HM21b, For80, HM24, HM20a, Mei23b, Mei23a]. Seinen [Ano46a, HM25b]. seiner [vBHM13, HM19b, Mei28b]. Seite [Kra81c]. Seiten [Her06, Hof97]. sekundären [Mei43a]. Selected [For80]. Selove [Whi00]. seltener [Mei41c]. Semiconductors [SRRM^{+}97]. Semiconductors/High [SRRM^{+}97]. Sensation [Kra81b, Kra81a, Ras81]. sentence [Ano46b]. separation [HJ43]. September [Ano03, Har43]. Serie [KK97a]. Serienherausgeber [KK97a]. Series [HM35a, HMS38, KK97b, Mei28a, Gri91, HM10, HM22, Mei11a, MH13, Mei25, MHS37, MSH38]. set [Dur03]. Seventh [Far01]. Sexl [Nig03]. shade [Chi94]. Shadow [Kra81a, Kra81b]. shadows [Chi08]. shaped [Mac11]. Shell [Mei50a, Mei50b]. shift [HM20b]. shores [Lah91]. short [HM08a, HM09b, Ker91, MH31a, MH32]. short-lived [HM08a]. short-wave [MH31a, MH32]. Shortwave [MH30b, MH31b, MH30a, MHK33, MK33]. sich [For80, Mei06b, Mei06c]. Sichtbarmachung [Mei24b, Mei26e]. Side [For80]. siècle [Ren04]. Siegbahn [Kac96, Tru96]. Siegfried [Sch98]. Sime [AS96, Hof97, KK97a, Rou96, Wal96, Bey96, Gis97, Gol96, Joh96, KK97b, Kel96, Kri99, Nye97, Rad99, Whi00]. Simes [Mor11, Mor11]. Simple [Mei43a, Mei43c, Mei11b, Mei43d, Mei43b]. simultaneously [HM20b]. SIN [MTVB88, MTVB88]. Single [Mei45a, HJ43]. Sir [Osg66]. Sites [HJP17]. Situation [Sim07a]. sixtieth [HM31a]. Sixty [FR13c, Bar99]. Skandinavien [DF14]. slower [Mei48]. smallest [Fis10, Fis12]. Smith [KZN^{+}88]. Smyth [Kae45]. Society [Bas62, Mei28a, Neu21, Bas62, Fre07, Lei09b, Mai07, Mei26e, Vog06, Wol10]. Sohn [Bri61, Lin60, Sch60]. Solvay [Far01]. Some [Mei45a, HM08c, Mei06b, Mei06c, Mei11b, Mei41c, Ols63, HMS35, Mei26b, Mei48, Mei54]. Sommer [Bre36]. Sonderforschungsbereich [KZN^{+}88]. Sonnet [Org46]. Soon [Lev97]. soul [Seg07]. Sound [Fri74b]. Source [Kae39, Lei68]. Sources [KHFA67]. space [Gou98]. Spaltung [Hah43, Mei50b]. späten [For80]. special [KZN^{+}88]. Specialist [Mor98]. species [HM33, HM36]. Spectra [HM25b, Mei22d, Mei22c, Mei43a, Mei43b, vBHM11a, Mei43d]. Spectrum [HM24, Mei23d, Mei23b, Mei23a, Mei28b, Mei28d, vBHM13, vBHM14, vBHM15, Mei28c, MP31, Mei34b]. speculation [Pei86]. Speeches [HR01]. speed [Bal14]. Spektra [Mei22c]. Spektren [vBHM11a, Mei22d]. Spektrum [vBHM13, vBHM14, vBHM15, vBHM12, MP31, Mei28b]. Spezialinventar [For80]. Split [Ano39]. Splits [Kae39]. splitting [Hah43, MSH38]. spread [vLM27]. Springer [Dan36]. Staatsbibliothek [B^{+}79, Lem03, Lem03]. Stage [Dur03]. staircase [Fis10, Fis12]. Stand [Ano31]. Start [SRRM^{+}97, WSB^{+}99]. Starting [PC02]. State [B^{+}79]. states [JKS15, Ber16d]. Stationen [Hof78]. Stations [Hof78]. statistical [LPS^{+}15]. Statt [Ano46a, Ano46b]. Stature [McL99]. status [Mei60]. Stefan [Mei27a]. Stellung [HM13b, Her80, Her10]. Stellungswechsel [Ano79a]. Sternstunden [Büh98b, SRRM^{+}97]. Stiftung [DRF90b]. Stiftung/Jürgen [DRF90a, DRF90b]. still [Kae48]. Stock [Lei09b]. Stockholm [Bar05, Ano56, Har02]. Störfall [Pop16]. Stories [Far05, Swe06]. Story [Kae45, KZN^{+}88, MP01, Ree05, Ker98, Ree15a]. Strahl [Mei22d, Mei22c]. Strahlen [vBHM11a, vBHM11b, vBHM11c, vBHM13, vBHM14, vBHM15, HM08c, HM08b, HM12b, HM20b, HM23, Mei10, vBHM12, MH13, Mei15b, MF24, Mei24d, Mei25, Mei26d, MF26b, Mei28c, Mei29e, Mei29d, MO30, Mei32a, MP33b, MP33a, Mei37, Mei43a, HM09d, Mei06a, Mei07, Mei09a, Mei09b, Mei15b, Mei22e, Mei22b, Mei24a, Mei26c, Mei28b, Mei29b, Mei29a, MH31a, MW33, MHK33, MK33, MP33a, Mei34c, MD35d, MSH38, Mei42b]. Strahlenspektrum [HM24, Mei23b, Mei23a, Mei28c, Mei28b, Mei28d]. Strahlenspektrums [Mei23d]. Strahlenumwandlungen [Mei28c]. Strahlspektren [HM25b, Mei43a]. Strahlung [HM09e, HM10, Mei15a, Mei25, MH32, Mei22c, Mei29e, MH30b, MH30a, MH31b]. Strassmann [MH67, CSW97a, Kra81c, Ras81, Kra81c, CSW97a, FH81, Kra81b, Kra81a, Kra81c, MH67, Sim03, SK80a, SK80b, Ras81, SK80a, SK80b]. Straßmann-Biographie [Ras81]. strategies [Bal14]. Streugesetz [MH31a]. Streuung [Mei26d, MH32, MK33, Mei34c, MHK33]. Streuungsformel [MH30b, MH30a]. structure [Dan36, Mei21c, Mei26a, Mei26d, Mei27b, MP31, MP33a, MD35c]. struggle [Seg07]. struggles [McG93b, McG98b]. Studienförderpreis [DRF90a, DRF90b]. Studiengang [DRF90a, DRF90b]. Studies [Bar05, Cra02, KK97b, KK97a, Rad99, Bey49, Cra92]. Study [Mei43a, Fis10, Fis12]. studying [Sch91]. Stuttgart [Her06, MTVB88]. Stuttgarter [MTVB88]. substance [HM18, Mei18b]. Substances [Mei10, Mei22d, HM20b, HM21a]. Substanzen [Mei10, HM20b, HM21a, Mei22d]. Successful [Ken54]. Suche [KZN^{+}88]. Suicide [Hen16]. suitable [Mei28e]. Summary [Mei10]. summer [Bre36]. superheavy [JKS15, SN17]. Superstrings [KZN^{+}88]. supplemented [Hof23]. Supposé [Gru04, Gru04]. supposed [Kan12]. supraleitendem [DRF90a, DRF90b]. sustained [And73]. Sweden [HJP17, Sim94b, Sim14a]. Swedish [CSW97b, Dur03, FH04]. Swiss [Bre36, Mei36a]. Symposion [Ano79a]. Symposium [Ano79a]. System [Mai07, Mei34a, HM13b]. Systematik [MD35k]. Systems [MD35k]. Szivessy [FKEE48]. Szivessy/Paul [FKEE48].

Tab [Ras81]. table [HM13b]. Tagen [MLT01, MLT00]. taken [CSW97a]. Tale [CSW97b, CSW96, Sce13]. Taschenbuch [Nig03]. Tauber [For80]. Teaching [PC02]. Team [Kra81c, Kra83]. Teamarbeit [SK80a, SK80b]. Teamwork [SK80b, SK80a]. Tech [SRRM^{+}97]. technical [MWZG61]. Technik [DRF90a, DRF90b, SRRM^{+}97, vM55]. Technik/Sanitt [SRRM^{+}97]. Techniques [KZN^{+}88]. Technischen [Bre36, Mei36a]. Technology [Bre36, vM55, Mei36a]. Teil [For80]. Teilchen [Fis10, Fis12]. Telegamma [Rob95]. tells [FS06]. Temperatures [KZN^{+}88]. Temperatures/Jaenicke [KZN^{+}88]. Terminänderung [Ano79a]. Tesla [DRF90a, DRF90b]. Test [Ken54]. Testing [MH30b]. Teubner [Mei27a]. Text [Mei32b]. textbook [Mei32b]. ThC [MF26b, MP31, MP33b, MP33a, MF26b, MP31, MP33b, MP33a]. Their [HM23, Mei22c, MF26b, Mei45a, Cal17a, FPLR59b, HMS34, Lin60, McG93b, McG98b, Mei21c, Mei26e, Mei31b, RCRC97]. them [Ols63]. Theodor [B^{+}79, B^{+}79]. Theodor-Zink-Museum [B^{+}79]. Theoretical [Mor98, CSW97b]. theoretischen [Mor98]. Theorie [MP31]. Theories [SRRM^{+}97]. Theorist [SM08]. Theory [BM84, For80, Gea14, KZN^{+}88, Sim96b, MP31]. there [Ols63]. Thermal [Mei40]. These [Mei43a, Mei43b, Mei43c, Gol96, Mei43d]. Third [Den99, KLR13, Sim06]. thirteen [Bey49]. Thomas [All96, Bar05]. Thor [HM36]. Thorium [vBHM15, HM09b, HM10, HM35a, HS89, Mei12, Mei18a, Mei39a, Mei40, Mei41b, vBHM11c, HM12b, HM36, Mei11a, MH13, MSH38, MF39b, HM09b, HM10, Mei18a]. Thoriumreihe [HM10, Mei11a, MH13]. Thoriums [HM35a, MSH38, vBHM11c, HM09b, HM12b, vBHM12, Mei12]. Thorsen [KZN^{+}88]. those [Rom88]. Thousand [WM15]. Three [And73, Ell08]. Tiger [BM84]. Time [Ano16, KZN^{+}88, Hof23, Kra78b, MLT00]. Times [Gav04, Kra78a, Rec02, See56, Cro01a]. timing [Kei14]. Todesurteil [Ano46a, Ano46b]. Tokyo [Bar05]. Tölz [Ano46a, Ano46b, Ano46b]. Tölz/Dr [Ano46a, Ano46b]. Tot [Ano46a]. Tot/Erste [Ano46a]. Trackdown [Ols63]. Traditions [KLR13]. Trakl [Chi08]. Trans [HMS37, HM37, HMS38, HM38]. Trans-Urane [HMS37, HM37, HM38]. Trans-uraniums [HM37, HM38]. Trans-Uranreihen [HMS38]. transfer [Mai07]. Transformation [HM35a, HMS38, MH36, HM35b, HM35c, HM37, HM38, Mei36b, MHS37, MSH38]. transformations [Dan36, Hah43, KLR13, MD35c]. transitions [Mah22]. translated [All96, For80]. Transport [SRRM^{+}97]. Transurane [Ano46a, HMS34, MH67, Mei37]. Transuranelementen [Mei36c]. Transuranic [Sto97, HMS34, Kan12, Mei36c]. Transuranium [HMS38, Sim00, HMS37, Mei37, MH67]. Traum [WSB^{+}99]. Trendelenburg [ML56, ML56]. Trends [FPLR59b, Lin60]. Trenneffektes [HJ43]. trials [Wal96]. Tribute [Rec91a, Rec91b]. Triumphs [Kri99]. True [MP01, Hen16]. Truman [Ano46a, Ano46b, Ano46b]. truth [RS80]. Tunnel [DRF90a, DRF90b]. TV [Gri91]. Twentieth [Bar05, Dar04, Lig05]. Twentieth-Century [Bar05, Dar04, Lig05]. Two [Lau51, McL99]. Type [MF39a, MF68, MF69]. types [Mei21c]. Typical [HM09e]. typische [HM09e].

U [Bal12, Mei23a, KZN^{+}88, Ras81, Sch60, WSB^{+}99, Mei23a]. U.S.A [Mei28a]. Überreichung [Ano49a, Ano49b]. Übersicht [Mor98]. Uddmanska [HJP17]. Ufern [Lah91]. UK£ [Gav04]. Ukraine [PC02]. umstritten [DRF90a, DRF90b]. umstritten/Internationale [DRF90a, DRF90b]. Umwandlung [HM25a, HM35a, HM35b, HM35c, Mei34f]. Umwandlungen [HM09a, RM09]. Umwandlungs [HMS36]. Umwandlungs-Prozesse [HMS36]. Umwandlungsprodukt [HMS38]. Umwandlungsprodukte [HMS35, HM37, HM38]. Umwandlungsprozesse [Mei36b, MH36, MSH38]. Umwandlungsreihen [MHS37]. unbounded [Kae48]. uncanny [Her06]. United [Ber16d]. Universal [Bar05]. Universe [Ano03, LT13, KZN^{+}88]. Universe/ [KZN^{+}88]. Universität [Ano46a, Vog06, Wol10, Bey09]. Universität/Russische [Ano46a]. University [All96, Bar05, Gav04, Gav05, Hof97, KK97a, PC02, Rou96, Wal96, Wol10, Lau51, Vog06]. Universums [KZN^{+}88]. Universums/Davies [KZN^{+}88]. Unpublished [Hen16]. Unravelling [FR13c]. unseren [MLT00, MLT01]. Untersuchung [Mei43a]. Untersuchungen [Mei36c]. Unveröffentlichte [Hen16]. unverwüstliche [Her06]. Up/Gambini [SRRM^{+}97]. Upon [MH36]. Uran [vBHM14, HM10, HM13a, HM14, HM23, HM25a, HM36, HMS35, HMS36, HM20a, Mei36b]. Uran- [HM10]. Uran-Mineralien [HM20a]. Urane [HMS37, HM37, HM38]. Uranium [Acz09, Ano39, HM23, Hah62, HS89, Kae39, MH36, MF39a, MF39c, Mei40, Mei45a, MF68, MF69, PdK23, vBHM14, HM10, HM13a, HM14, HM21b, HM25a, HMS35, HM35b, HM35c, HM36, HMS36, HM37, HM38, HS39, Hah43, KF01, HM20a, Mei36b, MHS37, MF39b, Sim07a]. uraniums [HM37, HM38]. Uranmineralien [HM21b]. Uranreihen [HMS38]. Urans [HM35b, HM35c, HMS36, HS39, Hah43, MH36, MHS37, HM37, HM38]. Uranspaltung [KF01, Sim07a]. Urkraft [KZN^{+}88]. Ursprung [HM19a]. USA [Ano46a]. USA./Prof [Ano46a]. Uses [Ano39]. Using [MF24, Mei45a]. Ute [All96]. UX [Mei23b, Mei23b].

V [Ano46a, Ano79b, Bre36, HM32, Mei27a, Mei32b]. V-1 [Ano46a]. V-1/B [Ano46a]. V-1/Zerstörung [Ano46a]. vapor [Mei28e]. various [Mei28e]. Vast [Ano39]. velocity [Mei29d]. Verantwortung [Hof05, KF01, Lam79, Wal03]. Verdrängung [Wal03]. Verhalten [HS39, MF26b, Mei41c, HMS37]. Verhaltensweisen [HMS34]. Verhältnisses [Mor11]. Verlag [Bri61, Dan36, Ras81]. Verlage [Ano46a]. Verlagsgesellschaft [Mei29f]. Verlegt [Ano46a]. verliehen [Ano79a]. verliehen/Broschüre [Ano79a]. vermehrte [Mei27a]. Verschiebungsregel [HM20b]. verschiedene [MF26b, Mei28e]. verschiedenen [Mei21c, Mei34e]. verschiedener [HJ43]. Versuche [Mei28e, MP34]. Versus [Rob95]. Verteilung [HM12b, MH13]. Verwundbarkeit [DRF90a, DRF90b]. vi [Bar05]. Vienna [Ren04]. vier [Trö09]. Viertel [Ren04]. Vieweg [Bri61, Lin60, Sch60]. viii [Bar05]. visualization [Mei24b, Mei26e]. Vogt [Wol10]. Vol [KK97a, KZN^{+}88, Mei29f, Osg66, Rad99]. vollständig [Hof23]. Volt [Ano39]. Volts [Ano39]. Volume [KK97b, Rei72, For80]. Vom [WSB^{+}99, B^{+}79, Vog06, Wol10, Bey09]. Vor [Hah67]. Vorgängen [Mei29c]. vorstehenden [Mei22a]. Vortrag [Mei26e]. Vorträge [Bre36, HR01, Mei36a, Mei11c].

W [Bar05, Bre36, For80, Kra02, Mei29f, Whi00]. Wagner [WHFB53].

Wagner/Gustav [WHFB53]. Walden [FKEE48]. Walter [Lei09b]. War [Ber16d, Sim12, Mai07, Sim07a]. war-relevant [Mai07]. waren [MH67]. Warmeleitung [Mei06d]. Warnow [For80]. wars [Acz09]. wartime [CSW96]. Warum [CSW97a]. was [MH67, MH67]. Washingtoner [Ano46a]. wave [MH31a, MH32]. wavelength [Mei22e]. way [Der07]. WE-Heraeus-Stiftung [DRF90a]. WE-Heraeus-Stiftung/Jürgen-Geiger-Preis/Verwundbarkeit [DRF90a]. Weapons [Ber16b]. Wechselbeziehungen [Mei31c]. Wechselwirkung [MP32]. Week [Kae39]. Weg [Der07, Her86]. Wege [Mei63b]. Weggefährten [Mei05]. weight [Mei31a]. Weinheim [Ras81]. Weinheim-Deerfield [Ras81]. Weitere [MWZG61, HMS35, Hof23, MP34]. weiterhin [DRF90a, DRF90b]. Weizsäcker [KF01, KF01]. Wellenlänge [Mei22e]. Welt [Ano12]. Weltgeschichte [Ano46a, Ano46b, Her86]. Weltgeschichte/Schmetterlinge [Ano46a, Ano46b]. Weltkrieg [Sim07a]. went [Ols63]. werden [Ker91, CSW97a, Mei25, MHS37]. were [Bey49]. Werk [Vor10, DS93]. Werner [Cap57a, Cap57b, KF01, Rec91a, Rec91b, Cap57b, KF01, Rec91a, Rec91b]. Wetterdienst [Ano46a]. Wetterdienst/Radiozünder [Ano46a]. Wheeler [Rad04]. where [Ols63]. whirl [Der07]. who [Ano12, Dem15, Dem15]. whole [Sim07a]. Wichtiges [WSB^{+}99]. Wieder [Ano46a]. Wieland [MWF55]. Wieland/Hans [MWF55, MWF55]. Wien [Mei29f]. Wild [LMOS15, Pei86]. wilde [LMOS15]. Wilhelm [Bey09, HSMH36, Hah40, Mai07, Mei54, Neu21, Sim05, Sim07a, Vog06, Wol10]. Wilhelm-Institut [HSMH36]. Wilhelm-Instituts [Hah40]. Will [Ker91]. William [Gav04, Rec02]. Williams [Bre36]. Willy [WHFB53]. Wilson [MF24, Mei28e, vLM27]. Wilson-Aufnahmen [vLM27]. Wilsonschen [MF24, Mei28e]. Wirbel [Der07]. Wird [Ker91, DRF90a, DRF90b, Her86]. Wirken [Kra81b, Kra81a, Ras81]. Wirklichkeit [MH67]. Wirkungsgrades [HJ43]. Wismuts [Mei15a]. Wissenschaft [Gri91, vM55, EMvL79, FS06, Gru04, Mei03, MLT01, Rif90, Rif92]. Wissenschaften [Neu21]. Wissenschaftler [Ano46a, Lam79]. Wissenschaftlerin [Ker91]. Wissenschaftlers [Hof05, KF01]. wissenschaftlichen [Vor10]. Wissenschaftsakademie [LMOS15]. Wissenschaftsgeschichte [Her80, Her10, Pop16]. Wissenstransfer [Mai07]. Wissensw [WSB^{+}99]. Wizard [Swe06]. Wohnsitz [Ano46a]. Wohnsitz/Lise [Ano46a]. Woman [Mei03, Der07, Gru04, Wal96]. Women [AS11, Chi08, Cun13, DS93, DRRMH^{+}09, Far05, Hab79, KSFN90, MB94, Mic48, Stu96, Swe06, Cal17a, Föl91b, Gou98, Jon90, McG93b, McG98b, Mei60, RCRC90, RCRC97, Rom88, Sch91, You08, Chi94, Whi00]. Words [Hei69]. Work [DS93, Kra81a, Lau51, Mar54, Tel55, HM25a, Hah40, Kra81b, Mei22a, Mei26d, Vor10]. working [HM21a]. works [Ano46b, KZN^{+}88]. Workshop [DRF90a, DRF90b]. World [Sim07a, WM15, Ano46b, Ano12, Ber11, Con16, Her86, Mac11, Ber16d, Sim12]. writing [MWZG61]. Wrong [Mei62, Mei63a, Mei63b].

X [HM10, HM13a, HM23, Kae39, Mei23a, Bar05, vBHM14, vBHM15, HM10, HM13a, HM23, HM25a, Rob95]. X-Rays [Kae39, Rob95]. XI [Bri61, Sch60]. XIII [KK97a]. xv [Bar05]. xvii [Bar05].

Year [FR13c]. Years [Ano71, Bar05, Mla98, Mon98, SKDA15, Ano46b, Bar99, Ern92, Ern93, Gre84, Hah67, Kae48, Sch91, Sim14b]. yield [HJ43]. York [Bal12, Bar05, Kra02, Lin60, Mei28a]. Yukawa [Jam04b, Gav05].

Z [You08]. zahlr [Ras81]. zahlreiche [Hof97]. Zeilinger [Fis10, Fis12]. Zeit [Kra78b, KZN^{+}88, See56, For80]. zeitgenössische [Hof23]. zeitliche [Kei14]. Zeitschrift [DRF90a, DRF90b]. Zentrifuge [HJ43]. Zentrifugenanordnungen [HJ43]. Zentrum [DRF90a, DRF90b]. Zerfall [MD35a, MD35b]. Zerfalls [Mei21c]. Zerfallsprodukte [HM09b, Mei09a, Mei09b, vBHM13, Mei11b]. Zerfallsprodukten [HM25b]. Zerfallsprozessen [Mei24c]. Zerfallsschema [Mei12, Mei45b]. Zerstörung [Ano46a]. Zerstreuung [Mei07]. Zertrümmerung [MD35j]. Ziel [WSB^{+}99]. Zink [B^{+}79]. Zuflucht [DF14]. Zugehört [Gru04]. Zukunft [Zim79]. Zum [Kra78b, Ano79a, B^{+}79, Bey09, Bri61, Fis10, Fis12, For80, HM31a, Kra81c, Lem03, Mei39b, Mei49, Mei52b, Mei59, Mei64d, Mei79, MTVB88, Sch60, Str54, Vog06, Wol10, WSB^{+}99, Ano79b, Vor10]. Zuordnung [HM23]. Zur [Ano46a, HM32, Lam79, SK80a, SK80b, Wei90, Ano46b, Bri61, FPLR59a, Gri91, HM09b, HJ43, KF01, Lip59, Mei22a, Mei26b, Mei43a, Mei63b, Mor98, Neu21, Sch60, vG32, HM13b, HM33, Mei52b, Str54]. Zürich [Bas62, Bre36, Bre36]. Zürich/Phys [Bas62]. Zusammenarbeit [Hah67]. Zusammenfassender [Mei10]. zusammengelegt [MTVB88]. zusammengelegt/Stuttgarter [MTVB88]. Zusammenhang [Mei22c, Mei22b, Mei24a]. Zwei [Ano46a]. Zweite [Mei27a]. Zweiten [Sim07a]. zwischen [Lam79, Mei22b, Mei24a, Mei31c, MP32]. Zyklotrons [Ano46a]. Zyklotrons/Um [Ano46a].

References

Auger:1936:KGN

[ABB^{+}36] P. Auger, G. Bernardini, W. Bothe, J. Clay, J. D. Cockcroft, J. R. Dunning, R. Fleischmann, S. Franchetti, H. Geiger, H. v. Halban, L. Meitner, M. L. E. Oliphant, P. Preiswerk, and E. J. Williams. Kernphysik. (German) [Nuclear physics]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1936. Edited by E. Bretscher.

Abbott:1999:BRL

[Abb99] Steve Abbott. Book review: Lise Meitner and the Dawn of the Nuclear Age, by Patricia Rife . Mathematical Gazette, 83(498):

538–539, November 1999. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Aczel:2009:UWS

[Acz09] Amir D. Aczel. Uranium wars: the scientific rivalry that created the nuclear age. Palgrave Macmillan, New York, NY, USA, 2009. ISBN 0-230-61374-8. 248 + 8 pp. LCCN QC773.A1 A28 2009.

Allen:1996:BUH

[All96] Garland E. Allen. Book review: Biologists under Hitler, Ute Deichmann, translated by Thomas Dunlap. Pp 335. Harvard University Press, 1996. ISBN 0-674-07404-1. Endeavour, 20(4): 178, January 1996. CODEN ENDEAS. ISSN 0160-9327 (print),

1873-1929 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0160932797889417.

Alper:1989:FFA

[Alp89] Tikvah Alper. False fission attribution. Physics World, 2(3):14, March 1989. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-

7058 (electronic). URL http://stacks.iop.org/2058-7058/
2/i=3/a=14.

Anderson:1973:TQA

[And73] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3–7, March/April 1973. ISSN 0041-9508. URL http://library.

ucsd.edu/dc/object/bb3960604v.

Andersen:1998:CSR

[And98] Hanne Andersen. Characteristics of scientific revolutions. Endeavour, 22(1):3–6, January 1998. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016093279801093X.

Anonymous:1924:EENa

[Ano24a] Anonymous. Ergebnisse der Exakten Naturwissenschaften. (German) [Results of Exact Sciences], volume 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1924.

Anonymous:1924:EENb

[Ano24b] Anonymous. Ergebnisse der Exakten Naturwissenschaften. (German) [Results of Exact Sciences]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1924.

Anonymous:1931:RKN

[Ano31] Anonymous. Die radioaktiven Konstanten nach dem Stand von 1930. (German) [Radioactive constants as of 1930]. Physikalische Zeitschrift, 32:569–??, ???? 1931. CODEN PHZTAO. ISSN 0369-982X. Mitglieder (Members): M. Curie, A. Debierne, A. S. Eve, H. Geiger, O. Hahn, S. C. Lind, St. Meyer, E. Rutherford, E. Schweidler. Experten (Experts): J. Chadwick, I. Joliot-Curie, K. W. F. Kohlrausch, A. F. Kovarik. L. W. McKeehan, L. Meitner U. H. Schlundt.

Anonymous:1938:MLM

[Ano38] Anonymous. Miss Lise Meitner. Nature, 142(3602):856–866, November 12, 1938. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v142/n3602/pdf/142865d0.pdf.

Anonymous:1939:VEF

[Ano39] Anonymous. Vast energy freed by uranium atom: Split, it produces 2 ‘cannonballs’, each of 100,000,000 electron volts: Hailed as epoch making; new process, announced at Columbia, uses only 1–30 volt to liberate big force. New York Times, ??(??): 18, January 31, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/102759255/.

Anonymous:1946:RBTa

[Ano46a] Anonymous. Radiumfund Bei Bad Tölz/Dr. Eckener Arbeitet/Präsident Truman Zur Forschung/Goethe-Preis für Max Planck/Todesurteil Gegen Chemiker/Das Grösste Forschungsunternehmen Der Weltgeschichte/Schmetterlinge Statt Atombomben/Nobel-Preis für Medizin/Pasteur 50 Jahre Tot/Erste Kunde Von V-1/B. Baruch Über Die Forschung in Deutschland/Max Planck Verlegt Seinen Wohnsitz/Lise Meitner an der Washingtoner Universität/Russische Entdeckung in der Atomforschung/Wieder Astronomisches Recheninstitut/Zwei Berliner Verlage Wieder Eröffnet/88 Deutsche Wissenschaftler Nach Amerika/Nobel-Preis für Deutschen Gelehrten/Neuer Bayrischer Wetterdienst/Radiozünder Gegen V-1/Zerstörung der Japanischen Zyklotrons/Um Die Atomkontrolle/Ehrung Russischer Gelehrter/Farben-Fernsehen in USA./Prof. Hahn Wieder in Deutschland/Amerikanischer Bericht Zur Deutschen Atomforschung/Schiffssteuerung Durch Radar/Zwei Neue Transurane?/ Experiment Beim Bikini-Atoll. Physikalische Blätter, 2(1):17– 22, 1946. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1946:RBTb

[Ano46b] Anonymous. Radiumfund bei Bad Tölz/Dr. Eckener arbeitet/ Präsident Truman zur Forschung/Goethe-Preis für Max Planck/ Todesurteil gegen Chemiker/das grösste Forschungsunternehmen der Weltgeschichte/Schmetterlinge statt Atombomben/Nobel-Preis für Medizin/Pasteur 50 jahr. (German) [Radium discovery near Bad Tölz /Dr. Eckener works /President Truman on research /Goethe Award for Max Planck /Death sentence against chemist /The largest research firm in world history /Butterflies instead of nuclear bombs /Nobel Prize for Medicine /Pasteur 50 years anniversary]. Physik Journal, 2(1):17–22, 1946. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1949:UPMa

[Ano49a] Anonymous. Überreichung der Planck-Medaille an Otto Hahn und Lise Meitner durch Max von Laue. (German) [Presentation of the Planck Medal to Otto Hahn and Lise Meitner by Max von Laue]. Physikalische Blätter, 5(10):471–473, 1949. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1949:UPMb

[Ano49b] Anonymous. Überreichung der Planck-Medaille an Otto Hahn und Lise Meitner durch Max von Laue. (German) [Presentation of the Planck Medal to Otto Hahn and Lise Meitner by Max von Laue]. Physik Journal, 5(10):471–473, October 1949. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1956:LML

[Ano56] Anonymous. Lise Meitner: die Leiterin des Kernforschungsinstitutes in Stockholm. (German) [Lise Meitner: the Director of the Nuclear Research Institute in Stockholm]. Du : kulturelle Monatsschrift, 16(1):59, 1956. ISSN 0012-6837. URL https:/

/www.e-periodica.ch/digbib/view?pid=dkm-001:1956:16::

680#680. Only captioned photograph of Lise Meitner and Otto Hahn.

Anonymous:1959:HMI

[Ano59a] Anonymous. Hahn–Meitner-Institut. Physik Journal, 15(7):325, July 1959. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1959:PPH

[Ano59b] Anonymous. Porträt Professor O. Hahn und Professor L. Meitner. (German) [Portrait of …]. Zeitschrift für Physik, 154(3):i, June 1959. CODEN ZEPYAA. ISSN 0044-3328.

Anonymous:1966:NBF

[Ano66] Anonymous. News in brief: Fermi Award. Science, 153(3737): 723, August 12, 1966. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1967:ALM

[Ano67] Anonymous. Award for Lise Meitner. International Atomic Energy Agency Bulletin, 9(2):17, ???? 1967. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http://

www.iaea.org/Publications/Magazines/Bulletin/Bull092/

09201401717.pdf. Enrico Fermi Award for 1966.

Anonymous:1969:AGE

[Ano69] Anonymous. Akademische Gedenkfeier zu Ehren von Otto Hahn und Lise Meitner am 21. Februar 1969 in Berlin. (German) [Academic memorial in honor of Otto Hahn and Lise Meitner on 21 February 1969 in Berlin]. Maz-Planck-Gesellschaft [zur Förderg. d. Wissenschaften], München, West Germany, 1969. ???? pp. LCCN QC15 .A38.

Anonymous:1971:YNB

[Ano71] Anonymous. 50 years Niels Bohr Institute, Copenhagen. Europhysics News, 2(7):3–4, ???? 1971. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL https://www.

europhysicsnews.org/articles/epn/abs/1971/07/epn19710207p3/
epn19710207p3.html.

Anonymous:1979:TDA

[Ano79a] Anonymous. Terminänderung!/Deutsche Arbeitsgemeinschaft Akustik (DAGA)/Feier der 100. Geburtstage von Albert Einstein, Otto Hahn, Lise Meitner und Max von Laue/Einstein-Symposion Berlin/Hubert-Schardin-Medaille verliehen/Broschüre zum Stellungswechsel. (German) [Date change! /German Association for Acoustics (DAGA) /celebration of the 100th birthdays of Albert Einstein, Otto Hahn, Lise Meitner and Max von Laue /Einstein Symposium Berlin /Hubert Schardin medal awarded /Brochure on the change of position]. Physik Journal, 35(1):43– 46, January 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1979:GAE

[Ano79b] Anonymous. Zum 100. Geburtstag von Albert Einstein, Max v. Laue, Otto Hahn und Lise Meitner. (German) [On 100th birthdays …]. Physik Journal, 35(5):189–191, May 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1992:NEN

[Ano92] Anonymous. New elements named. Europhysics News, 23(9): 180, ???? 1992. CODEN EUPNAS. ISSN 0531-7479 (print),

1432-1092 (electronic). URL https://www.europhysicsnews.
org/articles/epn/pdf/1992/09/epn19922309p180.pdf.

Anonymous:1994:EOL

[Ano94] Anonymous. Errata: The Origin of the Liquid-Drop Model and the Interpretation of Nuclear Fission. Perspectives on Science, 2(2):254, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Correction of printers error in bottom three equations from page 90.

Anonymous:1997:HMI

[Ano97] Anonymous. Hahn-Meitner-Institut Berlin. Physik Journal, 53 (12):1249–1260, December 1997. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1999:MM

[Ano99] Anonymous. 1000 makers of the millennium. DK Pub., New York, NY, USA, 1999. ISBN 0-7894-4709-6, 0-7513-5664-6. 256 pp. LCCN CT107 .A16 1999.

Anonymous:2003:PIC

[Ano03] Anonymous, editor. Proceedings of the International Conference “Enrico Fermi and the Universe of Physics”, Rome, September 29–October 2, 2001. Accademia Nazionale dei Lincei and Instituto Nazionale di Fisica Nucleare, Rome, Italy, 2003. ISBN 88-

8286-032-9. LCCN ???? URL http:old.enea.it/produzione_
scientifica/pdf.../V2003_AttiFermi.pdf.

Anonymous:2004:LMP

[Ano04] Anonymous. 2004 Lise Meitner Prize. Europhysics News, 35(3): 97, May/June 2004. CODEN EUPNAS. ISSN 0531-7479 (print),

1432-1092 (electronic). URL https://www.europhysicsnews.
org/articles/epn/pdf/2004/03/epn2004-35-3.pdf.

Anonymous:2010:HMI

[Ano10] Anonymous. Hahn–Meitner-Institut für Kernforschung Berlin. (German) [Hahn–Meitner Institute for Nuclear Research Berlin]. Nachrichten aus Chemie und Technik, 9(14):223, April 2010. CODEN NCHTAD. ISSN 0027-738X.

Anonymous:2012:MWB

[Ano12] Anonymous. Menschen, die die Welt bewegten [Kurzfassungen]. (German) [People who moved the world [Abstracts]]. Menschen, die die Welt bewegten. Reader’s Digest Deutschland, Schweiz, Österreich, Stuttgart, Germany, 2012. ISBN 3-89915-827-X. 469

pp. LCCN ???? URL http://d-nb.info/1019418915/04.

Anonymous:2013:ENL

[Ano13] Anonymous. EPS-NPD Lise Meitner Prize 2012. Europhysics News, 44(1):6, 2013. CODEN EUPNAS. ISSN 0531-7479 (print),

1432-1092 (electronic). URL https://www.europhysicsnews.
org/articles/epn/pdf/2013/01/epn2013-44-1.pdf.

Anonymous:2016:MTA

[Ano16] Anonymous. The march of time: Atomic power! Periscope Film LLC archive, April 17, 2016. URL https://www.youtube.com/

watch?v=iTJJ7tkD5L4.

Ajzenberg-Selove:1996:BRR

[AS96] Fay Ajzenberg-Selove. Book review: Ruth Lewin Sime, Lise Meitner: A Life in Physics. Physics Today, 49(6):55–58, June 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Apotheker:2011:EWC

[AS11] Jan Apotheker and Livia Simon Sarkadi, editors. European Women in Chemistry. Wiley-VCH Verlag GmbH & Co. KGaA, Weinheim, Germany, 2011. ISBN 3-527-32956-0, 3-527-63645-5 (e-book), 3-527-63647-1 (e-book). 239 pp. LCCN QD20 .E976 2011.

Beck:1979:GGA

[B^{+}79] F. (Friedrich) Beck et al., editors. Gedächtnisausstellung zum 100 Geburtstag von Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner: in der Staatsbibliothek Preussischer Kulturbesitz, Berlin, vom 1. März–12. April 1979: im Theodor-Zink-Museum, Kaiserslautern, vom 21. Juni–31. Juli 1979. (German) [Commemorative exhibition for the 100th birthday of Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner at the State Library of Prussian Cultural Heritage, Berlin, from 1 March to 12th April 1979: the Theodor Zink Museum, Kaiserslautern, 21 June– 31 July 1979]. Physik Kongress-Ausstellungs- und Verwaltungs GmbH, Bad Honnef, West Germany, 1979. 151 pp. LCCN QC16.E5 G43 1979.

Bales:2012:BRR

[Bal12] Ellen Bales. Book review: Radioactivity: A History of a Mysterious Science, Marjorie C. Malley, Oxford U. Press, New York, 2011. $21.95 (267 pp.). ISBN 978-0-19-976641-3. Physics Today, 65(10):58, October 2012. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Baldwin:2014:KRP

[Bal14] Melinda Baldwin. ‘Keeping in the race’: physics, publication speed and national publishing strategies in Nature, 1895– 1939. British Journal for the History of Science, 47(2):257– 279, June 2014. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://journals.cambridge.
org/abstract_S0007087413000381. Published online, but not

yet assigned to a journal volume.

Bankston:2004:LMA

[Ban04] John Bankston. Lise Meitner and the Atomic Age. Unlocking the secrets of science. Mitchell Lane Publishers, Bear, DE, USA, 2004. ISBN 1-58415-206-0. 48 pp. LCCN QC774.M4 B36 2004.

Bartlett:1999:JSY

[Bar99] Albert A. Bartlett. Just sixty years ago. The Physics Teacher, 37(1):30–31, January 1999. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Barron:2000:LMD

[Bar00] Rachel Stiffler Barron. Lise Meitner: discoverer of nuclear fission. Great scientists. Morgan Reynolds, Greensboro, NC, USA, 2000. ISBN 1-883846-52-8. 112 pp. LCCN QC774.M4 B37 2000.

Bartholomew:2005:ERO

[Bar05] James R. Bartholomew. Essay review: One hundred years of the Nobel Science Prizes: Elisabeth Crawford (Editor). Historical Studies in the Nobel Archives: The Prizes in Science and Medicine. viii + 161 pp., Tokyo: Universal Academy Press, 2002. Elisabeth Crawford. The Nobel Population, 1901–1950: A Census of the Nominators and Nominees for the Prizes in Physics and Chemistry. vi + 420 pp., Tokyo: Universal Academy Press, 2002. Mauro Dardo. Nobel Laureates and Twentieth-Century Physics. x + 515 pp., Cambridge: Cambridge University Press, 2004. Robert Marc Friedman. The Politics of Excellence: Behind the Nobel Prize in Science. xv + 400 pp., notes, index. New York: W. H. Freeman, 2001. István Hargittai. The Road to Stockholm: Nobel Prizes, Science, and Scientists. xvii + 342 pp., Oxford/New York: Oxford University Press, 2002. George Thomas Kurian. The Nobel Scientists: A Biographical Encyclopedia. 675 pp., Amherst, N.Y.: Prometheus Books, 2002. Isis, 96 (4):625–632, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/498605;   http://www.jstor.org/stable/

3652242. See [CHU87, Cra92, Fri01, Cra02, Har02, Kur02, Dar04].

Bas:1962:JPG

[Bas62] E. B. Bas. 75 Jahre Physikalische Gesellschaft Zürich/Phys. Ges. in der DDR/Physical Society London/Max Planck-Gesellschaft/ Hahn-Meitner-Institut für Kernforschung Berlin. (German) [75th anniversary of the physical society zürich /Phys. Ges. in the DDR /Physical Society London /Max Planck Society /Hahn Meitner Institute for Nuclear Research Berlin]. Physik Journal, 18(8):374–375, August 1962. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Baschin:1922:BGR

[BBT^{+}22] O. Baschin, H. Bluntschli, Thürnau, H. Cloos, K. Freudenberg, L. Meitner, and Fritz Braun. Besprechungen. (German) [Review]. Die Naturwissenschaften, 10(9):205–210, March 1922.

CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Beer:1928:BGR

[BEE^{+}28] Arthur Beer, R. Emden, P. P. Ewald, Friedrich Drenckhahn, Lise Meitner, W. Grotrian, and Arn Berliner. Besprechungen. (German) [Review]. Die Naturwissenschaften, 16(18):316–321, May 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bentzen:2000:LMN

[Ben00] Soren M. Bentzen. Lise Meitner and Niels Bohr: A historical note. Acta Oncologica (Stockholm), 39(8):1002, 2000. CODEN ACTOEL. ISSN 0284-186X (print), 1651-226X (electronic). See [Den99].

Bernstein:2011:MCW

[Ber11] Jeremy Bernstein. A memorandum that changed the world. American Journal of Physics, 79(5):440–446, May 2011. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bernstein:2016:NGF

[Ber16a] Jeremy Bernstein. Nazi Germany’s failed program. In One Physicist’s Guide to Nuclear Weapons [Ber16b], chapter 4, pages 4:1– 4:11. ISBN 0-7503-1308-0. LCCN U264 .B47 2016. URL http:

//iopscience.iop.org/book/978-0-7503-1308-7.

Bernstein:2016:PF

[Ber16b] Jeremy Bernstein. One Physicist’s Guide to Nuclear Weapons. 2053-2563. IOP Publishing, Bristol, UK, 2016. ISBN 0-7503-1308-0. LCCN U264 .B47 2016. URL http://iopscience.iop.

org/book/978-0-7503-1308-7.

Bernstein:2016:PFa

[Ber16c] Jeremy Bernstein. The physics of fission. In One Physicist’s Guide to Nuclear Weapons [Ber16b], chapter 3, pages 3:1–3:7. ISBN 0-7503-1308-0. LCCN U264 .B47 2016. URL http://

iopscience.iop.org/book/978-0-7503-1308-7.

Bernstein:2016:USG

[Ber16d] Jeremy Bernstein. The United States and Great Britain in World War II. In One Physicist’s Guide to Nuclear Weapons [Ber16b], chapter 2, pages 2:1–2:5. ISBN 0-7503-1308-0. LCCN U264 .B47

2016. URL http://iopscience.iop.org/book/978-0-7503-
1308-7.

Beyer:1949:FNP

[Bey49] Robert T. (Robert Thomas) Beyer, editor. Foundations of nuclear physics: facsimiles of thirteen fundamental studies as they were originally reported in the scientific journals. Dover, New York, NY, USA, 1949. 272 pp. LCCN QC173 .B485. See original paper [HS39] and later translation [Gra64].

Beyler:1996:BRL

[Bey96] Richard H. Beyler. Book review: Lise Meitner: A Life in Physics by Ruth Lewin Sime. Isis, 87(4):746–747, December 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235247.

Beyler:2009:BRH

[Bey09] Richard Beyler. Book review: Vom Hintereingang zum Hauptportal? Lise Meitner und ihre Kolleginnen an der Berliner Universität und in der Kaiser-Wilhelm-Gesellschaft. Isis, 100(1): 189–190, March 2009. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/599686.

Bothe:1933:EAI

[BFG^{+}33] W. Bothe, H. Fränz, W. Gerlach, O. Hahn, G. Kirsch, L. Meitner, St. Meyer, F. Paneth, K. Philipp, K. Przibram, and H. Geiger, editors. Elektronen · Atome Ionen. (German) [Electrons · Atoms Ions]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1933.

Brown:1984:LTM

[BM84] Laurie M. Brown and Donald F. Moyer. Lady or tiger? — The Meitner–Hupfeld effect and Heisenberg neutron theory. American Journal of Physics, 52(2):130–136, February 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bretscher:1936:KVG

[Bre36] E. Bretscher, editor. Kernphysik. Vorträge gehalten am Physikalischen Institut der Eidgenössischen Technischen Hochschule Zürich im Sommer 1936 (30. Juni-4. Juli) von P. Auger, G. Bernardini, W. Bothe, J. Clay, J. D. Cockcroft, J. R. Dunning, R. Fleischmann, S. Franchetti, H. Geiger, H. von Halban jun., L. Meitner, M. L. E. Oliphant, P. Preiswerk und E. J. Williams. Hrsg. v. E. Bretscher. (German) [Nuclear physics. Lectures held at the Physics Institute of the Swiss Federal Institute of Technology in Zurich in the summer of 1936 (June 30–July 4) by P. Auger, G. Bernardini, W. Bothe, J. Clay, J. D. Cockcroft, J. R. Dunning, R. Fleischmann, S. Franchetti, H. Geiger, H. von Halban, Jr., L. Meitner, M. L. E. Oliphant, P. Preiswerk and E. J. Williams. Edited by E. Bretscher]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1936. 141 + 68 pp.

Brill:1961:RBP

[Bri61] R. Brill. [Rezension:] Beiträge zur Physik und Chemie des 20. Jahrhunderts, Lise Meitner, Otto Hahn, Max von Laue zum 80. Geburtstag, herausgeg. von O. R. Frisch, F. A. Paneth, F. Laves und P. Rosbaud Verlag Friedrich Vieweg & Sohn, Braunschweig 1959, 1. Aufl., XI, 285 S., geb. DM 29.50. (German) [[Book review:] Contributions to the Physics and Chemistry of the 20th Century, Lise Meitner, Otto Hahn, Max von Laue’s 80th Birthday, ed. by O. R. Frisch, F. A. Paneth, F. Laves and P. Rosbaud Verlag Friedrich Vieweg & Sohn, Braunschweig 1959, 1st ed., xi + 285 pp., DM 29.50]. Angewandte Chemie, 73(4):152, February 1961. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Bruckner:2000:MFF

[Brü00] Reinhold Brückner. Meitner and the forgotten fragment. Physics World, 13(5):20, May 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.

org/2058-7058/13/i=5/a=23. See [Kra00].

Buhrke:1998:IAE

[Büh98a] Thomas Bührke. “Ich habe die Atombombe nicht entworfen.” Lise Meitner (1878–1968). (German) [“I did not design the atomic bomb.” Lise Meitner (1878–1968)]. In Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner] [Büh98b], pages 231–255. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). LCCN ????

Buhrke:1998:NAS

[Büh98b] Thomas Bührke. Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner], volume 1202 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 1998. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). 258 pp. LCCN ????

Born:1923:BGR

[BvLM23] M. Born, M. v. Laue, and Lise Meitner. Besprechungen. (German) [Review]. Die Naturwissenschaften, 11(47):937–938, November 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Calvin:2017:BCF

[Cal17a] Scott Calvin. Beyond Curie: four women in physics and their remarkable discoveries, 1903 to 1963. IOP concise physics. Morgan and Claypool Publishers, San Rafael, CA, 94903, USA, 2017. ISBN 1-68174-644-1 (print), 1-68174-645-X (e-book), 1-68174-647-6 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). LCCN QC774.A2 C353 2017. URL http://iopscience.iop.

org/book/978-1-6817-4645-6.

Calvin:2017:I

[Cal17b] Scott Calvin. Introduction. In Beyond Curie: four women in physics and their remarkable discoveries, 1903 to 1963 [Cal17a], chapter 1, pages 1:1–1:4. ISBN 1-68174-644-1 (print), 1-68174-645-X (e-book), 1-68174-647-6 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). LCCN QC774.A2 C353 2017. URL http:

//iopscience.iop.org/book/978-1-6817-4645-6.

Calvin:2017:LM

[Cal17c] Scott Calvin. Lise Meitner. In Beyond Curie: four women in physics and their remarkable discoveries, 1903 to 1963 [Cal17a], chapter 3, pages 3:1–3:30. ISBN 1-68174-644-1 (print), 1-68174-645-X (e-book), 1-68174-647-6 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). LCCN QC774.A2 C353 2017. URL http:

//iopscience.iop.org/book/978-1-6817-4645-6.

Cap:1957:AMPa

[Cap57a] F. Cap. Arthur March/Der Pour le Merite für Lise Meitner und Werner Heisenberg. Physikalische Blätter, 13(7):317–318, July 1957. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19570130705/abstract.

Cap:1957:AMPb

[Cap57b] F. Cap. Arthur March/Der Pour le Mérite für Lise Meitner und Werner Heisenberg. (German) [Arthur March/the Order of Merit for Lise Meitner and Werner Heisenberg]. Physik Journal, 13(7): 317–318, July 1957. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Chiu:1994:FIS

[Chi94] Charles S. Chiu. Frauen im Schatten. (German) [Women in the shade]. J and V, Wien, Austria, 1994. ISBN 3-224-17669-5. 228 pp. LCCN CT3310 .C45 1994.

Chiu:2008:WSM

[Chi08] Ch. S. Chiu. Women in the shadows: Mileva Einstein-Marić, Margarete Jeanne Trakl, Lise Meitner, Milena Jesenská, and Margarete Schütte-Lihotzky, volume 40 of Austrian culture. Peter Lang, New York, NY, USA, 2008. ISBN 0-8204-8856-9 (hardcover). viii + 203 pp. LCCN CT3310 .C4513 2008. URL http:

//www.loc.gov/catdir/toc/ecip0819/2008023764.html. In-

troduction and translation by Edith Borchardt.

Crawford:1987:NPC

[CHU87] Elisabeth T. Crawford, J. L. Heilbron, and Rebecca Ullrich. The Nobel population 1901–1937: a census of the nominators and nominees for the prizes in physics and chemistry, volume 11; 4 of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1987. ISBN 0-918102-15-4 (paperback). 337 pp. LCCN QC28 .C74 1987.

Conkling:2016:RHI

[Con16] Winifred Conkling. Radioactive!: how Irène Curie and Lise Meitner revolutionized science and changed the world. Algonquin Young Readers, an imprint of Algonquin Books of Chapel Hill, Chapel Hill, NC, USA, 2016. ISBN 1-61620-415-X, 1-61620-555-5 (e-book). 227 pp. LCCN QC774.J65 C66 2016. URL http://

alsos.wlu.edu/information.aspx?id=4609.

Crawford:1969:LMA

[Cra69] Deborah Crawford. Lise Meitner, atomic pioneer. Crown Publishers, New York, NY, USA, 1969. 192 pp. LCCN QC774.M4 C7 1969.

Crawford:1992:NIS

[Cra92] Elisabeth T. Crawford. Nationalism and internationalism in science, 1880–1939: four studies of the Nobel population. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40386-3 (hardcover). xii + 157 pp. LCCN Q126.9

.C73 1992. URL http://www.loc.gov/catdir/description/
cam024/91033702.html;   http://www.loc.gov/catdir/samples/
cam031/91033702.html;   http://www.loc.gov/catdir/toc/
cam028/91033702.html.

Crawford:2002:HSN

[Cra02] Elisabeth T. Crawford, editor. Historical Studies in the Nobel Archives: the Prizes in Science and Medicine, volume 31 of Uppsala studies in the history of science. Universal Academy Press, Tokyo, Japan, 2002. ISBN 4-946443-69-X. ISSN 0282-1036. vii + 161 pp. LCCN AS911.N9 .H57 2002.

Cropper:2001:GPL

[Cro01a] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-

5. xii + 500 pp. LCCN QC15 .C76 2001. URL http://www.

loc.gov/catdir/enhancements/fy0637/2001021611-d.html;
 http://www.loc.gov/catdir/enhancements/fy0724/2001021611-
b.html.

Cropper:2001:PFL

[Cro01b] William H. Cropper. Physics and friendships: Lise Meitner. In Great physicists: the life and times of leading physicists from Galileo to Hawking [Cro01a], pages 330–343. ISBN 0-19-513748-

5. LCCN QC15 .C76 2001. URL http://www.loc.gov/catdir/
enhancements/fy0637/2001021611-d.html;   http://www.
loc.gov/catdir/enhancements/fy0724/2001021611-b.html.

Crawford:1996:NTW

[CSW96] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of wartime injustice. Nature, 382(6590):393–395, August 1, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v382/n6590/abs/382393a0.html.

Crawford:1997:KIP

[CSW97a] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Die Kernspaltung und ihr Preis. Warum nur Otto Hahn den Nobelpreis erhielt, Otto Frisch, Lise Meitner und Fritz Straßmann dagegen nicht berücksichtigt werden. (German) [Fission and its price. Why only Otto Hahn received the Nobel Prize: Otto Frisch, Lise Meitner and Fritz Strassmann are not taken into consideration]. Kultur & Technik. Zeitschrift des Deutschen Museums München, 21(2):30–35, ???? 1997. ISSN 0344-

5690. URL http://www.deutsches-museum.de/fileadmin/
Content/data/Insel/Information/KT/heftarchiv/1997/21-

2-30.pdf. Translation from English by Dieter Beisel of [CSW96].

Crawford:1997:NTP

[CSW97b] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of postwar injustice: Recently released Swedish documents reveal why Lise Meitner, codiscoverer of nuclear fission, did not receive the 1946 Physics Prize for her theoretical interpretation of the process. Physics Today, 50(9):26–32, September 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Crawford:1998:LMN

[CSW98] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Lise Meitner und der Nobelpreis. (German) [Lise Meitner and the Nobel Prize]. Physik in unserer Zeit, 29(6):234–241, 1998. CODEN PHUZAH. ISSN 0031-9252.

Cunningham:2013:WPA

[Cun13] Beth A. Cunningham. Women in physics [abstract only]. AIP Conference Proceedings, 1517:173, 2013. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Danneel:1936:RAA

[Dan36] H. Danneel. [Rezension:] Der Aufbau der Atomkerne. Von Lise Meitner und Max Delbrück: Natürliche und künstliche Kernumwandlungen. Verlag Julius Springer, Berlin 1935. Preis geh. RM. 4,50. (German) [[Book review:] The structure of atomic nuclei. By Lise Meitner and Max Delbrück: Natural and artificial nuclear transformations. Publisher Julius Springer, Berlin 1935.

Price RM. 4.50]. Angewandte Chemie, 49(9):171–172, February 1936. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Dardo:2004:NLT

[Dar04] M. (Mauro) Dardo. Nobel laureates and twentieth-century physics. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83247-0, 0-521-54008-9 (paperback). xi + 533 pp.

LCCN QC7 .D27 2004. URL http://www.loc.gov/catdir/
description/cam041/2004049240.html;   http://www.loc.
gov/catdir/toc/cam041/2004049240.html.

Deichmann:1996:BUH

[Dei96] Ute Deichmann. Biologists under Hitler. Harvard University Press, Cambridge, MA, USA, 1996. ISBN 0-674-07404-1. xviii + 468 pp. LCCN QH305.2.G3 D4513 1996. Translated by Thomas Dunlap. The Epilogue consists of a letter written by physicist Lise Meitner from England to her former colleague Otto Hahn, in Berlin; see [All96].

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “The letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125– 152, ???? 2015. CODEN RQSCAN. ISSN 0035-2160.

URL https://www.unamur.be/sciences/philosoc/revueqs/
textes-en-ligne/rqs_186_1_demortier.

Denekamp:1999:ETM

[Den99] Juliana Denekamp. Entering the third millennium after a century of ionizing radiation in science and oncology. Acta Oncologica (Stockholm), 38(7):819–821, 1999. CODEN ACTOEL. ISSN 0284-186X (print), 1651-226X (electronic). See comments and historical correction, and brief response by Denekamp [Ben00].

Derado:2007:IWA

[Der07] Thea Derado. Im Wirbel der Atome: Lise Meitner — eine Frau geht ihren Weg: ein biografischer Roman. (German) [In the whirl of atoms: Lise Meitner — a woman goes her way: a biographical novel]. Kaufmann, Lahr, Germany, 2007. ISBN 3-7806-3059-1. 239 pp. LCCN ????

Dahl:2014:SZJ

[DF14] Izabela A. Dahl and Jorunn Sem Fure. Skandinavien als Zuflucht für jüdische Intellektuelle 1933–1945. (German) [Scandinavia as a Refuge for Jewish Intellectuals 1933–1945]. Metropol-Verlag, Berlin, Germany, 2014. ISBN 3-86331-194-9. 326 pp. LCCN DS135.S32 S53 2014.

Dreisigacker:1990:ABFa

[DRF90a] E. Dreisigacker, F. Rieß, and G. Fischer. Aktuelles: Berliner Forschungsreaktor weiterhin umstritten/Internationale Fachausstellung Optik 1990 in Berlin/CERN: Finnland wird Mitglied/ Neue Zeitschrift der Moskauer Physikalischen Gesellschaft/ Bergakademie Freiberg richtet Studiengang Physik ein/Gesamtdeutsche DFG?/Deutsches Museum ehrt Meitner und Heisenberg/ DVT-Preis Technik und Öffentlichkeit /Das Raster-Tunnel-Mikroskop als Pinzette für Atome/Europäisches Astronautenzentrum in Köln/20,7 Tesla mit supraleitendem Magneten erreicht!/Zentrum für hochauflösende Elektronenmikroskopie im Forschungszentrum Jülich eingeweiht/Studienförderpreis der WE-Heraeus-Stiftung/Jürgen-Geiger-Preis/Verwundbarkeit der Industrielandschaft Europa/Risikostudie Kernkraftwerke/ Computer Workshop IV ein Bericht. Physikalische Blätter, 46 (6):189–198, June 1990. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19900460614/abstract.

Dreisigacker:1990:ABFb

[DRF90b] E. Dreisigacker, F. Rieß, and G. Fischer. Aktuelles: Berliner Forschungsreaktor weiterhin umstritten/Internationale Fachausstellung Optik 1990 in Berlin/CERN: Finnland wird Mitglied/ Neue Zeitschrift der Moskauer Physikalischen Gesellschaft/ Bergakademie Freiberg richtet Studiengang Physik ein/Gesamtdeutsche DFG?/Deutsches Museum ehrt Meitner und Heisenberg/ DVT Preis Technik und Öffentlichkeit /Das Raster Tunnel Mikroskop als Pinzette für Atome/Europäisches Astronautenzentrum in Köln/20,7 Tesla mit supraleitendem Magneten erreicht!/Zentrum für hochauflösende Elektronenmikroskopie im Forschungszentrum Jülich eingeweiht/Studienförderpreis der WE Heraeus Stiftung/Jürgen Geiger Preis/Verwundbarkeit der Industrielandschaft Europa/Risikostudie Kernkraftwerke/ Computer Workshop IV — ein Bericht. (German) [News: …]. Physik Journal, 46(6):189–198, June 1990. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Durstberger-Rennhofer:2009:WPA

[DRRMH^{+}09] Katharina Durstberger-Rennhofer, Monika Ritsch-Marte, Beverly Karplus Hartline, Renee K. Horton, and Catherine M. Kaicher. Women physicists in Austria. AIP Conference Proceedings, 1119(1), 2009. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Dick:1993:JFI

[DS93] Jutta Dick and Marina Sassenberg. Jüdische Frauen im 19. und 20. Jahrhundert: Lexikon zu Leben und Werk. (German) [Jewish Women in the 19th and 20th Centuries: Encyclopedia on Life and Work]. Rowohlt, Reinbek bei Hamburg, Germany, 1993. ISBN 3-499-16344-6. 415 pp. LCCN DS135.G5 A1445 1993.

Durrani:2003:SSM

[Dur03] Matin Durrani. Stage set for Meitner’s Swedish debut. Physics World, 16(4):6, April 2003. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.

iop.org/2058-7058/16/i=4/a=6. Based on the author’s book

[Fri01].

Einstein:1979:FGA

[EHM^{+}79] Albert Einstein, Otto Hahn, Lise Meitner, Max von Laue, et al. Feier der 100. Geburtstage von Albert Einstein, Otto Hahn, Lise Meitner, Max von Laue. (German) [Celebration of the 100th Birthdays of Albert Einstein, Otto Hahn, Lise Meitner, Max von Laue]. Paperback der Zeitschrift Naturwissenschaftliche Rundschau. Wissenschaftliche Verlagsgesellschaft, Stuttgart, West Germany, 1979. ISBN 3-8047-0594-4. 99 pp. LCCN QC15 .F44.

Ellmer:2008:DEK

[Ell08] Reinhold Ellmer. Die drei Entdecker der Kernspaltung. (German) [The three discovers of nuclear fission]. Nachrichten aus der Chemie, 56(12):1241–1243, December 2008. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic). See response [Trö09].

Einstein:1979:WGF

[EMvL79] Albert Einstein, Lise Meitner, and Max von Laue. Wissenschaft unter dem Gesetz der Freiheit. (German) [Science under the law of freedom]. Physik Journal, 35(5):192–199, May 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Ernst:1992:LMO

[Ern92] Sabine Ernst, editor. Lise Meitner an Otto Hahn: Briefe aus den Jahren 1912 bis 1924: Edition und Kommentierung. (German) [Lise Meitner to Otto Hahn: Correspondence from the years 1912 to 1924: editing and commentary], volume 65 of Quellen und Studien zur Geschichte der Pharmazie. Wissenschaftliche Verlagsgesellschaft mbH, Stuttgart, Germany, 1992. ISBN 3-8047-1254-1. vi + 267 pp. LCCN RS61 .Q8 Bd. 65. With a foreword by Fritz Krafft.

Ernst:1993:LMO

[Ern93] Sabine Ernst. Lise Meitner an Otto Hahn. Briefe aus den Jahren 1912 bis 1924. (German) [Lise Meitner to Otto Hahn. Letters from the years 1912 to 1924]. Chemie in unserer Zeit, 27(6): 296–302, December 1993.

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137–138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Fara:2005:SAG

[Far05] Patricia Fara. Scientists Anonymous: Great Stories of Women in Science. Wizard Books, Thriplow, Cambridge, UK, 2005. ISBN 1-84046-574-3 (paperback). 213 pp. LCCN Q141 .F27 2005.

Fresenius:1935:LGL

[FBLB^{+}35] L. Fresenius, H. Brückner, Landolt-Börnstein, E. Berl, H. Jörgensen, F. Krczil, W. Langenbeck, A. Weissberger, E. Proskauer, Fr. Schuster, Lise Meitner, M. Delbrück, and H. Sponer. Literatur. (German) [Literature]. Zeitschrift für analytische Chemie, 103 (3–4):119–122, March 1935. CODEN ABCNBP. ISSN 1618-2642 (print), 1618-2650 (electronic).

Friedlander:1981:OFS

[FH81] Gerhart Friedlander and Günter Herrmann. [Obituary:] Fritz Strassmann. Physics Today, 34(4):84–86, April 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Friedman:2004:MNF

[FH04] Robert Marc Friedman and Inger Hayman. Minns ni fröken Meitner? (Swedish) [Do you remember Miss Meitner?.]. Stockholm Radioteatern, 2004.

Fischer:2010:HQE

[Fis10] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2012:HQE

[Fis12] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fischer:1948:GHG

[FKEE48] Ch. Fischer, P. Kunze, Ihr J. Eggert, and F. Emde. Gerhard Hoffmann/Guido Szivessy/Paul Walden 85 Jahre/Gustav Mie 80 Jahre/Lise Meitner 70 Jahre/Max von Laue/Otto Hahn. (German) [Birthdays …]. Physik Journal, 4(8):347–350, August 1948. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Meitner:1911:RI

[FM11] J. Franck and Lise Meitner. Über radioaktive Ionen. (German) [On radioactive ions]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 13(17):671–675, September 15, 1911. CODEN VDPEAZ. ISSN 0372-5448. URL https://hdl.handle.net/

2027/coo.31924056113495?urlappend=;seq=687.

Folsing:1991:LMO

[Föl91a] Ulla Fölsing. Lise Meitner (Otto Hahn: Chemie-Nobelpreis 1945). In Nobel-Frauen: naturwissenschaftlerinnen im Porträt. (German) [Nobel women: natural scientists in the portrait] [Föl91b], pages 146–157. ISBN 3-406-34018-0. LCCN Q141 .F55 1991.

Folsing:1991:NFN

[Föl91b] Ulla Fölsing. Nobel-Frauen: naturwissenschaftlerinnen im Porträt. (German) [Nobel women: natural scientists in the portrait]. Verlag C. H. Beck, München, Germany, second edition, 1991. ISBN 3-406-34018-0. 214 pp. LCCN Q141 .F55 1991.

Forman:1980:BRE

[For80] Paul Forman. Book reviews: Einstein: A Centenary Volume, edited by A. P. French; Albert Einstein’s Theory of General Relativity, edited by Gerald E. Tauber; Einstein, by Louis de Broglie and others; Einstein. Sein Leben und seine Zeit, by Philipp Frank; Documentary publications: Albert Einstein Autobiographical Notes, translated and edited by Paul Arthur Schilpp; Elie Cartan, Albert Einstein, Letters on Absolute Parallelism 1929–1932, edited by Robert Debever; Aus meinen späten Jahren, by Albert Einstein; Albert Einstein in Berlin 1918–1933, Teil I: Darstellung und Dokumente; Teil II: Spezialinventar; Images of Einstein: a Catalog, compiled by Joan N. Warnow; Albert Einstein: the Human Side. New Glimpses from his Archives, edited by Helen Dukas and Banesh Hoffman; Einstein privat. Herta W. erinnert sich an die Jahre 1927 bis 1933, by Friedrich Herneck; Centenary exhibitions: Albert Einstein, 1879–1955. A Centenary Exhibit of Manuscripts, Books, and Portraits Selected from the Humanities Research Center Collections, by Albert C. Lewis; Gedächtnisausstellung zum 100. Geburtstag von Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner; Einstein = mc^{2}; Einstein 1879–1979. Exhibition . Bulletin of the Atomic Scientists, 36(3):49–55, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1959:BPC

[FPLR59a] O. R. Frisch, F. A. Paneth, F. Laves, and P. Rosbaud, editors. Beiträge zur Physik und Chemie des 20. Jahrhunderts. (German) [Contributions to the physics and chemistry of the 20th century]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1959. 285 pp. LCCN QC475 .B45 1959.

Frisch:1959:TAP

[FPLR59b] O. R. Frisch, F. A. Paneth, F. Laves, and P. Rosbaud, editors. Trends in atomic physics; essays dedicated to Lise Meitner, Otto Hahn, Max von Laue on the occasion of their 80th birthday: Atomic physics. Interscience Publishers, New York, NY, USA, 1959. 285 pp. LCCN QC475 .B45 1959a. An identical edition is published simultaneously under the title Beiträge zur Physik und Chemie des 20. Jahrhunderts.

Fernandez:2013:HMR

[FR13a] Bernard Fernandez and Georges Ripka. Hahn, Meitner and β radioactivity. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13c], page 184. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:LM

[FR13b] Bernard Fernandez and Georges Ripka. Lise Meitner. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13c], pages 182–184. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13c] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Freiesleben:2006:LMP

[Fre06] Hartwig Freiesleben. Lise Meitner Prize for Nuclear Science 2006. Europhysics News, 37(5):12, ???? 2006. CO-DEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL https://www.europhysicsnews.org/articles/
epn/pdf/2006/05/epn2006-37-5.pdf.

Freiesleben:2007:LMP

[Fre07] Hartwig Freiesleben. Lise Meitner Prize for Nuclear Science of the European Physical Society, 2008 [call for nominations]. Europhysics News, 38(6):5, ???? 2007. CO-DEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL https://www.europhysicsnews.org/articles/
epn/pdf/2007/06/epn2007-38-6.pdf.

Frisch:1939:PED

[Fri39] Otto Robert Frisch. Physical evidence for the division of heavy nuclei under neutron bombardment. Nature, 143 (3616):276, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

chemteam.info/Chem-History/Frisch-Fission-1939.html;
 http://www.nature.com/nature/journal/v143/n3616/pdf/
143276a0.pdf.

Friendly:1950:QD

[Fri50] Fred Friendly. The quick and the dead. Four-part NBC radio miniseries on the creation of the atomic bomb, narrated by American entertainer Bob Hope, with Helen Hayes playing Lise Meitner and with Enrico Fermi and other leading scientists from the Manhattan Project receiving

some mention., 1950. URL https://www.discogs.com/
National-Broadcasting-Company-The-The-Quick-And-The-
Dead-Volume-1-The-Atom-Bomb-/release/5350974;   https:
//www.questia.com/library/journal/1P3-1718351551/living-
with-the-bomb-fred-friendly-s-the-quick-and.

Frisch:1968:LMD

[Fri68] Otto R. Frisch. Lise Meitner dies; nuclear-physics pioneer. Physics Today, 21(12):101, December 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/scitation.aip.org/content/aip/magazine/physicstoday/
article/21/12/10.1063/1.3034630.

Frisch:1970:LM

[Fri70] Otto Robert Frisch. Lise Meitner. 1878–1968. Biographical Memoirs of Fellows of the Royal Society, 16:405–420, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/769597.

Frisch:1973:DNP

[Fri73] Otto Robert Frisch. Distinguished nuclear pioneer–1973. Lise Meitner. Journal of Nuclear Medicine, 14(6):365–371, June 1, 1973. CODEN JNMEAQ. ISSN 0161-5505 (print), 1535-5667

(electronic). URL http://jnm.snmjournals.org/content/14/

6/365.full.pdf+html. Reprinted from The Biographical Memoirs of Fellows of The Royal Society, Volume 16, November, 1970, pages 405–416.

Frisch:1974:LM

[Fri74a] Otto Robert Frisch. Lise Meitner. In Charles Coulston Gillispie, editor, Dictionary of scientific biography, pages 260–263. Scribner, New York, NY, USA, 1974.

Frisch:1974:ONF

[Fri74b] Otto Robert Frisch. The origin of nuclear fission. [Sound recording]. Spring Green Multimedia, Washington, DC, USA, 1974. LCCN QC790; RYB 6354. 1 cassette tape.

Friedman:2001:PEB

[Fri01] Robert Marc Friedman. The politics of excellence: behind the Nobel Prize in science. Times Books, New York, NY, USA, 2001. ISBN 0-7167-3103-7. xv + 379 pp. LCCN QC49 .F75 2001.

Friedman:2002:RMM

[Fri02] Robert Marc Friedman. ‘Remembering Miss Meitner’: an attempt to forge history into drama. Interdisciplinary Science Reviews, 27(3):202–210, ???? 2002. ISSN 0308-0188 (print), 1743-2790 (electronic).

Friedman:2012:RMM

[Fri12] Robert Marc Friedman. Remembering Miss Meitner: History, memory, and the future of physics. Europhysics Conference Abstracts (ECA), 36 (Part C)(??):??, ???? 2012. ISSN 0378-2271.

Friege:2017:MME

[Fri17] Johanna Friege. Meitner’s Matilda effect. Varsity, ??(??):??, Oc-

tober 12, 2017. URL https://www.varsity.co.uk/science/
13682.

Fischer:2006:EPF

[FS06] Ernst Peter Fischer and Klaus Sander. Ernst Peter Fischer erzählt, Paarläufe der Wissenschaft. (German) [Ernst Peter Fischer tells the pair runs of science]. Suppoš, Köln, Germany, 2006. ISBN 3-932513-68-1. LCCN ???? URL http:/

/deposit.ddb.de/cgi-bin/dokserv?id=2758637\%26prov=M\
%26dok\_var=1\%26dok\_ext=htm. 4 CDs.

Frisch:1967:DFH

[FW67] Otto Robert Frisch and John A. Wheeler. The discovery of fission: How it all began and mechanism of fission. Physics Today, 20(11):43–52, November 1967. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/

phtoad/v20/i11/p43_s1. Reprinted in [WP85, pages 272–281].

Frisch:1985:DF

[FW85] Otto R. Frisch and John A. Wheeler. The discovery of fission. In Weart and Phillips [WP85], pages 272–281. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [FW67].

Graetzer:1971:DNF

[GA71] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission: a documentary history, volume 20 of Van Nostrand Reinhold momentum books. Van Nostrand Reinhold, New York, NY, USA, 1971. viii + 120 pp. LCCN QC790 .G68.

Garrett:1963:DFP

[Gar63a] Alfred Benjamin Garrett. The discovery of the fission process: Lise Meitner and O. R. Frisch. In The flash of genius [Gar63b], pages 227–232. LCCN Q125 .G38.

Garrett:1963:FG

[Gar63b] Alfred Benjamin Garrett. The flash of genius. Van Nostrand, Princeton, NJ, USA, 1963. x + 249 pp. LCCN Q125 .G38.

Gavin:2004:BRP

[Gav04]  Patrick Gavin.  Book review: A physicists’ pantheon:

Great Physicists — The Life and Times of Leading Physicists from Galileo to Hawking by William H. Cropper. Oxford University Press, 2001. UK£24.99 (500 pages) ISBN 0-19-513748-5. Endeavour, 28(1):5, March 2004. CO-DEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0160932704000067.

Gavin:2005:BRP

[Gav05] Patrick Gavin. Book review: Physicists are human: Remarkable Physicists: from Galileo to Yukawa, Ioan James, Cambridge University Press (2004), £19.95 (389 pages, paperback), ISBN: 0-521-01706-8. Endeavour, 29(1):4–5, March 2005. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0160932704001322.

Gearhart:2014:FHE

[Gea14] Clayton A. Gearhart. The Franck–Hertz experiments, 1911–1914 experimentalists in search of a theory. Physics in Perspective (PIP), 16(3):293–343, September 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-014-0139-3.

Gispen:1997:BRL

[Gis97] Kees Gispen. Book review: Lise Meitner: A Life in Physics by Ruth Lewin Sime. Central European History, 30(3):449– 453, 1997. ISSN 0008-9389 (print), 1569-1616 (electronic). URL

http://www.jstor.org/stable/4546731.

Guerra:2012:DAR

[GLR12] Francesco Guerra, Matteo Leone, and Nadia Robotti. The discovery of artificial radioactivity. Physics in Perspective (PIP), 14(1):33–58, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0064-7;   http://www.
springerlink.com/content/v1733673628jr058/.

Gudden:1932:BGR

[GOB^{+}32] B. Gudden, Franz Ollendorff, Arn Berliner, E. Grossmann, Peter Pringsheim, H. Thiene, A. v. Brunn, and Lise Meitner. Besprechungen. (German) [Review]. Die Naturwissenschaften, 20 (44):806–812, October 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Goldschmidt:1987:PAF

[Gol87] Bertrand Goldschmidt. Pionniers de l’atome. (French) [Pioneers of the atom]. Stock, Paris, France, 1987. ISBN 2-234-02070-0. 484 + 8 pp. LCCN QC773 .G64 1987. URL http://gallica.

bnf.fr/ark:/12148/bpt6k4803658x. See also English transla-

tion [Gol90].

Goldschmidt:1990:AR

[Gol90] Bertrand Goldschmidt. Atomic Rivals. Rutgers University Press, New Brunswick, NJ, USA, 1990. ISBN 0-8135-1518-1 (hardcover), 0-8135-1519-X (paperback). xvii + 372 + 8 pp. LCCN QC773 .G6413 1990. English translation of [Gol87].

Goldberg:1996:BRF

[Gol96] Stanley Goldberg. Book review: With friends like these …: Lise Meitner: A Life in Physics, by Ruth Lewin Sime. Bulletin of the Atomic Scientists, 52(4):55–57, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:2001:BRP

[Gol01] Gary R. Goldstein. Book review: Patricia Rife. Lise Meitner and the Dawn of the Nuclear Age. Peace & Change, 26(1):95–105, January 2001. ISSN 0149-0508 (print), 1468-0130 (electronic).

Gould:1998:MSW

[Gou98] Paula A. Gould. Making space for women in the history of physics. Endeavour, 22(1):24–26, January 1998. CODEN EN-DEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Graetzer:1964:DNF

[Gra64] Hans G. Graetzer. Discovery of nuclear fission. American Journal of Physics, 32(9):9–15, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [HS39, Bey49, HS89].

Greulich:1984:HMI

[Gre84] W. Greulich. Hahn-Meitner-Institut für Kernforschung 25 Jahre alt. (German) [The Hahn–Meitner Institute for Nuclear Research is 25 years old]. Physik Journal, 40(8):282–283, August 1984. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Grieser:1991:KPW

[Gri91] Dietmar Grieser. Köpfe: Portraits der Wissenschaft; [das Buch zur Fernsehserie des ORF, (Folge 1–6)]. (German) [Heads: portraits of science; [the book for the TV series of the ORF, (episode 1–6)]]. ÖBV, Wien, Austria, 1991. ISBN 3-215-07723-X. 140 pp. LCCN Q149.A9 G75 1991x.

Gruss:2004:ZLM

[Gru04] Andrea Gruß. Zugehört: Lise Meitner — Die Frau in der Wissenschaft. Supposé, Köln 2003. 57 Min, Audio-CD, 18,- Euro. ISBN 3-932513-46-0. (German) [Listen: Lise Meitner — The woman in science. Supposè, Cologne 2003. 57 min, audio CD, 18,– Euro. ISBN 3-932513-46-0]. Nachrichten aus der Chemie, 52(6):711–712, June 2004. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic).

Haber:1979:WPS

[Hab79] Louis Haber. Women Pioneers of Science. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1979. ISBN 0-15-299202-2. xvii + 171 pp. LCCN Q141 .H215.

Hahn:1940:BAK

[Hah40] Otto Hahn. Bericht über die Arbeiten des Kaiser Wilhelm-Instituts für Chemie über “Präparat 38”. (German) [Report on the work of the Kaiser Wilhelm Institute for Chemistry on “Preparation 38”]. Report FA 002 514, Archives of the Deutsches Museum (ADM), Munich, Germany, December 10, 1940.

Hahn:1943:KAS

[Hah43] Otto Hahn. Künstliche atomumwandlungen und die spaltung des urans. (German) [Artificial atom transformations and the splitting of uranium]. Report FA 002 737, Archives of the Deutsches Museum (ADM), Munich, Germany, May 6, 1943.

Hahn:1962:EFU

[Hah62] Otto Hahn. Enrico Fermi and uranium fission. International Atomic Energy Agency Bulletin, 4(0):9–11, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL

http://www.iaea.org/Publications/Magazines/Bulletin/
Bull040su/04004700911su.pdf.

Hahn:1967:OHB

[Hah67] Otto Hahn. Otto Hahn in Berlin: Vor 60 Jahren begann die Zusammenarbeit mit Lise Meitner. (German) [Otto Hahn in Berlin: the collaboration with Lise Meitner began 60 years ago]. Physik Journal, 23(9):387–389, September 1967. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Hahn:1975:EEG

[Hah75] Otto Hahn. Erlebnisse und Erkenntnisse. (German) [Experiences and insights]. Econ-Verlag, Düsseldorf, West Germany, 1975. ISBN 3-430-13732-2. 320 pp. LCCN QD22.H2 A26. DM36.00. With an introduction by Karl-Erik Zimen. Edited by Dietrich Hahn.

Hahn:1979:OHB

[Hah79] Dietrich Hahn, editor. Otto Hahn: Begründer des Atomzeitalters: eine Biographie in Bildern und Dokumenten. (German) [Otto Hahn: founder of the atomic age: a biography in pictures and documents]. List Verlag, München, West Germany, 1979. ISBN 3-471-77841-1. 357 pp. LCCN QD22.H2 O87.

Hamilton:2002:LMP

[Ham02] Janet Hamilton. Lise Meitner: pioneer of nuclear fission. Great minds of science. Enslow Publishers, Berkeley Heights, NJ, USA, 2002. ISBN 0-7660-1756-7, 0-7660-1786-9. 128 pp. LCCN QC774 .M4 H35 2002.

Harteck:1943:BBB

[Har43] Paul Harteck. Bericht über die Besprechungen in Berlin am 27.9.43 und in Leuna am 28.9.43. (German) [Report on the meetings in Berlin on 27 September 1943 and in Leuna on 28 September 1943]. Report FA 002 796, Archives of the Deutsches Museum (ADM), Munich, Germany, October 1943.

Hargittai:2002:RSN

[Har02] István Hargittai. The road to Stockholm: Nobel Prizes, science, and scientists. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850912-X. xvii + 342 + 24 pp.

LCCN Q141 .H267 2002. URL http://www.loc.gov/catdir/
enhancements/fy0614/2002283888-d.html;   http://www.
loc.gov/catdir/enhancements/fy0724/2002283888-b.html;
 http://www.loc.gov/catdir/toc/fy033/2002283888.html.

Heisenberg:1969:GOH

[Hei69] Werner Heisenberg. Gedenkworte für Otto Hahn und Lise Meitner. (German) [Words of commemoration for Otto Hahn and Lise Meitner]. In ????, editor, Orden Pour le Mérite für Wissenschaften und Künste. Reden und Gedenkworte. (German) [Pour le Mérite for Arts and Sciences. Speeches and commemorative words], pages 111–119. Verlag Lambert Schneider, Heidelberg, West Germany, 1969. LCCN ????

Henning:2016:FDU

[Hen16] Eckart Henning. Freitod in Dahlem (1915): Unveröffentlichte Briefe von Edith Hahn und Lise Meitner über Dr. Clara Haber geb. Immerwahr. (German) [Suicide in Dahlem (1915): Unpublished letters of Edith Hahn and Lise Meitner on Dr. Ing. Clara Haber: always true]. Zeitschrift für anorganische und allgemeine Chemie, 642(6):432–436, March 2016. CODEN ZAACAB. ISSN 0044-2313 (print), 1521-3749 (electronic).

Hermann:1979:NPR

[Her79] Armin Hermann. The new physics: the route into the atomic age: in memory of Albert Einstein, Max von Laue, Otto Hahn, Lise Meitner. Inter Nationes, Bonn-Bad Godesberg, West Germany, 1979. ISBN 3-7879-0131-0 (paperback). 175 pp. LCCN QC773 .H4813.

Herneck:1980:SLM

[Her80] Friedrich Herneck. Über die Stellung von Lise Meitner und Otto Hahn in der Wissenschaftsgeschichte. (German) [On the position of Lise Meitner and Otto Hahn in the history of science]. Zeitschrift für Chemie (Stuttgart, Germany), 20(7):237– 243, July 1980. CODEN ZECEAL. ISSN 0044-2402. Reprinted in [Her10].

Hermann:1986:WAP

[Her86] Armin Hermann. Der Weg in das Atomzeitalter Physik wird Weltgeschichte. (German) [The path to the atomic age of physics becomes world history]. Moos, Gräfelfing bei München, Germany, second edition, 1986. ISBN 3-89164-021-8. 170 pp. LCCN QC7 .H473 1986.

Herrmann:2006:BLF

[Her06] Günter Herrmann. ,,Liebenswürdigkeit und fast unverwüstliche Heiterkeit… ”: Lise Meitner: Erinnerungen an Otto Hahn. Hrsg. von Dietrich Hahn. Hirzel, Stuttgart, 2005. 167 Seiten, brosch. 18,-Euro. ISBN 3-7776-1380-0. (German) [“Kindness and almost uncanny cheerfulness.”. Lise Meitner: Memories of Otto Hahn. Edited by Dietrich Hahn. Hirzel, Stuttgart, 2005. 167 pages, paperback 18 Euro. ISBN 3-7776-1380-0]. Nachrichten aus der Chemie, 54(12):1238, December 2006. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic).

Herneck:2010:SLM

[Her10] Friedrich Herneck. Über die Stellung von Lise Meitner und Otto Hahn in der Wissenschaftsgeschichte. (German) [On the position of Lise Meitner and Otto Hahn in the history of science]. Zeitschrift für Chemie (Stuttgart, Germany), 20(7):237–243, August 2010. CODEN ZECEAL. ISSN 0044-2402. Reprint of [Her80].

Hessenbruch:1996:BRL

[Hes96] Arne Hessenbruch. Book review: Lise Meitner —- A Life in Physics, ISBN 0-520-08906-5. European Journal of Physics, 17 (6):??, November 1996. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://stacks.iop.org/

0143-0807/17/i=6/a=021.

Hahn:1958:PEF

[HFK58] Otto Hahn, A. Flammersfeld, and W. Kroebel. Persönliche Erinnerungen an Frédéric Joliot/Lise Meitner 80 Jahre/Ernst Lamla 70 Jahre. (German) [Personal recollections of Frédéric Joliot /Lise Meitner’s 80th birthday /Ernst Lamla’s 70th birthday]. Physik Journal, 14(11):510–514, November 1958. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Hiebert:2000:BRL

[Hie00] Erwin N. Hiebert. Book review: Lise Meitner-Max von Laue Briefwechsel, 1938–1948 by Jost Lemmerich. Isis, 91(1):189– 190, March 2000. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
237626.

Harteck:1943:BTA

[HJ43] Paul Harteck and Johannes Jensen. Berechnung des Trenneffektes und der Ausbeute verschiedener Zentrifugenanordnungen zur Erho hung des Wirkungsgrades einer einzelnen Zentrifuge. (German) [Calculation of the separation effect and the yield of different centrifuge arrangements to increase the efficiency of a single centrifuge]. Report FA 002 592, Archives of the Deutsches Museum (ADM), Munich, Germany, February 1943.

Heinz:2017:EHS

[HJP17] Andreas Martin Heinz, Björn Jonson, and Imre Pázsit. EPS historic sites: Uddmanska House, Kungälv, Sweden. Europhysics News, 48(1):4–5, ???? 2017. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL https://www.

europhysicsnews.org/articles/epn/pdf/2017/01/epn2017-
48-1.pdf.

Hahn:1908:ACN

[HM08a] Otto Hahn and Lise Meitner. Actinium C, ein neues kurzlebiges Produkt des Actiniums. (German) [Actinium C, a new shortlived product of actinium]. Physikalische Zeitschrift, 9(20):649– 655, October 15, 1908. CODEN PHZTAO. ISSN 0369-982X.

URL https://hdl.handle.net/2027/mdp.39015068320947?
urlappend=;seq=677.

Hahn:1908:SAG

[HM08b] Otto Hahn and Lise Meitner. Über die β-Strahlen des Actiniums. (German) [On the β rays of actinium]. Physikalische Zeitschrift, 9(21):697–702, October 25, 1908. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/mdp.
39015068320947?urlappend=;seq=725.

Hahn:1908:ASE

[HM08c] Otto Hahn and Lise Meitner. Über die Absorption der β-Strahlen einiger Radioelemente. (German) [On the absorption of β rays of some radioelements]. Physikalische Zeitschrift, 9(10):321– 333, May 15, 1908. CODEN PHZTAO. ISSN 0369-982X.

URL https://hdl.handle.net/2027/mdp.39015068320947?
urlappend=;seq=347.

Hahn:1909:AUR

[HM09a] Otto Hahn and Lise Meitner. Die Ausstoßung radioaktiver Materie bei den Umwandlungen des Radiums. (German) [Exposure of radioactive matter by conversions of radium]. Physikalische Zeitschrift, 10(12):422, June 15, 1909. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/nyp.
33433090815253?urlappend=;seq=457. Comment on a work

by Sidney Ruß and Walter Makower [RM09].

Hahn:1909:NMH

[HM09b] Otto Hahn and Lise Meitner. Eine neue Methode zur Herstellung radioaktiver Zerfallsprodukte; Thorium D, ein kurzlebiges Produkt des Thoriums. (German) [A new method for producing radioactive decay products; Thorium D, a short lived thorium product]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 11(??):55–??, ???? 1909. CODEN VDPEAZ. ISSN 0372-5448.

Hahn:1909:NKN

[HM09c] Otto Hahn and Lise Meitner. Nachweis der komplexen Natur von Radium C. (German) [Evidence of the complex nature of Radium C]. Physikalische Zeitschrift, 10(20):697–703, October 15, 1909. CODEN PHZTAO. ISSN 0369-982X. URL https:

//hdl.handle.net/2027/nyp.33433090815253?urlappend=;

seq=740. See discussion [Mei09a].

Hahn:1909:ASG

[HM09d] Otto Hahn and Lise Meitner. Über das Absorptionsgesetz der β-Strahlen. (German) [On the absorption law of β rays]. Physikalische Zeitschrift, 10(24):948–950, December 1, 1909. CODEN PHZTAO. ISSN 0369-982X. URL https://hdl.handle.net/

2027/nyp.33433090815253?urlappend=;seq=1017.

Hahn:1909:TSE

[HM09e] Otto Hahn and Lise Meitner. Über eine typische β-Strahlung des eigentlichen Radiums. (German) [On a typical β radiation radiation]. Physikalische Zeitschrift, 10(21):741–745, November 1, 1909. CODEN PHZTAO. ISSN 0369-982X. URL https:

//hdl.handle.net/2027/nyp.33433090815253?urlappend=;

seq=792. See discussion [Mei09a].

Hahn:1910:NSB

[HM10] Otto Hahn and Lise Meitner. Eine neue β-Strahlung beim Thorium X; Analogien in der Uran- und Thoriumreihe. (German) [On new β radiation from Thorium X; analogies in the uranium and thorium series]. Physikalische Zeitschrift, 11(11): 493–497, June 1, 1910. CODEN PHZTAO. ISSN 0369-982X.

URL https://hdl.handle.net/2027/mdp.39015035421174?
urlappend=;seq=589.

Hahn:1912:GER

[HM12a] Otto Hahn and Lise Meitner. Grundlagen und Ergebnisse der radioaktiven Forschung. (German) [Fundamentals and results of radioactive research]. Forschungen und Fortschritte, 6(??):265– ??, ???? 1912. CODEN FOFOAI. ISSN 0367-2794.

Hahn:1912:VSE

[HM12b] Otto Hahn and Lise Meitner. Über die Verteilung der β-Strahlen auf die einzelnen Produkte des aktiven Niederschlags des Thoriums. (German) [On the distribution of β-rays on the individual products of active precipitation of thorium]. Physikalische Zeitschrift, 13(9):390–393, May 1, 1912. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/mdp.
39015023176806?urlappend=;seq=454.

Hahn:1913:UXG

[HM13a] Otto Hahn and Lise Meitner. Über das Uran X_{2}. (German) [On uranium X_{2}]. Physikalische Zeitschrift, 14(16):758– 759, August 15, 1913. CODEN PHZTAO. ISSN 0369-982X.

URL https://hdl.handle.net/2027/mdp.39015021268936?
urlappend=;seq=850.

Hahn:1913:FNK

[HM13b] Otto Hahn and Lise Meitner. Zur Frage nach der komplexen Natur des Radioactiniums und der Stellung des Actiniums im Periodischen System. (German) [On the question of the complex nature of radioactinium and the position of actinium in the periodic table]. Physikalische Zeitschrift, 14(16):752– 758, August 15, 1913. CODEN PHZTAO. ISSN 0369-982X.

URL https://hdl.handle.net/2027/mdp.39015021268936?
urlappend=;seq=844.

Hahn:1914:UGU

[HM14] Otto Hahn and Lise Meitner. Über das Uran Y. (German) [On uranium Y]. Physikalische Zeitschrift, 15(5):236–240, March 1, 1914. CODEN PHZTAO. ISSN 0369-982X. URL https://hdl.

handle.net/2027/uc1.c2831834?urlappend=;seq=268.

Hahn:1918:MAN

[HM18] Otto Hahn and Lise Meitner. Die Muttersubstanz des Actiniums, ein neues radioaktives Element von langer Lebensdauer. (German) [The parent substance of actinium, a new radioactive element of long life]. Physikalische Zeitschrift, 19(10):208–218, May 15, 1918. CODEN PHZTAO. ISSN 0369-982X. URL https://

hdl.handle.net/2027/hvd.hxgq92?urlappend=;seq=234.

Hahn:1919:UAG

[HM19a] Otto Hahn and Lise Meitner. Der Ursprung des Actiniums. (German) [The origin of actinium]. Physikalische Zeitschrift, 20(23): 529–??, December 1, 1919. CODEN PHZTAO. ISSN 0369-

982X. URL https://hdl.handle.net/2027/hvd.hxgq91?
urlappend=;seq=671.

Hahn:1919:PFN

[HM19b] Otto Hahn and Lise Meitner. Über das Protactinium und die Frage nach der Möglichkeit seiner Herstellung als chemisches Element. (German) [On protactinium and the question of the possibility of its production as a chemical element]. Die Naturwissenschaften, 7(33):611–612, August 1919. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1919:PLA

[HM19c] Otto Hahn and Lise Meitner. Über das Protactinium und die Lebensdauer des Actiniums. (German) [On protactinium and the ifetime of actinium]. Physikalische Zeitschrift, 20(6): 127–130, March 15, 1919. CODEN PHZTAO. ISSN 0369-

982X. URL https://hdl.handle.net/2027/hvd.hxgq91?
urlappend=;seq=175.

Hahn:1919:CEP

[HM19d] Otto Hahn and Lise Meitner. Über die chemischen Eigenschaften des Protaktiniums. I. Mitteilung Abscheidung des Protaktiniums aus Pechblende. (German) [On the chemical properties of protactinium. First communication. Deposition of protactinium from pitchblende]. Berichte der deutschen chemischen Gesellschaft (Abteilung A und B), 52(9):1812–1828, October 1919. CODEN BDCBAD. ISSN 1099-0682 (print), 1434-1948 (electronic).

Meitner:1920:PSL

[HM20a] Otto Hahn and Lise Meitner. Protactinium, seine Lebensdauer und sein Gehalt in Uran-Mineralien. (German) [Protactinium, its life and content in uranium minerals]. Physikalische Zeitschrift, 21(21–22):591–592, November 1–15, 1920. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/mdp.
39015086723080?urlappend=;seq=657.

Hahn:1920:AVG

[HM20b] Otto Hahn and Lise Meitner. Über die Anwendung der Verschiebungsregel auf gleichzeitig α- und β-Strahlen aussendende Substanzen. (German) [On application of the shift rule to substances emitting α and β rays simultaneously]. Zeitschrift für Physik, 2(1):60–70, February 1920. CODEN ZEPYAA. ISSN 0044-3328.

Hahn:1920:EPG

[HM20c] Otto Hahn and Lise Meitner. Über die Eigenschaften des Protactiniums. (German) [On the properties of protactinium]. Berichte der deutschen chemischen Gesellschaft (Abteilung A und B), 54 (1):202–??, ???? 1920. CODEN BDCBAD. ISSN 1099-0682 (print), 1434-1948 (electronic).

Hahn:1921:ARS

[HM21a] Otto Hahn and Lise Meitner. Über das Arbeiten mit radioaktiven Substanzen. (German) [On working with radioactive substances]. Die Naturwissenschaften, 9(18):316–318, May 1921. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1921:EPI

[HM21b] Otto Hahn and Lise Meitner. Über die Eigenschaften des Protaktiniums, II. Mitteilung: Seine Lebensdauer und sein Gehalt in Uranmineralien. (German) [On properties of protactinium, Second communication: its lifetime and its content in uranium minerals]. Berichte der deutschen chemischen Gesellschaft (Abteilung A und B), 54(1):69–77, January 1921. CODEN BD-CBAD. ISSN 1099-0682 (print), 1434-1948 (electronic).

Hahn:1922:NPP

[HM22] Otto Hahn and Lise Meitner. Notiz über den Protactiniumgehalt in Pechblenderückständen und das Abzweigungsverhältnis der Actiniumreihe. (German) [Note of the protactinium content in pitchblue residues and the branching ratio of the actinium series]. Zeitschrift für Physik, 8(1):202–204, December 1922. CODEN ZEPYAA. ISSN 0044-3328.

Hahn:1923:SUX

[HM23] Otto Hahn and Lise Meitner. Die γ-Strahlen von Uran X und ihre Zuordnung zu Uran X_{1} und Uran X_{2}. (German) [The γ rays of uranium X and their allocation to uranium X_{1} and uranium X_{2}]. Zeitschrift für Physik, 17(1):157–167, December 1923. CODEN ZEPYAA. ISSN 0044-3328.

Hahn:1924:SRS

[HM24] Otto Hahn and Lise Meitner. Das β-Strahlenspektrum von Radium und seine Deutung. (German) [The β radiation spectrum of radium and its interpretation]. Zeitschrift für Physik, 26(1): 161–168, December 1924. CODEN ZEPYAA. ISSN 0044-3328.

Hahn:1925:BAK

[HM25a] Otto Hahn and Lise Meitner. Bemerkungen zu einer Arbeit über die künstliche Umwandlung von Uran in Uran X. (German) [Comments on a work on the artificial conversion of uranium into uranium X]. Die Naturwissenschaften, 13(44):907, October 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1925:SRS

[HM25b] Otto Hahn and Lise Meitner. Die β-Strahlspektren von Radioactinium und seinen Zerfallsprodukten. (German) [The β radiation spectra of radioactinium and its decay products]. Zeitschrift für Physik, 34(1):795–806, December 1925. CODEN ZEPYAA. ISSN 0044-3328.

Hahn:1931:LRS

[HM31a] Otto Hahn and Lise Meitner. Lord Rutherford zum sechzigsten Geburtstag. (German) [Lord Rutherford’s sixtieth birthday]. Die Naturwissenschaften, 19(35):729, August 1931. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01522345.

Hahn:1931:NEP

[HM31b] Otto Hahn and Lise Meitner. Notiz über die Entdeckung des Protactiniums. Die Naturwissenschaften, 19(35):738, August 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1932:BOM

[HM32] Otto Hahn and Lise Meitner. Bemerkungen zu obiger Mitteilung: Zu A. v. Grosse: Zur Entdeckung und Isolierung der Elements 91. (German) [Remarks on the above communication: A. v. Grosse: On the discovery and isolation of element 91]. Die Naturwissenschaften, 20(21):363, May 20, 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL https://link.springer.com/article/10.1007/

BF01504929. See [vG32].

Hahn:1933:EBG

[HM33] Otto Hahn and Lise Meitner. Zur Entstehungsgeschichte der Bleiarten. (German) [On the genesis of lead species [isotopes]]. Die Naturwissenschaften, 21(12):237–238, March 1933. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1935:KUT

[HM35a] Otto Hahn and Lise Meitner. Die künstliche Umwandlung des Thoriums durch Neutronen: Bildung der bisher fehlenden radioaktiven 4n+1 Reihe. (German) [The artificial transformation of thorium by neutrons: Formation of the missing radioactive 4n + 1 series]. Die Naturwissenschaften, 23(20):320–??, May 17, 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1935:KUUa

[HM35b] Otto Hahn and Lise Meitner. Über die künstliche Umwandlung des Urans durch Neutronen. (German) [On the artificial transformation of uranium by neutrons]. Die Naturwissenschaften, 23 (2):37–38, January 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1935:KUUb

[HM35c] Otto Hahn and Lise Meitner. Über die künstliche Umwandlung des Urans durch Neutronen (II. Mitteil.). (German) [On the artificial transformation of uranium by neutrons (second communication)]. Die Naturwissenschaften, 23(14):230–231, April 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1936:KRA

[HM36] Otto Hahn and Lise Meitner. Künstliche radioaktive Atomarten aus Uran und Thor. (German) [Artificial nuclear radioactive species from uranium and thorium]. Zeitschrift für Angewandte Chemie, 49(12–13):127–128, December 1936. ISSN 0932-2132.

Hahn:1937:TUK

[HM37] Otto Hahn and Lise Meitner. Trans-Urane als künstliche radioaktive Umwandlungsprodukte des Urans. (German) [Transuraniums as artificial radioactive transformation products of uranium]. Forschungen und Fortschritte, 13(??):298–??, ???? 1937. CODEN FOFOAI. ISSN 0367-2794.

Hahn:1938:TUK

[HM38] Otto Hahn and Lise Meitner. Trans-Urane als künstliche radioaktive Umwandlungsprodukte des Urans. (German) [Transuraniums as artificial radioactive transformation products of uranium]. Scientia (Milan), 63(??):13–??, January 1938. CODEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic).

Hahn:1954:AFG

[HM54] Otto Hahn and Lise Meitner. Atomenergie und Frieden. (German) [Atomic energy and peace]. Technical report, United Nations International Atomic Energy Agency, Vienna, Austria, 1954.

Honigmann:1920:BGR

[HMB20] H. L. Honigmann, L. Meitner, and A. Berliner. Besprechungen. (German) [Review]. Die Naturwissenschaften, 8(26):512– 514, June 1920. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1934:TIC

[HMS34]  Otto Hahn, Lise Meitner, and Fritz Straßmann.  Die

Transurane und ihre chemischen Verhaltensweisen. (German) [The transuranic and their chemical behaviors]. Berichte der deutschen chemischen Gesellschaft (Abteilung A und B), 67(??): ??, ???? 1934. CODEN BDCBAD. ISSN 1099-0682 (print), 1434-1948 (electronic).

Hahn:1935:EWB

[HMS35] Otto Hahn, Lise Meitner, and Fritz Strassmann. Einige weitere Bemerkungen ber die künstlichen Umwandlungsprodukte beim Uran. (German) [Some more remarks about the artificial conversion products of uranium]. Die Naturwissenschaften, 23(31): 544–545, August 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hahn:1936:NUP

[HMS36] Otto Hahn, Lise Meitner, and Fritz Straßmann. Neue Umwandlungs-Prozesse bei Neutronen-Bestrahlung des Urans; Elemente jenseits Uran. (German) [New conversion processes in neutron irradiation of uranium; elements beyond uranium]. Berichte der deutschen chemischen Gesellschaft (Abteilung A und B), 69(4):905–919, April 1936. CODEN BDCBAD. ISSN 1099-0682 (print), 1434-1948 (electronic).

Hahn:1937:TUI

[HMS37] O. Hahn, L. Meitner, and F. Strassmann. Über die Trans-Urane und ihr chemisches Verhalten. (German) [On transuranium and its chemical behavior]. Berichte der deutschen chemischen Gesellschaft (Abteilung A und B), 70(6):1374–1392, June 1937. CODEN BDCBAD. ISSN 1099-0682 (print), 1434-1948 (electronic).

Hahn:1938:NLU

[HMS38] O. Hahn, L. Meitner, and F. Strassmann. Ein neues langlebiges Umwandlungsprodukt in den Trans-Uranreihen. (German) [A new durable transformation product in the transuranium series]. Die Naturwissenschaften, 26(29):475–476, July 1938. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hoffmann:1978:OHS

[Hof78] Klaus Hoffmann. Otto Hahn: Stationen aus dem Leben eines Atomforschers: Biografie. (German) [Otto Hahn: Stations in the life of a nuclear scientist: Biography]. Verlag Neues Leben, Berlin, West Germany, 1978. ISBN 3-355-00561-4. 374 pp. LCCN QD22.H2 H63.

Hoffmann:1997:LML

[Hof97] Dieter Hoffmann. Lise Meitner. A Life in Physics. Von Ruth Lewin Sime, University of California Press, Berkeley 1996. 528 Seiten, zahlreiche Bilder, $34,95. Physik in unserer Zeit, 28(1): 47, 1997. CODEN PHUZAH. ISSN 0031-9252.

Hoffmann:2005:FVO

[Hof05] Klaus Hoffmann. Forschung und Verantwortung — Otto Hahn: Konflikte eines Wissenschaftlers. (German) [Research and responsibility — Otto Hahn: Conflicts of a scientist], volume 6 of Mäzene, Stifter, Stadtkultur: Schriften der Frankfurter Bürgerstiftung und der Ernst Max von Grunelius-Stiftung. Kramer, Germany, Frankfurt am Main, 2005. ISBN 3-7829-0560-1 (paperback). 240 pp. LCCN ????

Hoffmann:2023:OEF

[Hof23] Dieter Hoffmann. Operation Epsilon: die Farm-Hall-Protokolle erstmals vollständig, ergänzt um zeitgenössische Briefe und weitere Dokumente der 1945 in England internierten deutschen Atomforscher. (German) [Operation Epsilon: the Farm Hall Protocols complete for the first time, supplemented by contemporary letters and other documents from the German nuclear researchers interned in England in 1945]. GNT-Verlag, Berlin, Germany, 2023. ISBN 3-86225-111-X (hardcover), 3-86225-508-5, 3-87134-082-0, 978-38622-5-1. 588 pp. LCCN QC773.A1.

Hook:2002:PSD

[Hoo02] Ernest B. Hook, editor. Prematurity in Scientific Discovery: on Resistance and Neglect. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 2002. ISBN 0-520-23106-6 (hardcover), 0-520-23107-4 (paper), 0-520-92773-7 (e-book), 0-585-44116-2 (e-book), 1-59734-830-9, 1-282-35672-0. xx + 378 pp. LCCN Q180.55.D57 P74

2002. URL http://muse.jhu.edu/books/9780520927735/;
http://www.jstor.org/stable/10.1525/j.ctt1pnwvf.

Hook:2003:GBI

[Hoo03] Ernest B. Hook. Gender bias and Ida Noddack. Science, 301 (5636):1045, August 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See response [Ore03].

Hermann:2001:MPV

[HR01] Armin Hermann and Hans Roos. Max Planck: Vorträge, Reden, Erinnerungen. (German) [Max Planck: Lectures, Speeches, Recollections]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-41274-3 (hardcover). viii + 224 pp. LCCN QC3 .P65 2001. URL http:

//link.springer.com/book/10.1007/978-3-642-56594-6;
http://www.gbv.de/dms/bowker/toc/9783540412748.pdf;
http://www.gbv.de/dms/fazrez/FNUW200109261084460.pdf;
 http://www.zentralblattmath.org/zmath/en/search/?an=
0978.01057.

Hahn:1939:NVB

[HS39] Otto Hahn and Fritz Strassmann. Über den Nachweis und das Verhalten der bei der Bestrahlung des Urans mittels Neutronen entstehenden Erdalkalimetalle. (German) [Concerning the evidence and existence of alkaline earth metals resulting from the neutron irradiation of uranium]. Die Naturwissenschaften, 27 (1):11–15, January 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A facsimile is also available in [Bey49, pages 87–91] and in [Gra64]. Abridged English translation in [GA71, pages 44–47].

Hahn:1989:PFA

[HS89] Otto Hahn and Fritz Strassmann. Proof of the formation of active isotopes of barium from uranium and thorium irradiated with neutrons; proof of the existence of more active fragments produced by uranium fission. Journal of Chemical Education, 66(5):362–363, May 1989. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic). URL http://pubs.acs.org/ doi/abs/10.1021/ed066p362. English translation of [HS39]. Special issue commemorating fify years of nuclear fission. See also [Gra64].

Hahn:1936:KWI

[HSMH36] Otto Hahn, ?. Stock, L. Meitner, and K. Hess. Kaiser Wilhelm-Institut für Chemie in Berlin-Dahlem. (German) [Kaiser Wilhelm Institute for Chemistry in Berlin-Dahlem]. In 25 Jahre Kaiser Wilhelm-Gesellschaft zur Förderung der Wissenschaften. (German) [25 Years of the Kaiser Wilhelm Society for the Advancement of Science], pages 164–175. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1936. ISBN 3-642-47098-X (print), 3-642-47338-5 (e-book). LCCN ????

James:2004:LM

[Jam04a] I. M. (Ioan Mackenzie) James. Lise Meitner (1878–1968). In Remarkable physicists: from Galileo to Yukawa [Jam04b], page ?? ISBN 0-521-81687-4, 0-521-01706-8 (paperback). LCCN QC15

.J36 2004. URL http://www.loc.gov/catdir/description/
cam032/2003055423.html;   http://www.loc.gov/catdir/
enhancements/fy0732/2003055423-b.html;   http://www.
loc.gov/catdir/samples/cam041/2003055423.html;   http:
//www.loc.gov/catdir/toc/cam032/2003055423.html.

James:2004:RPG

[Jam04b] I. M. (Ioan Mackenzie) James. Remarkable physicists: from Galileo to Yukawa. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-81687-4, 0-521-01706-8 (paperback). xv + 389 pp. LCCN QC15 .J36 2004. URL http://www.loc.gov/

catdir/description/cam032/2003055423.html;   http://
www.loc.gov/catdir/enhancements/fy0732/2003055423-b.
html;   http://www.loc.gov/catdir/samples/cam041/2003055423.
html;   http://www.loc.gov/catdir/toc/cam032/2003055423.
html.

Jensen:2000:CCN

[Jen00] Carsten Jensen. Controversy and Consensus: Nuclear Beta Decay 1911–1934, volume 24 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 3-0348-9569-0 (paperback), 3-7643-5313-9 (hardcover), 3-0348-8444-3 (e-book). xv + 217 pp. LCCN

QC793.5.B425 J46 2000. URL http://www.springerlink.com/

content/978-3-0348-8444-0. Carsten Jensen died of cancer a few months after presenting his doctoral dissertation in 1990 at the University of Copenhagen. Finn Aaserud, Helge Kragh, Erik Rüdinger, and Roger H. Stuewer produced this book as a slightly edited version of that work, supplying additional figures, but leaving the prose largely untouched.

Jachimowicz:2015:CLL

[JKS15] P. Jachimowicz, M. Kowal, and J. Skalski. Candidates for longlived high-K ground states in superheavy nuclei. Physical Review C (Nuclear Physics), 92(4), October 2015. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Johnstone:2005:EBI

[JMH^{+}05] Gary Johnstone, Aidan McArdle, Shirley Henderson, John Lithgow, and David Bodanis. Einstein’s big idea. WGBH Boston Video, Boston, MA, USA, letterbox edition, 2005. ISBN 1-59375-317-9. LCCN QC173.6 .N68 2005. 1 videodisc (ca. 112 min.).

Johnson:1996:BRP

[Joh96] Jeffrey A. Johnson. Book review: Posthumous credit: Lise Meitner: A Life in Physics, by Ruth Lewin Sime. Science, 272(5254): 42, April 5, 1996. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
2890763.

Johnsen:2017:AB

[Joh17] Julia E. Johnsen, editor. The atomic bomb. Deutsche Nationalbibliothek, Leipzig and Frankfurt am Main, Germany, 2017. 335 pp.

Jones:1990:ICW

[Jon90] L. M. Jones. Intellectual contributions of women to physics: Lise Meitner and Ida Noddack. In Kass-Simon et al. [KSFN90], pages 193–195. ISBN 0-253-20813-0 (paperback), 0-253-33264-8. LCCN Q130 .W67 1990.

Kacser:1996:LMM

[Kac96] Claude Kacser. Lise Meitner, Manne Siegbahn, and Adolf Hitler. American Journal of Physics, 63(2):106–107, February 1996. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/63/2/10.1119/1.17992. See [Sim94b] and rebut-

tal [Tru96].

Kaempffert:1939:WSW

[Kae39] Waldemar Kaempffert. This week in science: When uranium splits: Doubtful source of power; cancer and x-rays; neutron possibilities. New York Times, ??(??):D9, March 5, 1939. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/102937542.

Kaempffert:1945:SSB

[Kae45] Waldemar Kaempffert. Story of scientists’ ‘battle’ for atom bomb secret revealed in Smyth Report. New York Times, ??(??):8, August 16, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107074385/.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kant:2012:LMS

[Kan12] Horst Kant. Lise Meitner and the (supposed) transuranic elements. Annalen der Physik (Berlin), 524(6–7):A99–A102, July 2012. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Karlik:1979:MLM

[Kar79] Berta Karlik. In memoriam Lise Meitner. Physik Journal, 35 (2):49–52, February 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Keiser:2014:ZAE

[Kei14] Vera Keiser. Die zeitliche Abfolge der Erkenntnisse im Prozess der Entdeckung der Atomkernspaltung. (German) [The timing of the findings in the process of the discovery of nuclear atomic fission], June 15, 2014. URL http://www.

sf-project.de/morgenweck/downloads/die%2520zeitliche%
2520abfolge%2520der%2520erkenntnisse%2520im%2520prozess%
2520der%2520entdeckung%2520der%2520atomkernspaltung.
pdf.

Keller:1996:BRN

[Kel96] Evelyn Fox Keller. Book review: In and out of the nuclear family: Lise Meitner: A Life in Physics by Ruth Lewin Sime. The Women’s Review of Books, 13(12):5–6, September 1996.

Kenworthy:1954:DHB

[Ken54] E. W. Kenworthy. The drama of the hydrogen bomb — and Dr. Oppenheimer’s key role: Security case focuses attention on disputes that preceded first successful test of Hbomb at Pacific Proving Ground. New York Times, ??(??):E5, April 18, 1954. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/113104238/.

Kerner:1991:WMG

[Ker91] Charlotte Kerner. “Wird es mir gelingen, eine Wissenschaftlerin zu werden?” Ein Kurzporträt der Atomphysikerin Lise Meitner (1878–1968). (German) [“Will I be able to become a scientist?” A short portrait of the atomic physicist Lise Meitner (1878–1968)]. In Pionierinnen — Feministinnen — Karrierefrauen?: 90 Jahre Frauenstudium in Deutschland. (German) [Pioneers — feminists — career women?: 90 years of studying women in Germany] [Sch91], pages 105–114. ISBN 3-89085-419-2. LCCN HQ1181.G3 P65 1992.

Kerner:1998:LAL

[Ker98] Charlotte Kerner. Lise, Atomphysikerin: die Lebensgeschichte der Lise Meitner. (German) [Lise, nuclear physicist: the life story of Lise Meitner], volume 812 of Gulliver-Taschenbuch. Beltz und Gelberg, Weinheim, Germany, 1998. ISBN 3-407-78812-6, 3-407-80847-X. 197 + 8 pp. LCCN ????

Kraus:2001:UGE

[KF01] Elisabeth Kraus and Carl Joachim Friedrich. Von der Uranspaltung zur Göttinger Erklärung: Otto Hahn, Werner Heisenberg, Carl Friedrich von Weizsäcker und die Verantwortung des Wissenschaftlers. (German) [From uranium fission to the Göttingen Declaration: Otto Hahn, Werner Heisenberg, Carl Friedrich von Weizsäcker and the responsibility of the scientist]. Königshausen and Neumann, Würzburg, Germany, 2001. ISBN 3-8260-1987-3. 449 pp. LCCN ????

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kauffman:1997:BRLb

[KK97a] George B. Kauffman and Laurie M. Kauffman. Book review: Lise Meitner: A Life in Physics. (Serie: California Studies in the History of Science. Vol. 13. Serienherausgeber J. L. Heilbron.) Von R. L. Sime. University of California Press, Berkeley, 1996. XIII, 526 S., geb. 34.95$. — ISBN 0-520-08906-5. Angewandte Chemie, 109(15):1729–1730, August 1997. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Kauffman:1997:BRLa

[KK97b] George B. Kauffman and Laurie M. Kauffman. Book review: Lise Meitner: A Life in Physics. (Series: California Studies in the History of Science, Volume 13. Series editor: J. L. Heilbron.) By R. L. Sime. Angewandte Chemie, International Edition in English, 36(6):651–652, April 1997. CODEN ACIEAY. ISSN 0570-0833.

Katzir:2013:TTH

[KLR13] Shaul Katzir, Christoph Lehner, and Jürgen Renn, editors. Traditions and transformations in the history of quantum physics: HQ-3, Third International Conference on the History of Quantum Physics, Berlin, June 28–July 2, 2010, volume 5 of Max Planck research library for the history and development of knowledge. Proceedings. Edition Open Access, Berlin, Germany, 2013.

ISBN 3-8442-5134-0. LCCN QC173.98. URL http://www.

edition-open-access.de/proceedings/5/.

Krafft:1978:LMH

[Kra78a] Fritz Krafft. Lise Meitner: Her life and times — on the centenary of the great scientist’s birth. Angewandte Chemie, International Edition in English, 17(11):826–842, November 1978. CODEN ACIEAY. ISSN 0570-0833.

Krafft:1978:LMI

[Kra78b] Fritz Krafft. Lise Meitner und ihre Zeit — Zum hundertsten Geburtstag der bedeutenden Naturwissenschaftlerin. (German) [Lise Meitner and her time — on the centenary of the great physicist]. Angewandte Chemie, 90(11):876–892, November 1978. CO-DEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Krafft:1981:ISSd

[Kra81a] Fritz Krafft. Im Schatten der Sensation: Leben und Wirken von Fritz Strassmann. (German) [In the Shadow of Sensation: Life and Work of Fritz Strassmann]. Verlag Chemie, Weinheim, West Germany, 1981. ISBN 3-527-25818-3. xvii + 541 pp. LCCN QD22.S77 K7.

Krafft:1981:ISS

[Kra81b] Fritz Krafft. Im Schatten der Sensation: Leben und Wirken von Fritz Strassmann. (German) [In the shadow of sensation: the life and work of Fritz Strassmann]. Verlag Chemie, Weinheim, West Germany, 1981. ISBN 3-527-25818-3. xvii + 541 pp. LCCN QD22.S77 K7.

Krafft:1981:NTH

[Kra81c] Fritz Krafft. Nochmals zum Team Hahn-Meitner-Straßmann: Abschließende Bemerkungen zu der Gegendarstellung ,,Otto Hahn und Fritz Straßmann” von Dietrich Hahn, München, in Phys Bl. 37 (1981) Nr. 2, Seite 44–46. (German) [More on the Hahn–Meitner–Strassmann team: Concluding remarks on the reply ,,Otto Hahn and Fritz Strassmann ” by Dietrich Hahn, Munich, in Phys Bl. 37 (1981) no. 2, pages 44–46]. Physik Journal, 37(11):351–352, November 1981. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Krafft:1983:IEC

[Kra83] Fritz Krafft. Internal and external conditions for the discovery of fission by the Berlin Team. In Shea [She83], chapter 5, pages 135–165. ISBN 90-277-1584-X. LCCN QC773 .O87

1983. URL https://link.springer.com/chapter/10.1007/
978-94-009-7133-2_5.

Krafft:1988:LMB

[Kra88] Fritz Krafft. Lise Meitner: eine Biographie, volume 448 of HMI-B, 0175-8349. Hans-Meitner-Institut, Berlin, Germany, 1988. 24 pp. LCCN MLCS 88/06254 (Q).

Kragh:2000:LMN

[Kra00] Helge Kragh. Lise Meitner: a nuclear heroine? Physics World, 13(3):45–46, March 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.org/

2058-7058/13/i=3/a=31. See comment [Brü00].

Kragh:2002:BRP

[Kra02] Helge Kragh. Book review: The Politics of Excellence: Behind the Nobel Prize in Science, Robert Marc Friedman. W. H. Freeman, New York, 2001. $30.00 (379 pp.) ISBN 0-7167-3103-7. Physics Today, 55(3):63, March 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Krim:1999:BRA

[Kri99] Jacqueline Krim. Book review: Against all odds: The triumphs and defeats of Lise Meitner: Lise Meitner: A Life in Physics, by Ruth Lewin Sime. The Physics Teacher, 37(4):202, April 1999. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928

(electronic). URL http://scitation.aip.org/content/aapt/
journal/tpt/37/4/10.1119/1.880229.

Kass-Simon:1990:WSR

[KSFN90] G. (Gabriele) Kass-Simon, Patricia Farnes, and Deborah Nash, editors. Women of Science: Righting the Record. Indiana University Press, Bloomington, IN, USA, 1990. ISBN 0-253-20813-0 (paperback), 0-253-33264-8. xvi + 398 pp. LCCN Q130 .W67

Kurian:2002:NSB

[Kur02] George Thomas Kurian. The Nobel scientists: a biographical encyclopedia. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-57392-927-1. 420 pp. LCCN Q141 .K78 2002.

Krafft:1988:KLA

[KZN^{+}88] F. Krafft, H.-D. Zeh, H. P. Nilles, J. A. Maruhn, H. Rechenberg, F. Pobell, and A. Hofzumahaus. Kerner: Lise, Atomphysikerin. Die Lebensgeschichte der Lise Meitner/Hawking: Eine kurze Geschichte der Zeit: Die Suche nach der Urkraft des Universums/Davies u. Brown: Superstrings: A Theory of Everything?/ Hasse u. Myers: Geometrical Relationships of Macroscopic Nuclear Physics/Thorsen: Niels Bohr Collected Works, Vol. 8/ Richardson u. Smith: Experimental Techniques in Condensed Matter Physics at Low Temperatures/Jaenicke: Ergebnisse aus dem gleichnamigen Sonderforschungsbereich. (German) [Kerner: Lise, nuclear physicist. The Life Story of Lise Meitner / Hawking: A Brief History of Time: The Quest for the Primal Power of the Universe/ Davies & Brown: Superstrings: A Theory of Everything?. Myers: Geometrical Relationships of Macroscopic Nuclear Physics/Thorsen: Niels Bohr Collected Works, Vol. 8/ Richardson & Smith: Experimental Techniques in Condensed Matter Physics at Low Temperatures/Jaenicke: Results from the eponymous special research area]. Physik Journal, 44(12):466– 468, December 1988. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

LAnnunziata:2007:RIH

[L’A07] Michael F. L’Annunziata. Radioactivity: introduction and history. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-52715-X (hardcover). xxi + 609 pp. LCCN QC795 .L264 2007.

URL http://www.loc.gov/catdir/toc/fy0805/2008925731.

html. Foreword by Werner Burkart (Deputy Director General, International Atomic Energy Agency, Vienna).

Lahmer:1991:HMI

[Lah91] W. Lahmer. Hahn–meitner-institut im aufbruch zu neuen ufern: Der neue geschäftsführer des hahn-meitner-instituts im gespräch mit den Physikalischen Blättern. (German) [Hahn–Meitner Institute on the move to new shores: the new managing director of the Hahn–Meitner Institute in conversation with the Physical Letters]. Physik Journal, 47(11):1011–1012, November 1991. CO-DEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Lammert:1979:WZE

[Lam79] Eberhard Lammert. Der Wissenschaftler zwischen Erkenntnis und Verantwortung: Zur Feier der 100. Geburtstage von Albert Einstein, Otto Hahn, Lise Meitner und Max von Laue am 1. März 1979 in Berlin. (German) [The scientist between knowledge and responsibility: Celebrating the 100th birthdays of Albert Einstein, Otto Hahn, Lise Meitner and Max von Laue on March 1, 1979 in Berlin]. Physik Journal, 35(11):526–530, November 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Laurence:1951:DHR

[Lau51] William L. Laurence. Day of Hiroshima recalls atom race: Two leaders in atomic work at Columbia University. New York Times, ??(??):3, August 6, 1951. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/112188459/fulltextPDF.

Lazarus:1982:DF

[Laz82] Allan M. Lazarus. Discoveries of fission. Physics Today, 35(11): 106–108, November 1982. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Lee:2000:SGR

[Lee00] David Lee. Secretary General’s report. Europhysics News, 31 (1):21, ???? 2000. CODEN EUPNAS. ISSN 0531-7479 (print),

1432-1092 (electronic). URL https://www.europhysicsnews.
org/articles/epn/pdf/2000/01/epn2000-31-1.pdf.

Leicester:1968:SBC

[Lei68] Henry M. Leicester. Source Book in Chemistry, 1900–1950. Harvard University Press, Cambridge, MA, USA, 1968. ISBN 0-674-36669-7, 0-674-36670-0. xvii + 408 pp. LCCN QD3.

Leino:2009:CNL

[Lei09a] Matti Leino. Call for nominations for the 2010 Lise Meitner Prize for Nuclear Science. Europhysics News, 40(5):7, ???? 2009. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL https://www.europhysicsnews.org/articles/
epn/pdf/2009/05/epn_40-5.pdf.

Leino:2009:LMP

[Lei09b] Matti Leino. Lise Meitner Prize: The European Physical Society awarded, through its Nuclear Physics Board, the Lise Meitner Prize for Nuclear Science 2008 to professor Walter Greiner and professor Reinhard Stock. Europhysics News, 40(2):8, ???? 2009. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL https://www.europhysicsnews.org/articles/
epn/pdf/2009/02/epn2009-40-2.pdf.

Lemmerich:1998:LMM

[Lem98] Jost Lemmerich. Lise Meitner–Max von Laue: Briefwechsel 1938–1948. (German) [Lise Meitner–Max von Laue: correspondence 1938–1948], volume 22 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. ERS-Verlag1998, Berlin, Germany, 1998. ISBN 3-928577-32-8 (paperback). 560 pp. LCCN QC16.M45 A3 1998; QC774.M4 A4 1998.

Lemmerich:2003:LMG

[Lem03] Jost Lemmerich. Lise Meitner zum 125. Geburtstag: Ausstellungskatalog: Staatsbibliothek zu Berlin Preußischer Kulturbesitz, 7. November–13. Dezember 2003. (German) [Lise Meitner’s 125th Birthday: Exhibition Catalog: Staatsbibliothek zu Berlin Prussian Cultural Heritage, November 7–13. December 2003]. ERS-Verlag, Berlin, Germany, 2003. ISBN 3-928577-51-4. 150 pp. LCCN QC774.M4 L57 2003.

Lemmerich:2010:BFL

[Lem10] Jost Lemmerich. Bande der Freundschaft: Lise Meitner– Elisabeth Schiemann, Kommentierter Briefwechsel 1911–1947. (German) [Bond of Friendship: Lise Meitner–Elisabeth Schiemann, Annotated Correspondence 1911–1947], volume 61 of Österreichische Akademie der Wissenschaften. Verlag der Österreichischen Akademie der Wissenschaften, Wien, Austria, 2010. ISBN 3-7001-6847-0 (paperback). 368 pp. LCCN QC16.M45 A4 2010.

Levi:1997:ESN

[Lev97] Barbara Goss Levi. Elements 104–109: Soon to be named? Physics Today, 50(5):52–53, May 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Lightman:2005:DGB

[Lig05] Alan P. Lightman. The discoveries: great breakthroughs in twentieth-century science. Pantheon, New York, NY, USA, 2005. ISBN 0-676-97789-8, 0-375-42168-8 (hardcover). xviii + 553 + 16 pp. LCCN Q180.55.D57 .L53. URL http://www.loc.gov/

catdir/enhancements/fy0623/2005040854-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0623/2005040854-
d.html;   http://www.loc.gov/catdir/enhancements/fy0623/
2005040854-t.html.

Lindsay:1960:BRT

[Lin60] R. Bruce Lindsay. Book review: Trends in Atomic Physics: essays dedicated to Lise Meitner, Otto Hahn, Max von Laue on the occasion of their 80th birthday. Edited by O. R. Frisch, F. A. Paneth, F. Laves, P. Rosbaud. 285 pp. (Vieweg & Sohn, Germany) Interscience Publishers, Inc., New York, 1959. $7.50. Physics Today, 13(9):45, September 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/scitation.aip.org/content/aip/magazine/physicstoday/
article/13/9/10.1063/1.3057115.

Lipson:1959:BRB

[Lip59] H. Lipson. Book review: Beiträge zur Physik und Chemie des 20 Jahrhunderts, by Otto Robert Frisch, Friedrich Adolf Paneth, Friedrich Laves and Paul Rosbaud. Physics Bulletin, 10(12):313, December 1959. CODEN PHSBB4. ISSN 0031-9112. URL http:

//stacks.iop.org/0031-9112/10/i=12/a=016.

Leibfried:2015:BWE

[LMOS15] Stephan Leibfried, Christoph Markschies, Ernst Osterkamp, and Günter Stock. Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy]. De Gruyter Akademie Forschung, Berlin, Germany, 2015. ISBN 3-11-037598-2 (hardcover). 517 pp. LCCN AS182.B38 B477 2015.

Lemaitre:2015:NSS

[LPS^{+}15] Jean-François Lemaître, Stefano Panebianco, Jean-Luc Sida, Stéphane Hilaire, and Sophie Heinrich. New statistical scissionpoint model to predict fission fragment observables. Physical Review C (Nuclear Physics), 92(3), September 2015. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Leone:2010:FJI

[LR10] Matteo Leone and Nadia Robotti. Frédéric Joliot, Irène Curie and the early history of the positron (1932–33). European Journal of Physics, 31(4):975–987, July 2010. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://

stacks.iop.org/0143-0807/31/i=4/a=027.

Langanke:2013:MEU

[LT13] Karlheinz Langanke and Friedrich-Karl Thielemann. Making the elements in the universe. Europhysics News, 44(3):23–26, May 2013. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic).

MacGregor:2011:ERH

[Mac11] I. J. Douglas MacGregor. Ernest Rutherford: his genius shaped our modern world. Europhysics News, 42(5):18–21, September/ October 2011. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-

1092 (electronic). URL http://www.europhysicsnews.org/
articles/epn/abs/2011/05/epn2011425p18/epn2011425p18.
html.

Mahnke:2022:LMD

[Mah22] Heinz-Eberhard Mahnke. Lise Meitner, β-decay and nonradiative electromagnetic transitions. Notes and Records of the Royal Society of London, 76(1):107–116, March 20, 2022. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL https://royalsocietypublishing.org/doi/
pdf/10.1098/rsnr.2020.0036.

Maier:2007:GBW

[Mai07] Helmut Maier, editor. Gemeinschaftsforschung, Bevollmächtigte und der Wissenstransfer: die Rolle der Kaiser-Wilhelm-Gesellschaft im System kriegsrelevanter Forschung des Nationalsozialismus. (German) [Community research, plenipotentiaries and the transfer of knowledge: the role of the Kaiser Wilhelm Society in the system of Nazi-related research of war-relevant research], volume 17. Wallstein, Göttingen, Germany, 2007. ISBN 3-8353-0182-9 (paperback). LCCN Q180.G4 K38 2000 v.17; U395.G3 G45 2007.

Marsden:1954:RML

[Mar54] E. Marsden. The Rutherford Memorial Lecture, 1954. Rutherford — his life and work, 1871–1937. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 226(1166):283–305, 1954. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL http://rspa.

royalsocietypublishing.org/content/226/1166/283. Lec-

ture delivered in South Africa in May 1954.

McGrayne:1994:NPW

[MB94] Sharon Bertsch McGrayne and Nina Byers. Nobel Prize women in science. Physics Today, 47(7):63, July 1994. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

McGrayne:1993:LM

[McG93a] Sharon Bertsch McGrayne. Lise Meitner. In Nobel Prize women in science: their lives, struggles, and momentous discoveries [McG93b], page ?? ISBN 1-55972-146-4. LCCN Q141 .B42 1993.

McGrayne:1993:NPW

[McG93b] Sharon Bertsch McGrayne. Nobel Prize women in science: their lives, struggles, and momentous discoveries. Carol Publishing Group, Secaucus, NJ, USA, 1993. ISBN 1-55972-146-4. xi + 419 pp. LCCN Q141 .B42 1993.

McGrayne:1998:LM

[McG98a] Sharon Bertsch McGrayne. Lise Meitner. In Nobel Prize women in science: their lives, struggles, and momentous discoveries [McG98b], pages 37–63. ISBN 0-309-07270-0 (paperback), 0-8065-2025-6 (paperback), 0-9702256-0-1. LCCN Q141 .M358

1998. URL http://www.nap.edu/catalog/10016.html.

McGrayne:1998:NPW

[McG98b] Sharon Bertsch McGrayne. Nobel Prize women in science: their lives, struggles, and momentous discoveries. Joseph Henry Press, Washington, DC, USA, second edition, 1998. ISBN 0-309-07270-0 (paperback), 0-8065-2025-6 (paperback), 0-9702256-0-1. xi + 451 pp. LCCN Q141 .M358 1998. URL http://www.nap.edu/

catalog/10016.html.

McLay:1999:LME

[McL99] David B. McLay. Lise Meitner and Erwin Schrödinger: Biographies of two Austrian physicists of Nobel stature. Minerva [Washington, DC], 37(1):75–94, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A:1004536223625.

McLay:2001:BRL

[McL01] David B. McLay. Book review: Lise Meitner and the Dawn of the Nuclear Age by Patricia Rife. Isis, 92(3):624–625, September 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/3080783.

Meitner:1935:ZGDa

[MD35a] Lise Meitner and Max Delbrück. Der α-Zerfall. (German) [α decay]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 45–55. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:ZGDb

[MD35b] Lise Meitner and Max Delbrück. Der β-Zerfall. (German) [β decay]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 55–59. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:AAN

[MD35c] Lise Meitner and Max Delbrück. Der Aufbau der Atomkerne, natürliche und künstliche Kernumwandlungen. (German) [The structure of atomic nuclei, natural and artificial nuclear transformations]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935. 62 + 13

pp. URL http://posner.library.cmu.edu/Posner/books/
pages.cgi?call=530.12_M51A_1935.

Meitner:1935:SGR

[MD35d] Lise Meitner and Max Delbrück. Die γ-Strahlen. (German) [γ radiation]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 59–61. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:APG

[MD35e] Lise Meitner and Max Delbrück. Die allgemeine Problemlage. (German) [The general problem]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 61–62. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:RGR

[MD35f] Lise Meitner and Max Delbrück. Die Radioaktivität. (German) [Radioactivity]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 41–45. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:EGIa

[MD35g] Lise Meitner and Max Delbrück. Einleitung. (German) [introduction]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 1–6. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:EGIb

[MD35h] Lise Meitner and Max Delbrück. Einleitung. (German) [introduction]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 38–41. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:KRG

[MD35i] Lise Meitner and Max Delbrück. Künstliche Radioaktivität. (German) [Artificial radioactivity]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 17–27. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:KZG

[MD35j] Lise Meitner and Max Delbrück. Künstliche Zertrümmerung. (German) [Artificial fragmentation]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 6–17. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1935:SAG

[MD35k] Lise Meitner and Max Delbrück. Systematik der Atomkerne. (German) [Systems of atomic nuclei]. In Der Aufbau der Atomkerne. (German) [Structure of Atomic Nuclei], pages 27–38. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1935.

Meitner:1906:ASG

[Mei06a] Lise Meitner. Über die Absorption der α- und β-Strahlen. (German) [On the absorption of α- and β-rays]. Physikalische Zeitschrift, 7(17):588–590, September 1, 1906. CODEN PHZ-

TAO. ISSN 0369-982X. URL https://hdl.handle.net/2027/
nyp.33433062733229?urlappend=;seq=164.

Meitner:1906:EFSa

[Mei06b] Lise Meitner. Über einige Folgerungen, die sich aus den Fresnelschen Reflexionsformeln ergeben. (German) [On some of the implications of the Fresnel reflection formulas]. Sitzungsberichte der Kaiserlichen Akademie der Wissenschaften. Mathematisch-Naturwissenschaftliche Klasse, 115(6):859–869, 1906. CODEN SWWPAX. ISSN 0376-2629. URL https://books.google.

com/books?id=BFQhAQAAMAAJ&pg=859.

Meitner:1906:EFSb

[Mei06c] Lise Meitner. Über einige Folgerungen, die sich aus den Fresnelschen Reflexionsformeln ergeben. (German) [On some of the implications of the Fresnel reflection formulas]. S. Ber. preuss Akad. Wiss.; Math. Natur-Klasse, Abt. IIa, 115(??):259–286, June 1906.

Meitner:1906:WIK

[Mei06d] Lise Meitner. Warmeleitung in inhomogenen Korpern. II. (German) [Heat conduction in inhomogeneous bodies. II]. Phys. Inst. Wien, IIa, 115(??):125–137, February 1906. Dr. phil. dissertation.

Meitner:1907:ZSG

[Mei07] Lise Meitner. Über die Zerstreuung der α-Strahlen. (German) [On the dispersion of α rays]. Physikalische Zeitschrift, 8(15): 489–491, August 1, 1907. CODEN PHZTAO. ISSN 0369-982X.

URL https://hdl.handle.net/2027/nyp.33433062733237?
urlappend=;seq=73.

Meitner:1909:SZRa

[Mei09a] Lise Meitner. Strahlen und Zerfallsprodukte des Radiums. (German) [Rays and decay products of radium]. Physikalische Zeitschrift, 10(22):852–853, November 10, 1909. CODEN PHZTAO. ISSN 0369-982X. URL https://hdl.handle.

net/2027/nyp.33433090815253?urlappend=;seq=903.  See

[HM09c, HM09e].

Meitner:1909:SZRb

[Mei09b] Lise Meitner. Strahlen und Zerfallsprodukte des Radiums. (German) [Rays and decay products of radium]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 11(??):648–??, ???? 1909. CODEN VDPEAZ. ISSN 0372-5448.

Meitner:1910:BSR

[Mei10] Lise Meitner. Über die Beta-Strahlen der radioaktiven Substanzen, Zusammenfassender Bericht. (German) [On the beta radiation of radioactive substances, summary report]. Naturwissenschaftliche Rundschau, 25(27):337–340, July 7, 1910. CODEN NARSAC. ISSN 0028-1050. URL http://www.

digizeitschriften.de/dms/img/?PID=PPN385489110_0025%
7CLOG_0274.

Meitner:1911:RET

[Mei11a] Lise Meitner. Die radioaktiven Eigenschaften der Thoriumreihe. (German) [The radioactive properties of thorium series]. Naturwissenschaftliche Rundschau, 26(28):353–356, July 13, 1911. CODEN NARSAC. ISSN 0028-1050. URL

http://www.digizeitschriften.de/dms/resolveppn/?PID=
PPN385489110_0026%7CLOG_0280.

Meitner:1911:EEH

[Mei11b] Lise Meitner. Über einige einfache Herstellungsmethoden radioaktiver Zerfallsprodukte. (German) [On some simple methods of manufacturing radioactive decay products]. Physikalische Zeitschrift, 12(24):1094–1099, December 1, 1911. CODEN PHZ-

TAO. ISSN 0369-982X. URL https://hdl.handle.net/2027/
mdp.39015011417709?urlappend=;seq=1152.

Meitner:1911:VGR

[Mei11c] Lise Meitner. Vorträge aus dem Gebiet der Radioaktivität. (German) [Lectures from the field of radioactivity]. Physikalische Zeitschrift, 12(4):147–151, February 15, 1911. CODEN PHZ-

TAO. ISSN 0369-982X. URL https://hdl.handle.net/2027/
mdp.39015011417709?urlappend=;seq=183.

Meitner:1912:ZAN

[Mei12] Lise Meitner. Über das Zerfallsschema des aktiven Niederschlags des Thoriums. (German) [On the decay scheme of thorium active precipitation]. Physikalische Zeitschrift, 13(13):623–626, July 1, 1912. CODEN PHZTAO. ISSN 0369-982X. URL https://hdl.

handle.net/2027/nyp.33433090814983?urlappend=;seq=55.

Meitner:1915:SWP

[Mei15a] Lise Meitner. Über die α-Strahlung des Wismuts aus Pechblende. Physikalische Zeitschrift, 16(13):4–6, July 1, 1915. CO-DEN PHZTAO. ISSN 0369-982X. URL https://hdl.handle.

net/2027/njp.32101054770928?urlappend=;seq=24.

Meitner:1915:SRD

[Mei15b] Lise Meitner. Über die β-Strahlen von Ra D und das Absorptionsgesetz der β-Strahlen. (German) [On the β rays of Ra D and the absorption law of β rays]. Physikalische Zeitschrift, 16(15):272–276, August 1, 1915. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/njp.
32101054770928?urlappend=;seq=320.

Meitner:1918:LRM

[Mei18a] Lise Meitner. Die Lebensdauer von Radiothor, Mesothor und Thorium. (German) [The lifetime of radiothorium, mesothorium and thorium]. Physikalische Zeitschrift, 19(12):257–262, June 15, 1918. CODEN PHZTAO. ISSN 0369-982X. URL https://

hdl.handle.net/2027/hvd.hxgq92?urlappend=;seq=285.

Meitner:1918:MAN

[Mei18b] Lise Meitner. Die Muttersubstanz des Actiniums, Ein Neues Radioaktives Element von Langer Lebensdauer. (German) [The parent substance of actinium, a new radioactive element of long life]. Berichte der Bunsen-Gesellschaft für Physikalische Chemie, 24(11–12):169–, June 1, 1918. CODEN BBPCAX. ISSN

0005-9021. URL https://onlinelibrary.wiley.com/doi/10.
1002/bbpc.19180241107.

Meitner:1918:PGP

[Mei18c] Lise Meitner. Über das Protactinium. (German) [On protactinium]. Die Naturwissenschaften, 6(22):324–326, May 1918. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1921:RAGa

[Mei21a] Lise Meitner. Radioaktivität und Atomkonstitution. (German) [Radioactivity and atomic constitution]. In Festschrift der Kaiser Wilhelm Gesellschaft zur Förderung der Wissenschaften zu ihrem ehnjährigen Jubiläum Dargebracht von ihren Instituten, pages 154–161. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1921.

Meitner:1921:RAGb

[Mei21b] Lise Meitner. Radioaktivität und Atomkonstitution. (German) [Radioactivity and atomic constitution]. Die Naturwissenschaften, 9(22):423–427, June 1921. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1921:VAR

[Mei21c] Lise Meitner. Über die verschiedenen Arten des radioaktiven Zerfalls und die Möglichkeit ihrer Deutung aus der Kernstruktur. (German) [On the different types of radioactive decay and the possibility of their interpretation from nuclear structure]. Zeitschrift für Physik, 4(1):146–156, March 1921. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1922:NVA

[Mei22a] Lise Meitner. Notiz zur vorstehenden Arbeit: [Das Neuburgersche Kernmodell und die Erfahrung]. (German) [Note on the above work: [the Neuburger nuclear model and the experience]]. Physikalische Zeitschrift, 23(15):305, August 1, 1922. CODEN PHZTAO. ISSN 0369-982X. URL https://hdl.handle.net/

2027/mdp.39015086723239?urlappend=;seq=371.

Meitner:1922:ZZS

[Mei22b] Lise Meitner. Über den Zusammenhang zwischen β- und γ-Strahlen. (German) [On the relationship between β and γ rays]. Zeitschrift für Physik, 9(1):145–152, December 1922. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1922:SSI

[Mei22c] Lise Meitner. Über die α-Strahl-Spektra und ihren Zusammenhang mit der γ-Strahlung. (German) [On the α beam spectra and their correlation with γ radiation]. Zeitschrift für Physik, 11 (1):35–54, December 1922. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1922:ESS

[Mei22d] Lise Meitner. Über die Entstehung der β-Strahl-Spektren radioaktiver Substanzen. (German) [On the origin of the β radiation spectra of radioactive substances]. Zeitschrift für Physik, 9(1):131–144, December 1922. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1922:WSG

[Mei22e] Lise Meitner. Über die Wellenlänge der γ-Strahlen. (German) [On the wavelength of γ rays]. Die Naturwissenschaften, 10(16): 381–384, April 21, 1922. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1923:SUX

[Mei23a] Lise Meitner. Das β-Strahlenspektrum von U X_{1} und seine Deutung. (German) [The β radiation spectrum of U X_{1} and its interpretation]. Zeitschrift für Physik, 18(1):238, December 1923. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1923:SUS

[Mei23b] Lise Meitner. Das β-Strahlenspektrum von UX_{1} und seine Deutung. (German) [The β radiation spectrum of UX_{1} and its interpretation]. Zeitschrift für Physik, 17(1):54–66, December 1923. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1923:BRK

[Mei23c] Lise Meitner. Die Bedeutung der Radioaktivität für kosmische Prozesse. (German) [The importance of radioactivity for cosmic processes]. Zeitschrift für Angewandte Chemie, 36(2):9–11, January 1923. ISSN 0932-2132.

Meitner:1923:MDK

[Mei23d] Lise Meitner. Über eine mögliche Deutung des kontinuierlichen β-Strahlenspektrums. (German) [On a possible interpretation of the continuous β-ray spectrum]. Zeitschrift für Physik, 19(1): 307–312, December 1923. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1924:ZZS

[Mei24a] Lise Meitner. Der Zusammenhang zwischen β- und γ-Strahlen. (German) [The relationship between β and γ rays]. In Ergebnisse der Exakten Naturwissenschaften, volume 3, pages 160– 181. Springer-Verlag, Berlin, Germany / Heidelberg, Germany /

London, UK / etc., 1924. URL http://posner.library.cmu.
edu/Posner/books/pages.cgi?call=508_M51Z&layout=vol0/
part0/copy0&file=0167.

Meitner:1924:SAG

[Mei24b] Lise Meitner. Die Sichtbarmachung der Atome. (German) [The visualization of atoms]. Kaiser-Wilhelm Inst. Chem., Annual report, ??(??):??, ???? 1924.

Meitner:1924:ERZ

[Mei24c] Lise Meitner. Über die Energieentwicklung bei radioaktiven Zerfallsprozessen. (German) [On energy development in radioactive decay processes]. Die Naturwissenschaften, 12(50):1146–1150, December 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1924:RSB

[Mei24d] Lise Meitner. Über die Rolle der γ-Strahlen beim Atomzerfall. (German) [On the role of γ rays in atomic decay]. Zeitschrift für Physik, 26(1):169–177, December 1924. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1924:NFC

[Mei24e] Lise Meitner. Über eine notwendige Folgerung aus dem Comptoneffekt und ihre Bestätigung. (German) [On a necessary conclusion from the Compton effect and its confirmation]. Zeitschrift für Physik, 22(1):334–342, December 1924. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1925:SAN

[Mei25] Lise Meitner. Die γ-Strahlung der Actiniumreihe und der Nachweis, daß die γ-Strahlen erst nach erfolgtem Atomzerfall emittiert werden. (German) [The γ radiation of the actinium series and evidence that the γ rays are emitted only after the atomic decomposition has occurred]. Zeitschrift für Physik, 34(1):807– 818, December 1925. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1926:K

[Mei26a] Lise Meitner. Die Kernstruktur. (German) [Nuclear structure]. In Handbuch der Physik, volume 22, pages 124–?? Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1926.

Meitner:1926:EBI

[Mei26b] Lise Meitner. Einige Bemerkungen zur Isotopie der Elemente. (German) [Some comments on isotopes of the elements]. Die Naturwissenschaften, 14(30):719–720, July 1926. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1926:EBR

[Mei26c] Lise Meitner. Experimentelle Bestimmung der Reichweite homogener β-Strahlen. (German) [Experimental determination of the range of homogeneous β rays]. Die Naturwissenschaften, 14 (50–51):1199–1203, December 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1926:NAB

[Mei26d] Lise Meitner. Neuere Arbeiten ber die Streuung der α-Strahlen und den Aufbau der Atomkerne. (German) [Recent work on the dispersion of α rays and the structure of atomic nuclei]. Die Naturwissenschaften, 14(38):863–869, September 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1926:AIS

[Mei26e] Prof. Dr. Lise Meitner. Atomvorgänge und ihre Sichtbarmachung; Vortrag gehalten in der Münchner Chemischen Gesellschaft am 16. Juli 1925. (German) [Atomic processes and their visualization; Lecture held at the Munich Chemical Society on July 16, 1925]. F. Enke, Stuttgart, Germany, 1926. 32 pp.

Meitner:1927:RRS

[Mei27a] Meitner. [Rezension:] Radioaktivität. Von Stefan Meyer und Egon v. Schweidler. Zweite vermehrte Auflage. Leipzig 1927. B. G. Teubner Geb. 36,– M.. (German) [Radioactivity. By Stefan Meyer and Egon v. Schweidler. Second expand edition. Leipzig 1927. B. G. Teubner 36,- M.]. Zeitschrift für Angewandte Chemie, 40(36):1022, September 1927. ISSN 0932-2132.

Meitner:1927:AAG

[Mei27b] Lise Meitner. Über den Aufbau des Atominnern. (German) [On the structure of the atomic interior]. Die Naturwissenschaften, 15(16):369–378, April 1927. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1928:BRC

[Mei28a] Lise Meitner. [Book review]: The chemical effects of α-particles and electrons. Von L. C. Lind. Second (revised and enlarged) Edition American Chemical Society Monograph Series. The Chemical Catalog Comp. Inc. New York U.S.A. 1928. Zeitschrift für Angewandte Chemie, 41(46):1249–1250, November 1928. ISSN 0932-2132.

Meitner:1928:SRA

[Mei28b] Lise Meitner. Das β-Strahlenspektrum des Radiothors als Absorptions-spektrum seiner γ-Strahlen. (German) [The β radiation spectrum of radiothorium as absorption spectrum of its γ radiation]. Zeitschrift für Physik, 52(9–10):637–644, September 1928. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1928:SPE

[Mei28c] Lise Meitner. Das γ-Strahlenspektrum des Protactiniums und die Energie der γ-Strahlen bei α- und β-Strahlenumwandlungen. (German) [The γ =beam spectrum of protactinium and the energy of γ beams on α and β beam conversions]. Zeitschrift für Physik, 50(1–2):15–23, January 1928. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1928:SRE

[Mei28d] Lise Meitner. Das γ-Strahlenspektrum des Radiothors in Emission. (German) [The γ radiation spectrum of radiothorium in emission]. Zeitschrift für Physik, 52(9–10):645–649, September 1928. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1928:GDG

[Mei28e] Lise Meitner. Über geeignete Dampf–Gasgemische fur verschiedene Versuche nach der Wilsonschen Nebelmethode. (German) [On suitable vapor–gas mixtures for various experiments according to the Wilson cloud method]. Zeitschrift für physikalische Chemie (Leipzig), 139A(1):717–724, December 1928. CODEN ZPCLAH. ISSN 0942-9352 (print), 2196-7156

(electronic). URL https://www.degruyter.com/view/j/zpch.
1928.139.issue-1/zpch-1928-13944/zpch-1928-13944.xml.

Meitner:1929:SGRb

[Mei29a] Lise Meitner. Die γ-Strahlen. (German) [γ radiation]. In Lehrbuch der Physik, pages 2097–2118. Vieweg + Teubner Verlag, Braunschweig, Germany, 1929.

Meitner:1929:SGRa

[Mei29b] Lise Meitner. Die γ-Strahlen. (German) [γ rays]. In Müller– Pouillet Lehrbuch der Physik und Meteorologie. (German) [Müller–Pouillet Textbook of Physics and Meteorology], volume 2(3), pages 132–?? Friedrich Vieweg und Sohn, Braunschweig, Germany, eleventh edition, 1929.

Meitner:1929:HIB

[Mei29c] Lise Meitner. Die Höhenstrahlung und ihre Beziehung zu physikalischen und kosmischen Vorgängen. (German) [Cosmic radiation and its relation to physical and cosmic processes]. Zeitschrift für Angewandte Chemie, 42(14):345–351, April 1929. ISSN 0932-2132.

Meitner:1929:RSB

[Mei29d] Lise Meitner. Die Reichweite von β-Strahlen bestimmter Geschwindigkeit. (German) [The range of β-radiation of certain velocity]. Forschungen und Fortschritte, 5:321–322, 1929. CO-DEN FOFOAI. ISSN 0367-2794.

Meitner:1929:EPD

[Mei29e] Lise Meitner. Energieverteilung der primären β-Strahlen und die daraus zu folgende γ-Strahlung. (German) [Energy distribution of primary β rays and the resulting γ radiation]. Physikalische Zeitschrift, 30:515–??, ???? 1929. CODEN PHZTAO. ISSN 0369-982X.

Meitner:1929:RRH

[Mei29f] Lise Meitner. [Rezension:] Radioaktivität. Handbuch der Experimentalphysik. Von K. W. F. Kohlrausch. Herausgegeben von W. Wien und F. Harms. Bd. 15. Mit 285 Abbildungen. Akademische Verlagsgesellschaft m. b. H., Leipzig 1928. Geb. RM. 81,– . (German) [[Review:] Radioactivity. Handbook of Experimental Physics. By K. W. F. Kohlrausch. Edited by W. Wien and F. Harms. Vol. 15. With 285 illustrations. Academic Publishing Company Ltd., Leipzig 1928. Geb. RM. 81,–]. Zeitschrift für Angewandte Chemie, 42(15):392–393, April 1929. ISSN 0932-2132.

Meitner:1931:BAM

[Mei31a] Lise Meitner. Die Bedeutung des Atomgewichtes in der modernen Physik. (German) [The importance of atomic weight in modern physics]. Berichte der deutschen chemischen Gesellschaft (Abteilung A und B), 64(8):A149–A162, September 1931. CO-DEN BDCBAD. ISSN 1099-0682 (print), 1434-1948 (electronic).

Meitner:1931:IIN

[Mei31b] Lise Meitner. Über die Ionisierungswahrscheinlichkeit innerer Niveaus durch schnelle Korpuskularstrahlen und eine Methode zu ihrem Nachweis. (German) [On the ionization probability of internal levels by fast corpuscular rays and a method for their proof]. Die Naturwissenschaften, 19(23–25):497–499, June 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1931:WZM

[Mei31c] Lise Meitner. Über Wechselbeziehungen zwischen Masse und Energie. (German) [On interactions between mass and energy]. Zeitschrift des Vereines deutsches Ingenieure, 75(??):977–980, ???? 1931.

Meitner:1932:BSF

[Mei32a] Lise Meitner. Die Bedeutung der β und γ-Strahlen für die Atomforschung. (German) [The importance of β and γ rays for nuclear research]. Zeitschrift für Elektrochemie, 38(??):490–??, ???? 1932. CODEN ZEELAI. ISSN 0372-8382.

Meitner:1932:RLR

[Mei32b] Lise Meitner. [Rezension:] Lehrbuch der Radioaktivität. Von G. v. Hevesy und F. Paneth. Mit 50 Abbildungen im Text. Johann Ambrosius Barth, Leipzig 1931. Preis brosch. RM. 22,–; geb. RM. 24,–. (German) [[Review:] textbook of radioactivity. By G. v. Hevesy and F. Paneth. With 50 pictures in the text. Johann Ambrosius Barth, Leipzig 1931. Price brochure. RM. 22,–; b RM. 24,–]. Angewandte Chemie, 45(13):258–259, March 26, 1932. CO-DEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Meitner:1934:APS

[Mei34a] Lise Meitner. Atomkern und periodisches System der Elemente. (German) [Atomic nucleus and periodic system of elements]. Die Naturwissenschaften, 22(44):733–739, November 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1934:EPE

[Mei34b] Lise Meitner. Das Energiespektrum der positiven Elektronen aus Aluminium. (German) [The energy spectrum of the positive electrons of aluminum]. Die Naturwissenschaften, 22(22– 24):388–390, June 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1934:SHS

[Mei34c] Lise Meitner. Die Streuung harter γ-Strahlen. (German) [The scattering of hard γ-rays]. Die Naturwissenschaften, 22(11):174, March 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1934:NA

[Mei34d] Lise Meitner. Neuere Atomkernprozesse. (German) [Newer atomic nuclear processes]. Forschungen und Fortschritte, 10(??): 246–??, ???? 1934. CODEN FOFOAI. ISSN 0367-2794.

Meitner:1934:EKR

[Mei34e] Lise Meitner. Über die Erregung künstlicher Radioaktivität in verschiedenen Elementen. (German) [On the excitement of artificial radioactivity in different elements]. Die Naturwissenschaften, 22(22–24):420, June 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1934:UED

[Mei34f] Lise Meitner. Über die Umwandlung der Elemente durch Neutronen. (German) [On the conversion of the elements by neutrons]. Die Naturwissenschaften, 22(45):759, November 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1934:CFJ

[Mei34g] Lise Meitner. Über die von I. Curie und F. Joliot entdeckte künstliche Radioaktivität. (German) [On the artificial radioactivity discovered by I. Curie and F. Joliot]. Die Naturwissenschaften, 22(11):172–174, March 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1936:KVP

[Mei36a] Lise Meitner. Kernphysikalische Vorträge am Physikalischen Institut der Eidgenössischen Technischen Hochschule. (German) [Nuclear Physics Lectures at the Physical Institute of the Swiss Federal Institute of Technology]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1936. ???? pp.

Meitner:1936:KUB

[Mei36b] Lise Meitner. Künstliche Umwandlungsprozesse beim Uran. (German) [Artificial transformation processes in uranium]. In Egon Bretscher, editor, Kernphysik: Vorträge gehalten am Physikalischen Institut der E.T.H. Zürich im Sommer 1936. (German) [Nuclear Physics: Presented at the Physical Institute of E.T.H. Zurich in the summer of 1936], pages 24–42. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1936.

Meitner:1936:RCU

[Mei36c] Lise Meitner. Radioaktive und chemische Untersuchungen an der Transuranelementen. (German) [Radioactive and chemical investigations on the transuranic elements]. Zeitschrift für Angewandte Chemie, 49(38):692–??, September 19, 1936. ISSN 0932-

2132. URL https://onlinelibrary.wiley.com/doi/pdf/10.
1002/ange.19360493805.

Meitner:1937:STG

[Mei37] Lise Meitner. Über die β- und γ-Strahlen der Transurane. (German) [On β and γ rays of transuranium]. Annalen der Physik (1900), 421(3–4):246–250, 1937. ISSN 0003-3804 (print), 1521-3889 (electronic).

Meitner:1939:NPF

[Mei39a] Lise Meitner. New products of the fission of the thorium nucleus. Nature, 143(3624):637, April 15, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v143/n3624/
pdf/143637a0.pdf.

Meitner:1939:OHS

[Mei39b] Lise Meitner. Otto Hahn zum sechzigsten Geburtstag (8. März 1939). (German) [Otto Hahn’s 60th birthday (8 march 1939)]. Current Science, 8(5):204–205, May 1939. CODEN CUS-

CAM. ISSN 0011-3891. URL http://www.jstor.org/stable/
24205120.

Meitner:1940:CCS

[Mei40] Lise Meitner. Capture cross-sections for thermal neutrons in thorium, lead and uranium 238. Nature, 145(3672):422–423, March 16, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Meitner:1941:BRE

[Mei41a] Lise Meitner. Behaviour of rare earths under neutron bombardment. Arkiv för Matematik, Astronomi och Fysik, 17(??):18–??, ???? 1941. CODEN AMAFAX. ISSN 0365-4133.

Meitner:1941:RET

[Mei41b] Lise Meitner. The resonance energy of the thorium capture process. Physical Review, 60(1):58–60, July 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Meitner:1941:BVE

[Mei41c] Lise Meitner. Über das Verhalten einiger seltener Erden bei Neutronenbestrahlung. (German) [On the behavior of some rare earths in neutron irradiation]. Arkiv för Matematik, Astronomi och Fysik, 27(??):3–??, ???? 1941. CODEN AMAFAX. ISSN 0365-4133. Paper 17.

Meitner:1942:RES

[Mei42a] Lise Meitner. Radiations emitted by scandium 46. Arkiv för Matematik, Astronomi och Fysik, 28B(4(14)):6–??, ???? 1942. CODEN AMAFAX. ISSN 0365-4133.

Meitner:1942:BSE

[Mei42b] Lise Meitner. Über die von Sc 46 emittierten Strahlen. (German) [On the rays emitted by Sc 46]. Arkiv för Matematik, Astronomi och Fysik, 28B(??):14–??, ???? 1942. CODEN AMAFAX. ISSN 0365-4133.

Meitner:1943:EMM

[Mei43a] Lise Meitner. Eine Einfache Methode zur Messung der durch γ-Strahlen erregten sekundären Elektronen und der Einfluß dieser Elektronen auf die Untersuchung primärer β-Strahlspektren. (German) [A simple method for measuring secondary electrons fired by γ-radiation and the influence of these electrons on the study of primary β beam spectra]. Arkiv för Matematik, Astronomi och Fysik, 29A(??):1–14, ???? 1943. CODEN AMAFAX. ISSN 0365-4133. Paper 17.

Meitner:1943:ESM

[Mei43b] Lise Meitner. Erratum: Simple method for the investigation of secondary electrons excited byγ-rays and the interference of these electrons with measurements of primaryβ-ray spectra (Phys. Rev. 63, 73 (1943)). Physical Review, 63(9–10):384, May 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). See [Mei43c].

Meitner:1943:SMIb

[Mei43c] Lise Meitner. A simple method for the investigation of secondary electrons excited byγ-raysand the interference of these electrons with measurements of primaryβ-RaySpectra. Physical Review, 63(3–4):73–76, February 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). See erratum [Mei43b].

Meitner:1943:SMIa

[Mei43d] Lise Meitner. A simple method for the investigation of secondary electrons the interference of these electrons with measurements of primary β-ray spectra. Arkiv för Matematik, Astronomi och Fysik, 17(??):14–??, ???? 1943. CODEN AMAFAX. ISSN 0365-4133.

Meitner:1945:ASS

[Mei45a] Lise Meitner. An attempt to single out some fission processes of uranium by using the differences in their energy release. Reviews of Modern Physics, 17(2–3):287–291, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.17.287;   http://rmp.aps.org/abstract/RMP/
v17/i2-3/p287_1.

Meitner:1945:ZSG

[Mei45b] Lise Meitner. Das Zerfallsschema des Scandium 46. (German) [The decay scheme of scandium 46]. Arkiv för Matematik, Astronomi och Fysik, 32A(??):6–??, ???? 1945. CODEN AMAFAX. ISSN 0365-4133.

Meitner:1945:DSS

[Mei45c] Lise Meitner. Disintegration scheme of scandium 46. Arkiv för Matematik, Astronomi och Fysik, 32A(??):11–??, ???? 1945. CODEN AMAFAX. ISSN 0365-4133.

Meitner:1946:NA

[Mei46a] Lise Meitner. The nature of the atom. Fortune, ??(??):137–188, February 1946. CODEN FORTAP. ISSN 0015-8259.

Meitner:1946:RNP

[Mei46b] Lise Meitner. Radioactive nuclei produced by irradiation of copper with deuterons. Arkiv för Matematik, Astronomi och Fysik, 33A(1(3)):1–12, 1946. CODEN AMAFAX. ISSN 0365-4133.

Meitner:1948:EBE

[Mei48] Lise Meitner. Einige Bemerkungen zu den Einfangquerschnitten langsamer und schneller Neutronen bei schwereren Elementen. (German) [Some remarks on slower and faster neutron capture cross sections for heavier elements]. Annalen der Physik (1900), 438(1):115–119, 1948. ISSN 0003-3804 (print), 1521-3889 (electronic).

Meitner:1949:OHM

[Mei49] Lise Meitner. Geburtstag: Otto Hahn zum 8. März 1949. (German) [Birthday: Otto Hahn on 8 March 1949 [70th]]. Zeitschrift für Naturforschung, 4A(2):81, February 1949. CODEN ZNTFA2.

ISSN 0372-9516. URL https://www.degruyter.com/view/j/
zna.1949.4.issue-2/zna-1949-0201/zna-1949-0201.xml.

Meitner:1950:FNS

[Mei50a] Lise Meitner. Fission and nuclear shell model. Nature, 165 (4197):561, April 8, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature. com/articles/165561a0. English translation of German original [Mei50b].

Meitner:1950:SSA

[Mei50b] Lise Meitner. Spaltung und Schalenmodell des Atomkerns. (German) [Cleavage and shell model of the atomic nucleus]. Arkiv för fysik, 4(??):383 (165, 561), ???? 1950. CODEN AFYSA7. ISSN 0365-2440. English translation in [Mei50a].

Meitner:1952:ARA

[Mei52a] Lise Meitner. Die Anwendung des Rückstoßes bei Atomkernprozessen. (German) [The application of recoil in nuclear processes]. Zeitschrift für Physik, 133(1–2):140–152, September 1952. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1952:ERO

[Mei52b] Lise Meitner. Zur Entwicklung der Radiochemie: Otto Hahn zum 50 jährigen Doktor-Jubiläum. (German) [On the development of radiochemistry: Otto Hahn on the occasion of his 50th doctoral anniversary]. Angewandte Chemie, 64(1):1–4, January 1952. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Meitner:1954:EEK

[Mei54] Lise Meitner. Einige Erinnerungen an das Kaiser-Wilhelm-Institut für Chemie in Berlin-Dahlem. (German) [Some memories of the Kaiser Wilhelm Institute for Chemistry in Berlin-Dahlem]. Die Naturwissenschaften, 41(5):97–99, 1954. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1956:KAA

[Mei56] Lise Meitner. Kein Anteil an der Atombombe. (German) [No part in the atomic bomb]. In Seelig [See56], page 113. ISBN 3-322-84225-8. LCCN ???? With contributions and letters from Hedwig Born.

Meitner:1957:AGA

[Mei57] Lise Meitner. Das Atom. (German) [The atom]. In Schröder [Sch57], page ?? LCCN QC776 .A865 1957.

Meitner:1958:MPM

[Mei58] Lise Meitner. Max Planck als Mensch. (German) [Max Planck as a human being]. Die Naturwissenschaften, 45(17):406–408, September 1, 1958. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1959:OHG

[Mei59] Lise Meitner. Otto Hahn zum 80. Geburtstag am 8. März 1959. (German) [Otto Hahn’s 80th birthday on March 8, 1959]. Die Naturwissenschaften, 46(5):157–158, 1959. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1960:SWP

[Mei60] Lise Meitner. The status of women in the professions. Physics Today, 13(8):16–21, August 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Meitner:1962:RWR

[Mei62] Lise Meitner. Right and wrong roads to the discovery of nuclear energy. International Atomic Energy Agency Bulletin, 4(0):4–6, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-

2690 (electronic). URL http://www.iaea.org/Publications/
Magazines/Bulletin/Bull040su/04004790608su.pdf.

Meitner:1963:RWR

[Mei63a] Lise Meitner. Right and wrong roads to the discovery of nuclear energy. Advancement of Science, 19(??):363–365, ???? 1963. CODEN ADSCAH. ISSN 0001-866X. English translation of [Mei63b].

Meitner:1963:WIK

[Mei63b] Lise Meitner. Wege und Irrwege zur Kernenergie. (German) [right and wrong roads to the discovery of nuclear energy]. Naturwissenschaftliche Rundschau, 16(??):167–169, May ??, 1963. CODEN NARSAC. ISSN 0028-1050. English translation in [Mei63a].

Meitner:1964:LMLa

[Mei64a] Lise Meitner. Lise Meitner looks back. Advancement of Science, 21(??):39–46, ???? 1964. CODEN ADSCAH. ISSN 0001-866X.

Meitner:1964:LMLb

[Mei64b] Lise Meitner. Lise Meitner looks back. International Atomic Energy Agency Bulletin, 6(1):4–12, ???? 1964. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http://

www.iaea.org/Publications/Magazines/Bulletin/Bull061/
06101400412.pdf.

Meitner:1964:LB

[Mei64c] Lise Meitner. Looking back. Bulletin of the Atomic Scientists, 20(9):2–7, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meitner:1964:OHG

[Mei64d] Lise Meitner. Otto Hahn zum 85. Geburtstag. (German) [Otto Hahn’s 85th birthday]. Die Naturwissenschaften, 51(5):97, 1964. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1968:KAA

[Mei68] Lise Meitner. Kein Anteil an der Atombombe. (German) [No part of the atomic bomb]. IJBF Online Internationale Jahresbibliographie der Festschriften, ??(??):??, 1968. ISSN 0724-2298 (print), 1865-0244 (electronic). URL https://www.degruyter.

com/view/IJBF/ID160829487.

Meitner:1979:OHG

[Mei79] Lise Meitner. Otto Hahn zum 80. Geburtstag. (German) [Otto Hahn’s 80th birthday]. In Hahn [Hah79], pages 308–?? ISBN 3-471-77841-1. LCCN QD22.H2 O87.

Meitner:2001:MPM

[Mei01] Lise Meitner. Max Planck als Mensch. (German) [Max Planck as a human]. In Vorträge Reden Erinnerungen. (German) [Lectures Speeches Memories], pages 77–83. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Meitner:2003:FWH

[Mei03] Lise Meitner. Die Frau in der Wissenschaft. (German) [The Woman in science]. Supposé, Köln, Germany, 2003. ISBN 3-932513-46-0. LCCN ???? Audio CD (57 minutes). Produced by Klaus Sander and Jost Lemmerich.

Meitner:2005:LME

[Mei05] Lise Meitner. Lise Meitner: Erinnerungen an Otto Hahn; mit Beiträgen von Mitarbeitern und Weggefährten. (German) [Lise Meitner: Memories of Otto Hahn; with contributions from employees and companions]. S. Hirzel Verlag, Stuttgart, Germany, 2005. ISBN 3-7776-1380-0 (paperback). 167 pp. LCCN

???? URL http://catalog.hathitrust.org/api/volumes/

oclc/71048128.html. Edited by Dietrich Hahn.

Meitner:2017:NA

[Mei17] Lise Meitner. The nature of the atom. In Johnsen [Joh17], pages 22–36.

Meitner:1924:PNS

[MF24] Lise Meitner and Kurt Freitag. Photographischer Nachweis von α-Strahlen langer Reichweite nach der Wilsonschen Nebelmethode. (German) [Photographic evidence for long-range α-radiation using Wilson’s cloud method]. Die Naturwissenschaften, 12(31):634–635, August 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1926:BGC

[MF26a] Lise Meitner and Kurt Freitag. Berichtigung. (German) [Correction]. Zeitschrift für Physik, 38(6–7):574, June 1926. CODEN ZEPYAA. ISSN 0044-3328. See [MF26b].

Meitner:1926:STC

[MF26b] Lise Meitner and Kurt Freitag. Über die α-Strahlen des ThC + C’ und ihr Verhalten beim Durchgang durch verschiedene Gase. (German) [On the α rays of ThC + C’ and their behavior on passing different gases]. Zeitschrift für Physik, 37(7–8):481–517, July 1926. CODEN ZEPYAA. ISSN 0044-3328. See correction [MF26a].

Meitner:1939:DUN

[MF39a] Lise Meitner and Otto Robert Frisch. Disintegration of uranium by neutrons: a new type of nuclear reaction. Nature, 143(3615): 239–240, February 11, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v143/n3615/pdf/143239a0.pdf. This paper,

and [MF39c], both submitted 16 January 1939 (see [Mei62]), provided the first published explanation of nuclear disintegration, called ‘nuclear fission’ by Frisch, that was first observed experimentally by Hahn and Strassmann in December 1939 [HS39]. Frisch’s paper [Fri39] describes the first experimental confirmation. It was these results that Niels Bohr intended to hold confidential until their journal publication during his January 1939 trip to the USA, but his traveling companion Léon Rosenfeld [Ros72] wasn’t informed of that intent, and the news escaped and spread quickly.

Meitner:1939:PFUa

[MF39b] Lise Meitner and Otto Robert Frisch. On the products of the fission of uranium and thorium under neutron bombardment. Math.-fys. Meddr, 17(5):1–13, 1939.

Meitner:1939:PFUb

[MF39c] Lise Meitner and Otto Robert Frisch. Products of the fission of the uranium nucleus. Nature, 143(3620):471–472, March 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v143/n3620/pdf/143471a0.pdf. See note in [MF39a].

Meitner:1968:DUN

[MF68] Lise Meitner and Otto Robert Frisch. Disintegration of uranium by neutrons. A new type of nuclear reaction. In Source Book in Chemistry, 1900–1950 [Lei68], pages 244–246. ISBN 0-674-36669-7, 0-674-36670-0. LCCN QD3. URL https:

//www.degruyter.com/view/books/harvard.9780674366701/
harvard.9780674366701.c80/harvard.9780674366701.c80.
xml.

Meitner:1969:DUN

[MF69] Lise Meitner and O. R. Frisch. Disintegration of uranium by neutrons: a new type of nuclear reaction (reprinted from Nature, February 11, 1939). Nature, 224(5218):466–467, November 1, 1969. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v224/n5218/abs/224466a0.html.

Meitner:1927:BGR

[MG27] L. Meitner and W. Grotrian. Besprechungen. (German) [Review]. Die Naturwissenschaften, 15(27):562–564, July 1927. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1913:VSE

[MH13] Lise Meitner and Otto Hahn. Über die Verteilung der γ-Strahlen auf die einzelnen Produkte der Thoriumreihe. (German) [On distributing γ rays to each thorium series product]. Physikalische Zeitschrift, 14(18):873–877, September 15, 1913. CODEN PHZ-

TAO. ISSN 0369-982X. URL https://hdl.handle.net/2027/
mdp.39015021268936?urlappend=;seq=991.

Meitner:1930:PSKb

[MH30a] L. Meitner and H. H. Hupfeld. Prüfung der Streuungsformel von Klein und Nishina an kurzwelliger γ-Strahlung. (German) [Examining Klein and Nishina’s scattering formulas on shortwave γ radiation]. Physikalische Zeitschrift, 31:947–948, 1930. CODEN PHZTAO. ISSN 0369-982X.

Meitner:1930:PSKa

[MH30b] L. Meitner and H. H. Hupfeld. Über die Prüfung der Streuungsformel von Klein und Nishina an kurzwelliger γ-Strahlung. (German) [On testing Klein and Nishina’s scattering formula on shortwave γ-radiation]. Die Naturwissenschaften, 18(22):534– 535, May 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1931:SKS

[MH31a] L. Meitner and H. H. Hupfeld. Über das Streugesetz kurzwelliger γ-Strahlen. (German) [On the scattering law of short-wave γrays]. Die Naturwissenschaften, 19(37):775–776, September 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1931:AKS

[MH31b] Lise Meitner and H. H. Hupfeld. Über das Absorptionsgesetz für kurzwellige γ-Strahlung. (German) [On the absorbance law for shortwave γ radiation]. Zeitschrift für Physik, 67(3–4):147–168, March 1931. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1932:SKS

[MH32] Lise Meitner and H. H. Hupfeld. Über die Streuung kurzwelliger γ-Strahlung an schweren Elementen. (German) [On scattering of short-wave γ radiation on heavy elements]. Zeitschrift für Physik, 75(11–12):705–715, November 1932. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1936:NUB

[MH36] Lise Meitner and Otto Hahn. Neue Umwandlungsprozesse bei Bestrahlung des Urans mit Neutronen. (German) [New transformation processes upon irradiation of uranium with neutrons]. Die Naturwissenschaften, 24(10):158–159, March 1936. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Menke:1967:WWT

[MH67] H. Menke and G. Herrmann. Was waren die Transurane von O. Hahn, L. Meitner und F. Strassmann in Wirklichkeit?. (German) [What was the transuranium of O. Hahn, L. Meitner and F. Strassmann in reality?]. Angewandte Chemie, 79(22):1005, November 1967. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Meitner:1933:SKSa

[MHK33] Lise Meitner, H. H. Hupfeld, and H. Köster. Streuung kurzwelliger γ-Strahlen. (German) [Scattering of shortwave γ rays]. Helvetica Physica Acta, 6(6):445–450, 1933. CODEN

HPACAK. ISSN 0018-0238. URL https://www.e-periodica.
ch/digbib/view?pid=hpa-001:1933:6#447.

Meitner:1937:BGR

[MhKG37] Lise Meitner, A. K. hl, M. Kobel, and L. Geitler. Besprechungen. (German) [Review]. Die Naturwissenschaften, 25(37):603–606, September 1937. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1937:UUD

[MHS37] L. Meitner, O. Hahn, and F. Strassmann. Über die Umwandlungsreihen des Urans, die durch Neutronenbestrahlung erzeugt werden. (German) [On the transformation series of uranium produced by neutron irradiation]. Zeitschrift für Physik, 106(3–4): 249–270, March 1937. CODEN ZEPYAA. ISSN 0044-3328.

Michels:1948:WLP

[Mic48] Walter C. Michels. Women … in physics. Physics Today, 1 (8):16–19, December 1948. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Meitner:1931:BGR

[MJG^{+}31] L. Meitner, P. Jordan, W. Grotrian, I. Koppel, and A. Rosenheim. Besprechungen. (German) [Review]. Die Naturwissenschaften, 19(30):660–664, July 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1933:SKSb

[MK33] L. Meitner and H. Kösters. Über die Streuung kurzwelliger γ-Strahlen. (German) [On scattering of shortwave γ rays]. Zeitschrift für Physik, 84(3–4):137–144, March 1933. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1956:IJC

[ML56] Lise Meitner and E. Lübcke. Irène Joliot-Curie/Ferdinand Trendelenburg 60 jahre. (German) [Irène Joliot-Curie/Ferdinand Trendelenburg 60th birthdays]. Physik Journal, 12(6):269–270, June 1956. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Morgenweck-Lambrinos:2000:LMO

[MLT00] Vera Morgenweck-Lambrinos and Martin Trömel. Lise Meitner, Otto Hahn und die Kernspaltung: eine Legende aus unseren Tagen. (German) [Lise Meitner, Otto Hahn and nuclear fission: a legend of our time]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 8(1):65–76, March 2000. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02914175.

Morgenweck-Lambrinos:2001:WLN

[MLT01] Vera Morgenweck-Lambrinos and Martin Trömel. Wissenschaft und Legende: eine Nachbetrachtung zu “Lise Meitner, Otto Hahn und die Kernspaltung: eine Legende aus unseren Tagen”. (German) [Science and legend: a review of “Lise Meitner, Otto Hahn and the nuclear fission: a legend from our days”]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 9(1):29–40, March 2001. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (electronic). URL http://link.

springer.com/article/10.1007/BF02911846.

Meitner:1928:BGR

[MnBB28] Lise Meitner, A. G. ntherschulze, H. Boegehold, and E. Brennecke. Besprechungen. (German) [Review]. Die Naturwissenschaften, 16(12):205–208, March 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1930:ABE

[MO30] Lise Meitner and Wilhelm Orthmann. Über eine absolute Bestimmung der Energie der primären β-Strahlen von Radium E. (German) [On an absolute determination of the energy of the primary β rays of Radium E]. Zeitschrift für Physik, 60(3–4): 143–155, March 1930. CODEN ZEPYAA. ISSN 0044-3328.

Montwill:1998:DYN

[Mon98] A. Montwill. Book review: The Defining Years in Nuclear Physics, 1932–1960s, by M. Mladjenovic. European Journal of Physics, 19(6):??, November 1998. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://stacks.

iop.org/0143-0807/19/i=6/a=014.

Morgenweck:1998:URT

[Mor98] Vera Morgenweck. Übersicht über die Radioaktivitätsforschung und über die theoretischen Grundlagen bis zur Entdeckung der Kernspaltung: Zur Einführung in die Fachliteratur. (German) [Review of radioactivity research and the theoretical foundations of the discovery of nuclear fission: Introduction to specialist literature]. Materialien zum Vortrag im Seminar von Prof. Trömel, Institut für Anorganische Chemie, Universität Frankfurt a. Main, Spring 1998. URL http://www.

sf-project.de/morgenweck/downloads/geschichte%20der%
20radioaktivitaet_theoretische%20grundlagen.pdf.

Morgenweck:2011:RLS

[Mor11]  Vera Morgenweck.  Ruth Lewin Simes Darstellung des

Verhältnisses Lise Meitner–Otto Hahn. (German) [Ruth Lewin Simes’s representation of the relationship Lise Meitner–Otto Hahn]. Wissenschaftliche Hausarbeit für das Lehramt an Haupt- und Realschulen (L2) eingereicht dem Wissenschaftlichen Prüfungsamt für das Lehramt an Grund-Schulen, Haupt- und Realschulen in Frankfurt am Main, August 21, 2011. URL http:

//www.psychologie.uni-frankfurt.de/61744137/Antragsunterlagen_
fuer_die_Wissenschaftliche_Hausarbeit.pdf. Abgabeter-

min: 21. September 1998, Überarbeitet (2011).

Meitner:1931:STG

[MP31] Lise Meitner and Kurt Philipp. Das γ-Spektrum von ThC” und die gamowsche Theorie der α-Feinstruktur. (German) [The γ spectrum of ThC” and the Gamow theory of α fine structure]. Die Naturwissenschaften, 19(50):1007, December 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1932:WZN

[MP32] Lise Meitner and Kurt Philipp. Über die Wechselwirkung zwischen Neutronen und Atomkernen. (German) [On the interaction between neutrons and atomic nuclei]. Die Naturwissenschaften, 20(51):929–932, December 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1933:STT

[MP33a] Lise Meitner and Kurt Philipp. Die γ-Strahlen von ThC und ThC” und die Feinstruktur der α-Strahlen. (German) [The γ rays of ThC and ThC” and the fine structure of the α rays]. Zeitschrift für Physik, 80(5–6):277–284, May 1933. CODEN ZEPYAA. ISSN 0044-3328.

Meitner:1933:APE

[MP33b] Lise Meitner and Kurt Philipp. Die Anregung positiver Elektronen durch γ-Strahlen von ThC”. (German) [The excitation of positive electrons by γ-radiation of ThC”]. Die Naturwissenschaften, 21(24):468, June 1933. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1933:NAE

[MP33c] Lise Meitner and Kurt Philipp. Die bei Neutronenanregung auftretenden Elektronenbahnen. (German) [The electron orbits occurring during neutron excitation]. Die Naturwissenschaften, 21(15):286–287, April 1933. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1934:WVN

[MP34] Lise Meitner and Kurt Philipp. Weitere Versuche mit Neutronen. (German) [Further experiments with neutrons]. Zeitschrift für Physik, 87(7–8):484–497, July 1934. CODEN ZEPYAA. ISSN 0044-3328.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Meitner:1938:KUB

[MSH38] Lise Meitner, Fritz Strassmann, and Otto Hahn. Künstliche Umwandlungsprozesse bei Bestrahlung des Thoriums mit Neutronen; Auftreten isomerer Reihen durch Abspaltung von α-Strahlen. (German) [Artificial transformation processes on irradiation of thorium with neutrons; appearance of isomeric series by splitting off of α-rays]. Zeitschrift für Physik, 109(7–8):538– 552, July 1938. CODEN ZEPYAA. ISSN 0044-3328.

Meyer-Ter-Vehn:1988:ESP

[MTVB88] J. Meyer-Ter-Vehn and Th. Bührke. EIR und SIN zum Paul-Scherrer-Institut zusammengelegt/Stuttgarter Max-Planck-Institute erweitert/Geheimer Fortschritt bei der Kernfusion/Nobelpreis: Preisgeld erhöht/Océ-van der Grinten-Preis 1988/Grünes Licht für den Bau eines Infrarot-Observatoriums/Neu. (German) [EIR and SIN merged to form the Paul Scherrer Institute/Stuttgart Max Planck Institute expanded/Secret advances in nuclear fusion/Nobel Prize: prize money increased/Océ van der Grinten Prize 1988/Green light for the construction of an infrared observatory/news]. Physik Journal, 44(5):128–154, May 1988. CO-DEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Meitner:1933:IPS

[MW33] Lise Meitner and Kan Chang Wang. Der innere Photoeffekt der γ-Strahlen. (German) [The inner photoelectric effect of γrays]. Die Naturwissenschaften, 21(32):594, August 1933. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:1955:EOH

[MWF55] Lise Meitner, Heinrich Wieland, and W. Fricke. Die ersten Otto-Hahn-Preise an Lise Meitner und Heinrich Wieland/Hans Kienle 60 Jahre. (German) [The first Otto Hahn Prize to Lise Meitner and Heinrich Wieland/Hans Kienle’s 60th birthday]. Physik Journal, 11(11):518–520, November 1955. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Meitner:1961:BKD

[MWZG61] L. Meitner, W. H. Westphal, J. E. Zürich, and J. Groeneveld. Bonzenfreies Kolloquium/Drucktechnische Fragen eines Autors/ Weitere Bissigkeiten zu Schul- und Bildungsfragen. (German) [Bonnet Colloquium/an author’s technical writing questions/ further school education issues]. Physik Journal, 17(4):198–199, April 1961. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Navarro:2006:EAD

[Nav06] Jaume Navarro. Early attempts to detect the neutrino at the Cavendish Laboratory. Physics in Perspective (PIP), 8(1):64–82, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-005-0249-z.

Newhouse:1998:CMC

[NCSW98] Irene Newhouse, Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Caricature of Meitner countered by drawing on historical record. Physics Today, 51(2):15, February 1998. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Nelson:2014:ARE

[Nel14] Craig Nelson. The age of radiance: the epic rise and dramatic fall of the atomic era. Scribner, New York, NY, USA, 2014. ISBN 1-4516-6043-X (hardcover), 1-4516-6044-8 (paperback), 1-4516-6045-6 (e-book). 4a37 pp. LCCN QC773 .N45 2014.

Neuberg:1921:FKW

[Neu21] C. Neuberg, editor. Festschrift der Kaiser Wilhelm Gesellschaft zur Förderung der Wissenschaften zu ihrem ehnjährigen Jubiläum Dargebracht von ihren Instituten. (German) [Festschrift of the Kaiser Wilhelm Society for the Advancement of Science on the occasion of its anniversary Presented by its institutes]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1921.

Nelson:2008:CCR

[NFG^{+}08] S. L. Nelson, C. M. Folden III, K. E. Gregorich, I. Dragojević, Ch. E. Düllmann, R. Eichler, M. A. Garcia, J. M. Gates, R. Sudowe, and H. Nitsche. Comparison of complementary reactions for the production of ^{261,262}Bh. Physical Review C (Nuclear Physics), 78(2), August 2008. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Nelson:2009:CCR

[NGD^{+}09] S. L. Nelson, K. E. Gregorich, I. Dragojević, J. Dvořák, P. A. Ellison, M. A. Garcia, J. M. Gates, L. Stavsetra, M. N. Ali, and H. Nitsche. Comparison of complementary reactions in the production of Mt. Physical Review C (Nuclear Physics), 79(2), February 2009. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Nigro:2003:BRL

[Nig03] Maria Nigro. Book review: Lise Meitner, by Hannelore Sexl and Anne Hardy, Rowohlt Taschenbuch, 2002, 157 pp., illus., $8.50 (paperback). History and Philosophy of the Life Sciences, 25(4):

562–563, 2003. ISSN 0391-9714 (print), 1742-6316 (electronic).

URL http://www.jstor.org/stable/23333258.

Noddack:1934:EGE

[Nod34] Ida Noddack. Über das Element 93. (German) [On element 93 [neptunium]]. Zeitschrift für Angewandte Chemie, 47(37):653– 655, September 15, 1934. ISSN 0932-2132. URL http://en.

wikipedia.org/wiki/Neptunium;   http://onlinelibrary.
wiley.com/doi/10.1002/ange.19340473707/abstract;   http:
//www.chemteam.info/Chem-History/Noddack-1934.html. Ac-

cording to Frisch [FW67, page 48], “Ida Noddack, a German chemist, quite rightly pointed out that they might be lighter elements [after bombardment of uranium by neutrons]; but her comments (published in a journal not much read by chemists and hardly at all by physicists) were regarded as mere pedantry. She did not indicate how such light elements could be formed; her paper had probably no effect whatever on later work.”. English translation in [GA71, pages 16–20].

Nye:1997:BRL

[Nye97] Mary Jo Nye. Book review: Lise Meitner: A Life in Physics, by Ruth Lewin Sime. The American Historical Review, 102(4): 1128–1129, 1997. ISSN 0002-8762 (print), 1937-5239 (electronic).

URL http://www.jstor.org/stable/2170646.

Offereins:2011:LM

[Off11] Marianne Offereins. Lise Meitner (1878–1968). In Apotheker and Sarkadi [AS11], pages 69–74. ISBN 3-527-32956-0, 3-527-63645-5 (e-book), 3-527-63647-1 (e-book). LCCN QD20 .E976 2011.

Olsen:1963:TGS

[Ols63] Arthur J. Olsen. Trackdown of the German scientist: Nazism and defeat scattered the leaders of a once-great scientific establishment. Herewith a review of where they went and what some of them did when they got there. New York Times, ??(??): 214, September 22, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/116463805/.

Oreskes:2003:GBI

[Ore03] Naomi Oreskes. Gender bias and Ida Noddack: Response. Science, 301(5636):1045–1046, August 2003. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See [Hoo03].

Orgel:1946:SLM

[Org46] Irene Orgel. Sonnet: To Lise Meitner. American Scholar, 15(2): 146, 1946. ISSN 0003-0937 (print), 2162-2892 (electronic). URL

http://www.jstor.org/stable/41204773.

Osgood:1966:BRC

[Osg66] Thomas H. Osgood. Book review: The Collected Papers of Lord Rutherford of Nelson, Vol. 3: The Cavendish Laboratory, Under scientific direction of Sir James Chadwick. Physics Today, 19(5): 93–95, May 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Parker:2002:MKN

[PC02] Kerry Parker and Steven Chapman. On the map: Kharkiv National University of Radioelectronics, Ukraine. starting out: Preparation, preparation, preparation. teaching anecdotes: Lise Meitner. personality: The accidental astronomer. Physics Education, 37(3):264–268, May 2002. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.

iop.org/0031-9120/37/i=3/a=603.

Popp:2023:PGU

[PdK23] Manfred Popp and Piet de Klerk. The peculiarities of the German uranium project (1939–1945). Journal of Nuclear Engineering, 4(3):634–653, September 2023. CODEN ???? ISSN 2673-4362.

Pearson:2015:BDF

[Pea15] J. Michael Pearson. On the belated discovery of fission. Physics Today, 68(6):40–45, June 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See reply and response [SP15].

Peierls:1986:FFW

[Pei86] Rudolf Peierls. Fact or fission [wild speculation]. London Review of Books, 8(18):4, October 23, 1986. ISSN 0260-9592. URL

https://www.lrb.co.uk/v08/n18/letters#letter4.

Perutz:1997:LME

[Per97] Max F. Perutz. Lise Meitner und die Entdeckung der Kernspaltung. (German) [Lise Meitner and the discovery of nuclear fission]. Naturwissenschaftliche Rundschau, 50(12):463–470, December 1997. CODEN NARSAC. ISSN 0028-1050.

Philipp:1933:AGA

[PMFB33] Kurt Philipp, Lise Meitner, H. Fränz, and W. Bothe. Atomkerne. (German) [Atomic nucleii]. In Elektronen · Atome Ionen. (German) [Electrons · Atoms Ions], pages 90–200. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1933.

Popp:2016:HAS

[Pop16] Manfred Popp. Hitlers Atombombe. Störfall der Wissenschaftsgeschichte. (German) [Hitler’s atomic bomb. Accident in the history of science]. Spektrum der Wissenschaft (German translation of Scientific American), 12(16):12–21, ???? 2016. CO-DEN SPEKDI. ISSN 0170-2971. URL https://www.spektrum.

de/inhaltsverzeichnis/hitlers-atombombe-spektrum-der-
wissenschaft-12-2016/1373070.

Popp:2021:WHD

[Pop21] Manfred Popp. Why Hitler did not have atomic bombs. Journal of Nuclear Engineering, 2(1):9–27, March 2021. CODEN ???? ISSN 2673-4362.

Radvanyi:1999:AOL

[Rad99] Pierre Radvanyi. Analyses d’ouvrages: Lise Meitner: A life in physics, ^{≪}California Studies in the history of science ^{≫}, vol. 13 par Ruth Lewin Sime. Revue d’Histoire des Sciences, 52(1):170– 173, January 1999. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23633583.

Radvanyi:2004:AOL

[Rad04] Pierre Radvanyi. Analyses d’ouvrages: Lise Meitner and the dawn of the nuclear age. [1^{re} éd. 1990, en allemand] par Patricia Rife; John Archibald Wheeler. Revue d’Histoire des Sciences, 57 (1):220–224, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23634145.

Rasch:1981:SMB

[Ras81] J. Rasch. Eine Straßmann-Biographie: Im Schatten der Sensation — Leben und Wirken von Fritz Straßmann. Von F. Krafft. Verlag Chemie, Weinheim-Deerfield Beach, Florida-Basel 198, 541 S., 48 Abb., zahlr. Faksimiles u. Tab., DM 150,-. Nachrichten aus Chemie, Technik und Laboratorium, 29(10):715–716, October 1981. CODEN NCTLDI. ISSN 0341-5163.

Rayner-Canham:1990:PWN

[RCRC90] M. F. Rayner-Canham and G. W. Rayner-Canham. Pioneer women in nuclear science. American Journal of Physics, 58(11): 1036–1043, November 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://adsabs.harvard.

edu/abs/1990AmJPh..58.1036R;   http://scitation.aip.
org/content/aapt/journal/ajp/58/11/10.1119/1.16269.

Rayner-Canham:1997:DTS

[RCRC97] Marelene F. Rayner-Canham and Geoffrey Rayner-Canham, editors. A devotion to their science: pioneer women of radioactivity. McGill-Queen’s University Press and Chemical Heritage Foundation, Montréal, Québec, Canada and Philadelphia, PA, USA, 1997. ISBN 0-7735-1608-5 (McGill hardcover), 0-7735-1642-5 (McGill paper), 0-941901-16-5 (Chemical Heritage hardcover), 0-941901-15-7 (Chemical Heritage paper). xiv + 307 pp. LCCN QC15 .R39 1997.

Rechenberg:1991:ELMa

[Rec91a] H. Rechenberg. Ehrung für Lise Meitner und Werner Heisenberg: Aufnahme der Büsten von Lise Meitner und Werner Heisenberg in den Ehrensaal des Deutschen Museums. (German) [Tribute to Lise Meitner and Werner Heisenberg: Inauguration of the busts of Lise Meitner and Werner Heisenberg in the Hall of Honor of the German Museum]. Physikalische Blätter, 47(8):773, August 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19910470810/abstract.

Rechenberg:1991:ELMb

[Rec91b] H. Rechenberg. Ehrung für Lise Meitner und Werner Heisenberg: Aufnahme der Büsten von Lise Meitner und Werner Heisenberg in den Ehrensaal des Deutschen Museums. (German) [Tribute to Lise Meitner and Werner Heisenberg: Inauguration of the busts of Lise Meitner and Werner Heisenberg in the Hall of Honor of the German Museum]. Physik Journal, 47(8):773, August 1991. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Rechenberg:2002:BRG

[Rec02] Helmut Rechenberg. Book review: Great Physicists — The Life and Times of Leading Physicists from Galileo to Hawking, by William H. Cropper. European Journal of Physics, 23(6):651, November 2002. CODEN EJPHD4. ISSN 0143-0807 (print),

1361-6404 (electronic). URL http://stacks.iop.org/0143-
0807/23/i=6/a=702.

Reed:2005:PNF

[Ree05] Rosemarie Reed. The path to nuclear fission: The story of Lise Meitner and Otto Hahn. Filmakers Library documentary film (56

minutes), 2005. URL http://alsos.wlu.edu/information.
aspx?id=3042.

Reed:2014:MP

[Ree14] Bruce Cameron Reed. The Manhattan Project. Physica Scripta, 89(10):108003:1–108003:26, October 2014. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic). URL http://

stacks.iop.org/1402-4896/89/i=10/a=108003.

Reed:2015:ABS

[Ree15a] Bruce Cameron Reed. The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical game-changer. IOP concise physics. Morgan and Claypool Publishers and IOP Publishing, San Rafael, CA, USA and Bristol, UK, 2015. ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). 239 (est.) pp. LCCN QC773.3.U5 R443 2015eb. URL

http://iopscience.iop.org/book/978-1-6270-5991-6.

Reed:2015:BS

[Ree15b] Bruce Cameron Reed. The background science. In The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical game-changer [Ree15a], pages 2:1– 2:30. ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). LCCN QC773.3.U5 R443 2015eb. URL http://

iopscience.iop.org/book/978-1-6270-5991-6.

Reines:1972:CFO

[Rei72] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reingold:1984:MMA

[Rei84] Nathan Reingold. MGM meets the atomic bomb. The Wilson Quarterly (1976–2012), 8(4):154–163, Autumn 1984. ISSN 0363-

3276. URL http://www.jstor.org/stable/40256804.

Reinhardt:2007:CSC

[Rei07] Carsten Reinhardt. Chemical sciences in the 20th century: bridging boundaries. Wiley-VCH, New York, NY, USA, 2007. ISBN 1-281-84243-5, 3-527-30271-9, 3-527-61273-4 (e-book). xviii + 281 pp. LCCN QD15 .C44 2001. URL http://sfx.

ethz.ch/sfx\_locater?sid=ALEPH:EBI01\%26genre=book\%
26isbn=9783527302710\%26id=doi:10.1002/9783527612734.

Rentetzi:2004:CCS

[Ren04] Maria Rentetzi. The city as a context for scientific activity: creating the Mediziner-Viertel in fin-de-siècle Vienna. Endeavour, 28(1):39–44, March 2004. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Rife:1990:LML

[Rif90] Patricia Rife. Lise Meitner: ein Leben für die Wissenschaft. (German) [Lise Meitner: a life for science]. Claassen, Düsseldorf, Germany, 1990. ISBN 3-546-47786-3. 349 pp. Translation to German by Peter Jacobs.

Rife:1992:LML

[Rif92] Patricia Rife. Lise Meitner: ein Leben für die Wissenschaft. (German) [Lise Meitner: a life for science]. Claassen, Hildesheim, Germany, second edition, 1992. ISBN 3-546-00029-3. 350 pp. Translation to German by Peter Jacobs.

Rife:1999:LMD

[Rif99] Patricia Rife. Lise Meitner and the dawn of the nuclear age. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-3732-X. xviii + 434 pp. With a foreword by John Archibald Wheeler.

Rife:2006:LMD

[Rif06] Patricia Rife. Lise Meitner and the Dawn of the Nuclear Age. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2006. ISBN 0-8176-4559-4 (softcover), 0-8176-3732-X (hardcover). xviii + 432 + 20 pp. LCCN

QC774.M4 R54 1999. URL http://www.loc.gov/catdir/
enhancements/fy0812/2006935949-d.html;   http://www.
loc.gov/catdir/enhancements/fy0812/2006935949-t.html.

Foreword by John Archibald Wheeler.

Russ:1909:AUR

[RM09] Sidney Ruß and Walter Makower. Die Ausstoßung radioaktiver Materie bei den Umwandlungen des Radiums. (German) [Exposure of radioactive matter by conversions of radium]. Physikalische Zeitschrift, 10(16):568, August 15, 1909. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/nyp.
33433090815253?urlappend=;seq=609. Expansion of remarks

by Otto Hahn and Lise Meitner [HM09a] on our work.

Regener:1927:BGR

[RMJ27] E. Regener, L. Meitner, and P. Jordan. Besprechungen. (German) [Review]. Die Naturwissenschaften, 15(38):789–792, September 1927. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Robison:1995:RMX

[Rob95] Roger F. Robison. The race for megavoltage X-rays versus Telegamma. Acta Oncologica (Stockholm), 34(8):1055–1074, 1995. CODEN ACTOEL. ISSN 0284-186X (print), 1651-226X

(electronic). URL https://www.tandfonline.com/doi/abs/
10.3109/02841869509127233.

Romer:1988:EMW

[Rom88] Robert H. Romer. Editorial: 958 men, 93 women — how many Lise Meitners among those 865? American Journal of Physics, 56(10):873–874, October 1988. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Rosenfeld:1972:NR

[Ros72] Léon Rosenfeld. Nuclear reminiscences. In Reines [Rei72], pages 289–299. ISBN 0-87081-025-1. LCCN QC780 .C65. URL http:

//adsabs.harvard.edu/abs/1972cht..conf..289R.

Rouvray:1996:BRL

[Rou96] Dennis H. Rouvray. Book review: Lise Meitner: A life in physics, by R. L. Sime. Pp. 526. University of California Press, 1996. £30.00. ISBN 0-520-0890-5. Endeavour, 20(3):137, January 1996. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Rosenthal-Schneider:1980:RST

[RS80] Ilse Rosenthal-Schneider. Reality and scientific truth: discussions with Einstein, von Laue, and Planck. Wayne State University Press, Detroit, MI, USA, 1980. ISBN 0-8143-1650-6. 148 pp. LCCN QC6 .R5696. Edited by Thomas Braun. Foreword by Arthur I. Miller.

Saperstein:2014:BRA

[Sap14] Alvin Saperstein. Book review: The Age of Radiance: The Epic Rise and Dramatic Fall of the Atomic Era, Craig Nelson, Scribner, 2014. $29.99 (448 pp.). ISBN 978-1-4516-6043-2. Physics Today, 67(9):45, September 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Scerri:2013:TE

[Sce13] Eric R. Scerri. A tale of 7 elements. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-539131-4 (hardcover), 0-19-987503-0 (e-book). xxxiii + 270 pp. LCCN

QD467 .S3475 2013. URL http://catalogue.bnf.fr/ark:/
12148/cb436625450.

Schroder:1957:ANP

[Sch57] Gottfried Schröder, editor. Das Atom und die neue Physik. (German) [The atom and the new physics]. List Verlag, Berlin and Frankfurt, West Germany, 1957. LCCN QC776 .A865 1957.

Schwabe:1960:BPC

[Sch60] K. Schwabe. Beiträge zur Physik und Chemie des 20. Jahrhunderts. Lise Meitner, Otto Hahn, Max von Laue zum 80. Geburtstag. Herausgeg. von O. R. Frisch, F. A. Paneth, F. Laves. P. Rosbaud. XI + 285 S. m. 89 Abb. u. 3 Porträts. Braunschweig 1959. Friedr. Vieweg & Sohn. Preis geb. 29,50 DM. Zeitschrift für Angewandte Mathematik und Mechanik, 40(10–11):519–520, 1960. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Schluter:1991:PFK

[Sch91] Anne Schlüter. Pionierinnen — Feministinnen — Karrierefrauen?: 90 Jahre Frauenstudium in Deutschland. (German) [Pioneers — feminists — career women?: 90 years of studying women in Germany], volume 22 of Frauen in Geschichte und Gesellschaft. Centaurus-Verl.-Ges., Pfaffenweiler, Germany, 1991. ISBN 3-89085-419-2. 352 pp. LCCN HQ1181.G3 P65 1992.

Scheich:1997:SPM

[Sch97] Elvira Scheich. Science, politics, and morality: The relationship of Lise Meitner and Elisabeth Schiemann. Osiris (Series 2), 12 (??):143–168, ???? 1997. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301903.

Schlogl:1998:SF

[Sch98] Friedrich Schlögl. Siegfried Flügge. Physics Today, 51(8):77, August 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Seelig:1956:HZD

[See56] Carl Seelig, editor. Helle Zeit — dunkle Zeit: in memoriam Albert Einstein. (German) [Bright Times — Dark Times: in memory of Albert Einstein]. Europa Verlag, München, Germany and Zürich, Switzerland, 1956. ISBN 3-322-84225-8. 171 pp. LCCN ???? With contributions and letters from Hedwig Born.

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452 (1):xv–xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Segre:1993:MAM

[Seg93] Emilio Segrè. A mind always in motion: the autobiography of Emilio Segrè. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1993. ISBN 0-520-07627-3. xii + 332 pp. LCCN QC16.S35 A3 1993. URL

http://ark.cdlib.org/ark:/13030/ft700007rb;   http://
www.loc.gov/catdir/description/ucal041/92010722.html.

Segre:2007:FCS

[Seg07] Gino Segrè. Faust in Copenhagen: a struggle for the soul of physics. Viking, New York, NY, USA, 2007. ISBN 0-670-03858-X. x + 310 pp. LCCN QC15 .S427 2007. URL http://www.

loc.gov/catdir/enhancements/fy0743/2006052807-b.html;
 http://www.loc.gov/catdir/enhancements/fy0743/2006052807-
d.html.

Sexl:2002:LML

[SH02a] Lore Sexl and Anne Hardy. Leseprobe aus: Monographie Lise Meitner. (German) [Excerpt from: Lise Meitner monograph]. In Lise Meitner [SH02b], pages 7–18, 143–145. ISBN 3-499-50439-1.

LCCN QC774.M4 S49 2002. URL https://www.rowohlt.de/
fm/131/Mono_Meitner.pdf.

Sexl:2002:LM

[SH02b] Lore Sexl and Anne Hardy. Lise Meitner, volume 50439 of Rowohlts Monographien. Rowohlt Taschenbuch, Reinbek bei Hamburg, Germany, 2002. ISBN 3-499-50439-1. 153 pp. LCCN QC774.M4 S49 2002.

Shea:1983:OHR

[She83] William R. Shea, editor. Otto Hahn and the Rise of Nuclear Physics, volume 22 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1584-X. 252 pp. LCCN QC773 .O87 1983. URL http:/

/link.springer.com/book/10.1007/978-94-009-7133-2.

Sime:1989:LMD

[Sim89] Ruth Lewin Sime. Lise Meitner and the discovery of fission. Journal of Chemical Education, 66(5):373–375, May 1989. CO-DEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

URL http://pubs.acs.org/doi/abs/10.1021/ed066p373.

Sime:1990:LME

[Sim90] Ruth Lewin Sime. Lise Meitner’s escape from Germany. American Journal of Physics, 58(3):262–267, March 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://scitation.aip.org/content/aapt/journal/ajp/58/
3/10.1119/1.16196. See also [Sim94b].

Sime:1991:LMF

[Sim91a] Ruth Lewin Sime. Lise Meitner and fission: Fallout from the discovery. Angewandte Chemie, International Edition in English, 30(8):942–953, August 1991. CODEN ACIEAY. ISSN 0570-0833.

Sime:1991:LMK

[Sim91b] Ruth Lewin Sime. Lise Meitner und die Kernspaltung: ,,Fallout” der Entdeckung. (German) [Lise Meitner and nuclear fission: “Fallout” of discovery]. Angewandte Chemie, 103(8):956– 967, August 1991. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Sime:1994:BRL

[Sim94a] Ruth Lewin Sime. Book review: Lise Meitner an Otto Hahn: Briefe aus den Jahren 1912 bis 1924: Edition und Kommentierung by Sabine Ernst. Isis, 85(4):726, December 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235348.

Sime:1994:LMS

[Sim94b] Ruth Lewin Sime. Lise Meitner in Sweden 1938–1960: Exile from physics. American Journal of Physics, 62(8):695– 701, August 1994. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://scitation.aip.

org/content/aapt/journal/ajp/62/8/10.1119/1.17498. See

[Sim90, Kac96, Tru96].

Sime:1996:LML

[Sim96a] Ruth Lewin Sime. Lise Meitner: a life in physics, volume 13 of California studies in the history of science. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1996. ISBN 0-520-08906-5 (hardcover), 0-520-20860-9 (paperback), 0-585-05524-6 (e-book). xiii + 526 pp. LCCN QC774.M4 S56 1996. URL http://www.loc.gov/

catdir/bios/ucal051/95035246.html;   http://www.loc.
gov/catdir/description/ucal041/95035246.html.

Sime:1996:MFG

[Sim96b] Ruth Lewin Sime. Meitner, Frisch got to fission theory first — the rest found it in Nature. Physics Today, 49(7):92, July 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Sime:1998:LMD

[Sim98a] Ruth Lewin Sime. Lise Meitner and the discovery of nuclear fission. Scientific American, 278(1):80–85 (Intl. ed. 58– ??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
80.pdf.

Sime:1998:LMK

[Sim98b] Ruth Lewin Sime. Lise Meitner und die Kernspaltung. (German) [Lise Meitner and the discovery of nuclear fission]. Sterne und Weltraum, ??(5):80–??, May 1, 1998. CODEN STUWAN. ISSN 0039-1263 (print), 1434-520X (electronic). URL

https://www.spektrum.de/magazin/lise-meitner-und-die-
kernspaltung/824545.

Sime:2000:STE

[Sim00] Ruth Lewin Sime. The search for transuranium elements and the discovery of nuclear fission. Physics in Perspective (PIP), 2(1):48–62, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050036;   http://
www.springerlink.com/content/nr2t13tndn6t9t72/.

Sime:2001:LML

[Sim01] Ruth Lewin Sime. Lise Meitner: ein Leben für die Physik. (German) [Lise Meitner: a life in physics]. Insel Verlag, Frankfurt am Main, Germany, 2001. ISBN 3-458-17066-9. 662 pp. LCCN ????

Sime:2002:LMC

[Sim02] Ruth Lewin Sime. Lise Meitner: a 20th century life in physics. Endeavour, 26(1):27–31, March 2002. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Sime:2003:FFM

[Sim03] Ruth Lewin Sime. From Fermi to fission: Meitner, Hahn and Strassmann in Berlin. In Anonymous [Ano03], pages 133–144. ISBN 88-8286-032-9. LCCN ???? URL http:old.enea.it/

produzione_scientifica/pdf.../V2003_AttiFermi.pdf.

Sime:2005:EPP

[Sim05] Ruth Lewin Sime. From exceptional prominence to prominent exception: Lise Meitner at the Kaiser Wilhelm Institute for Chemistry. In Susanne Heim, editor, Forschungsprogramm Geschichte der Max-Planck-Gesellschaft im Nationalsozialismus, Ergebnisse 24, page ???? Max-Planck-Gesellschaft, Berlin, Germany, 2005.

Sime:2006:PMO

[Sim06] Ruth Lewin Sime. The politics of memory: Otto Hahn and the Third Reich. Physics in Perspective (PIP), 8(1):3–51, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-004-0248-5.

Sime:2007:UGS

[Sim07a] Ruth Lewin Sime. “Die ‘Uranspaltung’ hat da die ganze Situation gerettet”: Otto Hahn und das Kaiser-Wilhelm-Institut für Chemie im Zweiten Weltkrieg. (German) [“Uranium fission has saved the whole situation”: Otto Hahn and the Kaiser Wilhelm Institute for Chemistry in the Second World War]. In Maier [Mai07], pages 268–304. ISBN 3-8353-0182-9 (paperback). LCCN Q180.G4 K38 2000 v.17; U395.G3 G45 2007.

Sime:2007:SAE

[Sim07b] Ruth Lewin Sime. The search for artificial elements and the discovery of nuclear fission. In Chemical sciences in the 20th century: bridging boundaries [Rei07], pages 146–159. ISBN 1-281-84243-5, 3-527-30271-9, 3-527-61273-4 (e-book). LCCN

QD15 .C44 2001. URL https://onlinelibrary.wiley.com/
doi/pdf/10.1002/9783527612734.ch08.

Sime:2010:IHN

[Sim10] Ruth Lewin Sime. An inconvenient history: the nuclear-fission display in the Deutsches Museum. Physics in Perspective (PIP), 12(2):190–218, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-009-0013-x;   http://www.
springerlink.com/content/w716842562715257/.

Sime:2012:PFO

[Sim12] Ruth Lewin Sime. The politics of forgetting: Otto Hahn and the German Nuclear-Fission Project in World War II. Physics in Perspective (PIP), 14(1):59–94, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/link.springer.com/article/10.1007/s00016-011-0065-6;
 http://www.springerlink.com/content/k12202vg92147h68/

.

Sime:2014:AEL

[Sim14a] Ruth Lewin Sime. Auswanderung und Exil: Lise Meitner in Schweden 1938–1960. (German) [Emigration and exile: Lise Meitner in Sweden 1938–1960]. In Skandinavien als Zuflucht für jüdische Intellektuelle 1933–1945. (German) [Scandinavia as a Refuge for Jewish Intellectuals 1933–1945] [DF14], pages 32–53. ISBN 3-86331-194-9. LCCN DS135.S32 S53 2014.

Sime:2014:SPD

[Sim14b] Ruth Lewin Sime. Science and politics: The discovery of nuclear fission 75 years ago. Annalen der Physik (1900), 526(3–4): A27–A31, April 2014. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1002/andp.201400805.

Strassmann:1980:FBIa

[SK80a] Fritz Straßmann and Fritz Krafft. Ein frühes Beispiel interdisziplinärer Teamarbeit (I): Zur Entdeckung der Kernspaltung durch Hahn, Meitner und Straßmann. (German) [An early example of interdisciplinary teamwork (I): The discovery of fission by Hahn, Meitner and Strassmann]. Physik Journal, 36(4):85–89, April 1980. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Strassmann:1980:FBIb

[SK80b] Fritz Straßmann and Fritz Krafft. Ein frühes Beispiel interdisziplinärer Teamarbeit (II): Zur Entdeckung der Kernspaltung durch Hahn, Meitner und Straßmann. (German) [An early example of interdisciplinary teamwork (II): The discovery of fission by Hahn, Meitner, and Strassmann]. Physik Journal, 36(5):113– 118, May 1980. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Shaik:2015:LMM

[SKDA15] Sason Shaik, Miri Karni, David Danovich, and Yitzhak Apeloig. The Lise Meitner-Minerva Center for Computational Quantum Chemistry: 18 years of Israeli–German collaboration. Israel Journal of Chemistry, 55(11–12):1167–1176, November 2015. CODEN ISJCAT. ISSN 0021-2148.

Stein:1926:BGR

[SM26] Emmy Stein and Lise Meitner. Besprechungen. (German) [Review]. Die Naturwissenschaften, 14(14):297–298, April 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Spangenburg:2008:NBA

[SM08] Ray Spangenburg and Diane Kit Moser, editors. Niels Bohr: Atomic Theorist. Makers of modern science. Chelsea House, New York, NY, USA, revised edition, 2008. ISBN 0-8160-6178-5. xvi + 141 pp. LCCN QC16.B63 S63 2008.

Santhosh:2017:DCS

[SN17] K. P. Santhosh and C. Nithya. α-decay chains of superheavy 265−−279Mt isotopes. Physical Review C (Nuclear Physics), 96 (4), October 2017. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Sime:2015:LMD

[SP15] Ruth Lewin Sime and J. Michael Pearson. Lise Meitner and the discovery of fission. Physics Today, 68(10):10–11, October 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [Pea15].

Sittig:1997:LIP

[SRRM^{+}97] R. Sittig, H. Rechenberg, M. Ritsch-Marte, H. Schopper, S. Koch, H. J. Queisser, H. Nicolai, and M. Ziegler. Lübbig: The Inverse Problem/Bührke: Newtons Apfel — Sternstunden der Physik: Von Galilei bis Lise Meitner/Memel, Renneberg: Geschlechterverhältnisse in Medizin, Naturwissenschaft und Technik/Sanitt: Science as a Questioning Process/Haug, Jauho: Quantum Kinetics in Transport and Optics of Semiconductors/High Tech Start Up/Gambini, Pullin: Loops, Knots, Gauge Theories and Quantum Gravity/MuPAD. (German) [Book reviews: …]. Physik Journal, 53(7–8):716–719, July 1997. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Strassmann:1968:LME

[ST68] F. Straßmann and Erich Töpfer. Lise Meitner/Ernst Lamla 80 Jahre. (German) [Birthdays …]. Physik Journal, 24(11):508– 509, November 1968. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Stewart:2000:BRL

[Ste00] Albert B. Stewart. Book review: Lise Meitner and the Dawn of the Nuclear Age, by Patricia Rife. The Antioch Review, 58 (2):240–241, Spring 2000. ISSN 0003-5769 (print), 2326-9707

(electronic). URL http://www.jstor.org/stable/4614000.

Stolz:1989:OHL

[Sto89] Werner Stolz. Otto Hahn/Lise Meitner, volume 64 of Biographien hervorragender Naturwissenschaftler, Techniker und Mediziner. Vieweg + Teubner Verlag, Braunschweig, Germany, second edition, 1989. ISBN 3-322-00685-9, 3-322-82223-0. 97 pp.

Stone:1997:CTE

[Sto97] Richard Stone. Chemistry: Transuranic element names finally final. Science, 277(5332):1601, September 12, 1997. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Strassmann:1954:ERL

[Str54] F. Strassmann. Zur Erforschung der Radioaktivität Lise Meitner zum 75. Geburtstag. (German) [Researching radioactivity: Lise Meitner’s 75th birthday]. Angewandte Chemie, 66(4):93– 95, February 1954. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Stuewer:1985:BNF

[Stu85] Roger H. Stuewer. Bringing the news of fission to America. Physics Today, 38(10):48–56, October 1985. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v38/i10/p48_s1.

Stuewer:1994:OLD

[Stu94] Roger H. Stuewer. The origin of the liquid-drop model and the interpretation of nuclear fission. Perspectives on Science, 2(1):76–129, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://alsos.wlu.edu/

information.aspx?id=1736. See errata [Ano94].

Stuart:1996:WNS

[Stu96] B. H. Stuart. Women in nuclear science. Physics Education, 31 (2):116–120, March 1996. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.iop.org/

0031-9120/31/i=2/a=022.

Stuewer:2013:ACM

[Stu13] Roger H. Stuewer. An act of creation: the Meitner–Frisch interpretation of nuclear fission. In Katzir et al. [KLR13], pages 231–245. ISBN 3-8442-5134-0. LCCN QC173.98. URL http://

www.edition-open-access.de/proceedings/5/.

Sweeney:2006:ANM

[Swe06] Seamus Sweeney. Book review: Anonymous no more: Scientists Anonymous: Great Stories of Women in Science, Patricia Fara, Wizard Books (2005), £6.99 (213 pages, paperback), ISBN: 1-84046-574-3. Endeavour, 30(1):7–8, March 2006. CODEN EN-DEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Teller:1955:WMP

[Tel55] Edward Teller. The work of many people. Science, 121(3139): 267–275, February 25, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1681529.

Tromel:1999:LBL

[TF99] Martin Trömel and Rudolf Fleischmann. Leserbriefe zu: ,,Lise Meitner und der Nobelpreis”. (German) [Readers’ letters on: “Lise Meitner and the Nobel Prize”]. Physik in unserer Zeit, 30(2):81, 1999. CODEN PHUZAH. ISSN 0031-9252.

Tromel:2009:VEK

[Trö09] Martin Trömel. Die vier Entdecker der Kernspaltung. (German) [The four discovers of nuclear fission]. Nachrichten aus der Chemie, 57(1):57, January 2009. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic). URL https://www.gdch.

de/fileadmin/downloads/Publikationen/Nachrichten_aus_
der_Chemie/PDFs/kernspaltung.pdf. See [Ell08].

Trulock:1996:MHH

[Tru96] David W. Trulock. Meitner, Hahn, Hitler, and Siegbahn. American Journal of Physics, 64(5):523–524, May 1996. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://scitation.aip.org/content/aapt/journal/ajp/64/

5/10.1119/1.18271. See [Sim94b, Kac96].

Sime:2015:FIA

[uRLS15] Dieter Hoffmann und Ruth Lewin Sime. Freundschaft, Interdisziplinarität, Ausgrenzung: Lise Meitner und Otto Hahn. (German) [Friendship, interdisciplinarity, exclusion: Lise Meitner and Otto Hahn]. In Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy] [LMOS15], pages 390–417. ISBN 3-11-037598-2 (hardcover). LCCN AS182.B38 B477 2015.

Baeyer:1911:MSS

[vBHM11a] O. v. Baeyer, Otto Hahn, and Lise Meitner. Magnetische Spektren der β-Strahlen des Radiums. (German) [Magnetic spectra of β rays of radium]. Physikalische Zeitschrift, 12(24):1099–1101, December 1, 1911. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/mdp.
39015011417709?urlappend=;seq=1157.

Baeyer:1911:NSR

[vBHM11b] O. v. Baeyer, Otto Hahn, and Lise Meitner. Nachweis von β-Strahlen bei Radium D. (German) [Evidence of β Radium D radiation]. Physikalische Zeitschrift, 12(10):378–379, May 15, 1911. CODEN PHZTAO. ISSN 0369-982X. URL https:

//hdl.handle.net/2027/mdp.39015011417709?urlappend=;
seq=418.

Baeyer:1911:SAN

[vBHM11c] O. v. Baeyer, Otto Hahn, and Lise Meitner. Über die β-Strahlen des aktiven Niederschlags des Thoriums. (German) [On the β rays of thorium active precipitation]. Physikalische Zeitschrift, 12(8):273–279, April 15, 1911. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/mdp.
39015011417709?urlappend=;seq=311.

Meitner:1912:MSS

[vBHM12] O. v. Baeyer, Otto Hahn, and Lise Meitner. Das magnetische Spektrum der β-Strahlen des Thoriums. Physikalische Zeitschrift, 13(7):264–266, April 1, 1912. CODEN PHZTAO.

ISSN 0369-982X. URL https://babel.hathitrust.org/cgi/
pt?id=mdp.39015023176806;view=1up;seq=326.

Baeyer:1913:MSS

[vBHM13] O. v. Baeyer, Otto Hahn, and Lise Meitner. Das magnetische Spektrum der β-Strahlen des Radioactiniums und seiner Zerfallsprodukte. (German) [The magnetic spectrum of β rays of radioactinium and its decay products]. Physikalische Zeitschrift, 14(8):321–323, April 15, 1913. CODEN PHZTAO.

ISSN 0369-982X. URL https://hdl.handle.net/2027/mdp.
39015021268936?urlappend=;seq=379.

Baeyer:1914:MSS

[vBHM14] O. v. Baeyer, Otto Hahn, and Lise Meitner. Das magnetische Spektrum der β-Strahlen des Uran X. (German) [The magnetic spectrum of the β rays of uranium X]. Physikalische Zeitschrift, 15(13):649–650, July 1, 1914. CODEN PHZTAO. ISSN 0369-

982X. URL https://hdl.handle.net/2027/uc1.c2831834?
urlappend=;seq=693.

Baeyer:1915:MSS

[vBHM15] O. v. Baeyer, Otto Hahn, and Lise Meitner. Das magnetische Spektrum der β-Strahlen von Radiothor und Thorium X. (German) [The magnetic spectrum of β rays of radiothorium and thorium X]. Physikalische Zeitschrift, 16(13):6–7, July 1, 1915. CO-DEN PHZTAO. ISSN 0369-982X. URL https://hdl.handle.

net/2027/njp.32101054770928?urlappend=;seq=26.

Grosse:1932:EIE

[vG32] Aristid v. Grosse. Zur Entdeckung und Isolierung des Elements 91. (German) [Toward the discovery and isolation of element 91]. Die Naturwissenschaften, 20(21):362–363, May 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL https://link.springer.com/article/10.1007/

BF01504928. See remarks [HM32].

vonLaue:1927:BRW

[vLM27] Max von Laue and Lise Meitner. Die Berechnung der Reichweitestreuung aus Wilson-Aufnahmen. (German) [Calculation of the range spread from Wilson recordings]. Zeitschrift für Physik, 41(6–7):397–406, October 1927. CODEN ZEPYAA. ISSN 0044-3328.

vonLippmann:1923:BGR

[vLvRM^{+}23] Edmund O. von Lippmann, M. v. Rohr, Lise Meitner, Arnold Heim, W. Funk, Vorwort, and I. Koppel. Besprechungen. (German) [Review]. Die Naturwissenschaften, 11(42):857–862, October 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

vonMeitner:1955:RWT

[vM55] Lise von Meitner. Radioisotope in wissenschaft und technik. (German) [Radioisotopes in science and technology]. Applied Scientific Research, Section B, 4(1):68–72, December 1955. CO-DEN ASRBAC. ISSN 0365-7140 (print), 1872-8065 (electronic).

vonOertzen:2002:LMP

[vO02] Wolfram von Oertzen. Lise Meitner Prize 2002. Europhysics News, 33(5):180, September/October 2002. CO-DEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL https://www.europhysicsnews.org/articles/
epn/pdf/2002/05/epn2002-33-5.pdf.

Vogt:2006:HHL

[Vog06] Annette Vogt. Vom Hintereingang zum Hauptportal?: Lise Meitner und ihre Kolleginen an der Berliner Universität und in der Kaiser-Wilhelm-Gesellschaft. (German) [From the back entrance to the main portal?: Lise Meitner and her colleagues at the University of Berlin and in the Kaiser Wilhelm Society], volume 17 of Pallas Athene. Steiner, Stuttgart, Germany, 2006. ISBN 3-515-08881-4. ISSN 1439-9857. 550 + 16 pp. LCCN Q130 .V64 2007.

Vogt:2014:LMI

[Vog14] Annette Vogt. Lise Meitner und ihre Kolleginnen im Exil. (German) [Lise Meitner and her colleagues in exile]. In Skandinavien als Zuflucht für jüdische Intellektuelle 1933–1945. (German) [Scandinavia as a Refuge for Jewish Intellectuals 1933–1945] [DF14], pages 54–80. ISBN 3-86331-194-9. LCCN DS135.S32 S53 2014.

Vormum:2010:WWL

[Vor10] Günther Vormum. Zum wissenschaftlichen Werk von Lise Meitner und Otto Hahn. (German) [On the scientific work of Lise Meitner and Otto Hahn]. Zeitschrift für Chemie (Stuttgart, Germany), 20(7):243–247, August 2010. CODEN ZECEAL. ISSN 0044-2402.

Walker:1996:BRT

[Wal96] Mark Walker. Book review: The trials of a woman’s life in science: Lise Meitner: a Life in Physics, by Ruth Lewin Sime, 1996, University of California Press, 526pp. Physics World, 9 (5):51, May 1996. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://stacks.iop.org/2058-
7058/9/i=5/a=27.

Walker:2003:OHV

[Wal03] Mark Walker. Otto Hahn: Verantwortung und Verdrängung. (German) [Otto Hahn: responsibility and repression]. Forschungsprogramm “Geschichte der Kaiser-Wilhelm-Gesellschaft im Nationalsozialismus” (Research Program “History of the Kaiser Wilhelm Society in the National Socialist Era”), Berlin, Germany,

2003. 64 pp. LCCN ???? URL www.mpiwg-berlin.mpg.de/
KWG/Ergebnisse/Ergebnisse10.pdf.

Walker:2006:OHR

[Wal06] Mark Walker. Otto Hahn: Responsibility and repression. Physics in Perspective (PIP), 8(2):116–163, May 2006. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://link.springer.com/article/10.1007/s00016-
006-0277-3.

Watkins:1983:LMB

[Wat83] Sallie A. Watkins. Lise Meitner and the beta-ray energy controversy: an historical perspective. American Journal of Physics, 51(6):551–553, June 1983. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Watkins:1997:LMF

[Wat97] Sallie A. Watkins. Lise Meitner: the foiled Nobelist. In Rayner-Canham and Rayner-Canham [RCRC97], chapter 16, pages 163– 191. ISBN 0-7735-1608-5 (McGill hardcover), 0-7735-1642-5 (McGill paper), 0-941901-16-5 (Chemical Heritage hardcover), 0-941901-15-7 (Chemical Heritage paper). LCCN QC15 .R39

1997. URL http://www.jstor.org/stable/j.ctt809w2.

Weiner:1972:MNP

[Wei72] Charles Weiner. 1932 — moving into the new physics. Physics Today, 25(5):40–49, May 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.

aip.org/content/aip/magazine/physicstoday/article/25/

5/10.1063/1.3070853. Reprinted in [Wei85].

Weiner:1985:MNP

[Wei85] Charles Weiner. 1932 — moving into the new physics. In Weart and Phillips [WP85], pages 332–339. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Wei72].

Weiss:1990:HMM

[Wei90] Dr.rer.nat. Burghard Weiss. Hahn und Meitner, Merton und Matthäus. Zur Namengebung einer deutschen Großforschungseinrichtung. (German) [Hahn and Meitner, Merton and Matthäus. The naming of a German research facility]. Berichte zur Wissenschaftsgeschichte, 13(4):219–231, ???? 1990. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Westfall:2001:LMD

[Wes01] Catherine Westfall. Lise Meitner and the Dawn of the Nuclear Age, by Patricia Rife (review). Technology and Culture, 42(2):375–377, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL http:

//www.jstor.org/stable/25147727;   https://muse.jhu.
edu/pub/1/article/33746.

Wheeler:2009:MF

[Whe09] John A. Wheeler. Mechanism of fission. Physics Today, 62(4): 35–38, April 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Winterberg:1996:MHN

[WHF^{+}96] Friedwardt Winterberg, Günter Herrmann, Igor Fodor, Lincoln Wolfenstein, and Mark E. Singer. More on how Nazi Germany failed to develop the atomic bomb. Physics Today, 49(1):11– 83, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wallot:1953:KWW

[WHFB53] J. Wallot, H. Hönl, A. Flammersfeld, and E. Brüche. Karl Willy Wagner/Gustav Mie 85 Jahre/Lise Meitner 75 jahre/Alexander Meissner 70 Jahre. (German) [Birthdays …]. Physik Journal, 9 (11):508–511, November 1953. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Whitten:2000:BRW

[Whi00] Barbara Whitten. Book reviews: Women Scientists in America: Before Affirmative Action, 1940–1972 by Margaret W. Rossiter; A Matter of Choices: Memoirs of a Female Physicist by Fay Ajzenberg-Selove; Lise Meitner: A Life in Physics by Ruth Lewin Sime. NWSA Journal, 12(3):203–208, Autumn 2000. ISSN 1040-0656 (print), 1527-1889 (electronic). URL http://www.

jstor.org/stable/4316773.

Ward:2015:RWF

[WM15] Tim Ward and Domenic Mastrippolito. Race for the world’s first atomic bomb: A thousand days of fear. [Los Alamos National] Laboratory produced documentary on the Man-

hattan Project., 2015. URL http://www.governmentattic.
org/11docs/NNSAfilmVideoList_2014.pdf;   https://www.
youtube.com/watch?v=xwpgmEvlRpM. Written by Michael Wil-

son. The hour-long program features interviews with some of those who worked on the Project Hans Bethe, Norman Ramsey, Phil Morrison, Edward Teller and others. The program profiles the creation of the Manhattan Project the obstacles that were faced, life in the “secret city,’ and events that led up to and followed the Trinity Test.

Wolff:2010:RHH

[Wol10] Stefan L. Wolff. Rezension: Vom Hintereingang zum Hauptportal? Lise Meitner und ihre Kolleginnen an der Berliner Universität und in der Kaiser-Wilhelm-Gesellschaft von Annette Vogt. (German) [Book review: From the back entrance to the main portal? Lise Meitner and her colleagues at the University of Berlin and in the Kaiser Wilhelm Society by Annette Vogt]. Berichte zur Wissenschaftsgeschichte, 33(4):441–443, December 2010. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Wurth:1999:BTS

[WSB^{+}99] Wilfried Wurth, Ruth Lewin Sime, Christian Bloch, Karl Joachim Schmidt-Tiedemann, Jan Louis, Klaus Lüders, Erhard Seiler, Karl-Heinz Rieder, Klaus Mainzer, Jürgen Schnack, and Filip Floegel. Bürger: Der Traum des Seglers bei Flaute/Lemmerich: Lise Meitner — Max von Laue Briefwechsel 1938–1948/Messing u. Huber: Die Doktorarbeit: Vom Start zum Ziel — Leitfaden für Promotionswillige/Quadbeck-Seeger: Faszination Innovation. Wichtiges und Wissensw. (German) [Book reviews: …]. Physik Journal, 55(7–8):105–110, July 1999. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Yount:2008:ZWS

[You08] Lisa Yount. A to Z of women in science and math. A to Z of women; Facts on File library of world history. Facts On File, New York, NY, USA, revised edition, 2008. ISBN 0-8160-6695-7 (hardcover). xv + 368 pp. LCCN ???? URL http://www.loc.

gov/catdir/toc/ecip0720/2007023966.html.

Zamfir:2014:LMP

[Zam14] Nicolae-Victor Zamfir. 2014 Lise Meitner Prize for Nuclear Science. Europhysics News, 45(5–6):8, ???? 2014. CO-DEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL https://www.europhysicsnews.org/articles/
epn/pdf/2014/05-06/epn201445-5-6.pdf.

Zimen:1979:OHL

[Zim79] Karl Erik Zimen. Otto Hahn, Lise Meitner und die Kernspaltung im Ausblick auf die Zukunft. (German) [Otto Hahn, Lise Meitner and nuclear fission in the perspective of the future]. Physik Journal, 35(5):200–210, May 1979. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Zilsel:1932:BGR

[ZSM32] E. Zilsel, Curt Stern, and L. Meitner. Besprechungen. (German) [Review]. Die Naturwissenschaften, 20(25):472–476, June 1932.

REFERENCES 139 CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).